A Complete Bibliography of Publications in Biometrics: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 September 2024 Version 2.01

Title word cross-reference

#33109 [1160].

1 [288]. 2 [671]. 2^{3} [827]. 2 × 2 [1420, 39, 895, 1038, 528, 1186]. 2 × 2 × 2 [330, 188]. 3^{2} [827]. 3^{4} [615]. ^{2} [760, 1049]. ^{3} [88]. ^{4} [212]. ^{n} [211]. χ^{2} [818]. F [748]. K [697, 1320]. Q [1099]. r × c [1069]. S [715]. S^{2}/X̄ [92]. Σ [760]. ∑(O − E)^{2}/E [574]. t [779, 1320]. T^{2} [414].

-based [1184]. -method [715]. -Ratio [1320]. -sample [697]. -Way [288].

/E [760].

10 [1154]. 1944 [763]. 1944-45 [763]. 1959 [385]. 1959-61 [385]. 1965 [194]. 1977 [1293]. 1979 [1275].

2 [353, 647]. 2. [1341]. 29 [1241].

3 [1309]. 3-Way [272]. 30 [1045, 1229]. 31 [1093]. 32 [885, 1064, 935, 955, 956, 1240, 973]. 33 [1160]. 35 [1339]. 3rd [512].

45 [763].

58 [1160].

61 [385]. 68 [309, 82].

8 [1225]. 81a [1339]. 8th [884].

A.O.V. [811]. abbreviated [69]. Abortion [201]. Abridged [370]. Abrupt [540]. absence [786]. Absolute [1383]. absorption [1329, 965]. Abstracts [9, 10, 11, 144, 145, 146, 147, 297, 298, 299, 474, 475, 476, 602, 603, 604, 720, 721, 722, 862, 863, 982, 983, 984, 1117, 1118, 1119, 1120, 1265, 1266, 1267, 194]. Abundance [1262, 640, 1322, 1343, 1344, 706]. Acceptance [192, 202]. Accessing [1237]. According [438]. Accounts [196]. Acoustic [559]. Action [632, 438]. Acute [1348, 557]. Adaptive [1096]. Additional [646]. Additive [1163]. Additives [72]. Additivity [305, 1059, 441, 579]. Adjacent [238, 240]. Adjunct [1399]. Adjustment [438, 581]. Adler [848]. Admits [437]. Advanced [369]. Advantages [1058]. Adverse [271]. Aerial [1312]. Affected [1236]. Affecting [1233]. Affine [1161]. Affinities [526]. Afifi [477, 1363]. Aflatoxin [950]. after [1149, 375]. Age [168, 35, 501, 387, 1376, 891, 676]. age-dependent [891, 676]. Agents [595]. Aggregation [964, 713]. Agreement [1071, 1072, 979, 786]. Agricultural [900, 904, 1179]. Agriculture [268, 517]. agronomiques [89]. Aid [8, 1173, 1177]. Aids [746]. Air [1348]. Aitchison [341, 177]. Ajustement [438]. al. [306]. Albert [303, 389, 78]. ale [149]. Alfred [600]. Alfredo [1002]. ALGOL [309, 82]. Algorithm [1053, 570]. Algorithme [570]. Algorithms [849]. Allele [1130, 45, 46, 1016, 1160, 359, 825]. Alleles [138]. Allelic [100]. Allocation [790, 400, 116, 1374, 1075]. Allometry [229]. Allow [1389]. allowing [891]. Alternative [1185, 238, 104, 687, 1074]. Alternatives [1249, 319]. Amble [904]. American [1270]. Among [357, 754, 1316, 1063, 1071]. Analogies [976]. Analogue [858]. analyse [774, 565]. Analyses [1018, 472]. Analysing [208, 243]. Analysis [716, 1260, 305, 307, 314, 477, 482, 728, 867, 1122, 1420, 34, 344, 1148, 178, 1012, 41, 1155, 638, 503, 353, 752, 756, 642, 1020, 645, 516, 776, 1028, 908, 1037, 1043, 1323, 1324, 392, 527, 206, 394, 791, 793, 1050, 71, 209, 6, 7, 213, 1058, 923, 541, 216, 1334, 675, 928, 1193, 413, 678, 229, 1341, 814, 93, 235, 417, 418, 818, 1067, 238, 1203, 1349, 1350, 95, 420, 1205, 96, 558, 821, 423, 1353, 1078, 103, 105, 249, 251, 689, 252, 254, 1208, 1358, 568, 436, 1209, 1362, 1363, 1369, 1370, 1373, 1086, 1219, 1220]. Analysis [950, 1387, 277, 701, 961, 453, 1395, 1235, 1398, 285, 456, 590, 1238, 462, 1247, 595, 290, 1105, 1252, 1107, 1407, 135, 856, 978, 857, 1412, 980, 1414, 1415, 1114, 1416, 860, 1115, 889, 894, 55, 788, 927, 942, 690, 832, 692, 565, 443, 1254, 1413, 859, 715, 18, 19, 747, 776, 1261, 498, 747, 1299, 1300, 50, 761, 1025, 1026, 193, 1047, 533, 661, 1190, 1367, 1418]. analytical [811]. Analyzing [1348, 1389, 1404]. Ancient [648]. Anderson [776, 533, 462]. Andrews [445]. Anesthesia [195]. Anesthetic [1148]. angular [809]. Animal [1262, 47, 1022, 904, 96, 560, 562, 1359, 1210, 447, 706]. Animals [1323]. Anita [586]. Annotated [381]. Announcements [26, 27, 28, 29, 169, 170, 171, 172, 337, 338, 339, 340, 490, 491, 492, 493, 620, 621, 622, 623, 736, 737, 738, 739, 875, 876, 877, 998, 999, 1000, 1001, 1137, 1138, 1139, 1140, 1286, 1287, 1288, 1289]. Annual [153, 154, 155]. ANOVA [748, 639, 1189, 938, 705]. Anthropology [304]. Antigen [43]. AOV [1064, 990, 933]. Aplicatii [149]. Apparent [467]. Apparentement [570]. Applicable [79, 237]. Application [612, 1125, 1129, 1418, 787, 539, 1057, 1062, 1071, 1077, 1220, 959, 584, 1235, 1238, 1403, 1243, 1251, 1105, 715, 1318, 1013, 557, 1384, 591, 1241]. Applications [1323, 663, 674, 89, 1083, 839, 449, 1391, 589, 1399, 1217, 348, 503, 516, 789, 421, 840, 1121, 898, 191]. Applied [1268, 32, 632, 500, 1299, 906, 391, 916, 1150, 498, 242, 689, 1367, 1234]. Approach [483, 1125, 1269, 184, 923, 1193, 413, 238, 696, 442, 456, 1242, 1074, 477, 514, 219, 1363]. Approaches [1293, 1043, 680, 686]. Appropriate [754, 1208, 971]. Approximate [182, 1159, 653, 1320, 1063, 573, 708, 711, 329]. Approximating [75]. Approximation [389, 574, 858, 1384]. Approximations [531, 567, 1347]. Arbitrary [987, 659, 1045]. Area [961]. Argument [320]. Arising [754, 1034, 234, 101]. Armitage [321]. Arnason [1341]. Arnold [1029]. Arrangements [305]. Array [1172]. Arthropod [1406]. Ascertainment [1236, 1100]. Ashford [561]. Aspects [504, 651, 211, 1197, 271, 1097, 1394, 1411, 1198, 270]. Assay [1020, 911, 109]. Assays [3, 419, 1077, 829, 159]. Assessment [1303, 526, 1071]. Assignment [871, 837, 1096, 971]. Associated [375, 819]. Association [20, 60, 1042, 195, 810, 247, 1212, 1213, 283]. Assorative [138]. Assortatively [642]. Assumption [56]. Assumptions [1100]. Asymmetrical [289]. Asymptotic [345, 356, 458, 690, 634]. Atherosclerosis [424]. Atkins [266]. Atlanta [557]. Attinger [463]. Attitudes [377]. Attributable [974]. Attribute [1202, 1369]. Auditor [1275]. Augmented [782, 783]. Aural [395]. Auswahl [971]. Auswertungsmethoden [221]. Autocorrelated [1155]. Autocorrelation [650]. Automatic [438]. Automatique [438]. Autopsy [1209]. Auxiliary [612, 1036, 937, 1070, 584]. Available [1128, 966, 1101, 1102, 1240]. Avoidance [1057]. Avoidance-Modified [1057]. Azen [477, 1363].

B [308, 317, 1124, 50, 51, 361, 901, 903, 1171, 198, 379, 794, 1190, 1191, 82, 1352, 558, 1362, 704, 1232]. Baboons [596]. Background [367]. Bacteria [818]. Bahn [586]. Bailey [1232, 1211]. Baker [1309]. Balaam [510]. Balance [1201]. Balanced [36, 40, 264, 714, 932]. Balancing [871, 912, 837]. Bancroft [176, 380, 553]. band [891]. band-return [891]. Bands [502]. Baniuk [1341]. Bánki [1174]. Barbensi [740]. Barlow [537]. Barner [1195]. Barnett [755]. Barra [1171]. Barrie [588]. Bartholomew [537]. Bartlett [665]. Based [61, 62, 532, 540, 802, 687, 276, 1184, 809, 942]. Basic [173, 76, 586]. Basis [190, 1412, 397, 1413]. Batschelet [767, 422]. Baum [1051]. Bayes [345, 555]. Bayesian [483, 1011, 349, 1074, 442, 469, 383, 661]. Bayo [385]. Be [1214, 495, 496, 625, 626, 627, 628, 741, 742, 743, 744, 878, 879, 880, 1004, 1005, 1006, 1007, 1143, 1144, 1145, 1146, 1290, 1291, 1292]. Bean [64]. Beck [1029]. Becker [797]. Beehive [564, 853]. before [375]. Begon [1262]. behavior [345, 458]. Behavioral [1050, 77, 135]. Behaviorial [455]. Behaviour [1295, 1181]. Benjamin [50, 51]. Berkeley [519, 399, 535]. Bertrand [1167, 1332]. Best [37, 808, 288]. Beta [539, 1369, 323]. Beta-Binomial [539, 1369]. Beta-Function [323]. Between [551, 438, 730, 987, 1161, 758, 644, 1320, 1038, 190, 195, 786, 659, 1045, 804, 809, 1192, 87, 238, 423, 1212, 119, 1378, 576, 959, 1233, 976, 137, 294, 860]. between-Group [1192]. Bevan [265]. Bhat [500]. Bias [609, 1312, 1321, 387, 1052, 580, 581, 471, 531, 955, 956, 957, 958]. biased [1217]. BIB [36]. Bibby [1230]. BIBD [593]. Bibliography [533, 1054, 1112, 905, 381, 1195]. Bickel [445]. Bijnen [761]. Binary [636, 1202, 944, 857]. Binomal [804]. Binomial [1258, 1159, 49, 653, 207, 394, 539, 1203, 104, 567, 1369, 1085, 1388, 286, 1257, 68, 397, 930]. Bio [159, 3]. Bio-Assays [3, 159]. Bioassay [483, 729, 631, 562, 442, 1251, 1174]. Bioassays [356, 599, 1255]. Bioavailability [693, 714, 1411, 800]. Biochemistry [1310]. Biocomputational [1399]. Bioequivalence [977]. Biography [303]. Biological [632, 1027, 780, 197, 66, 665, 674, 1066, 115, 266, 270, 1396]. Biologists [1262, 847]. Biology [308, 315, 1295, 1154, 663, 798, 1200, 421, 845, 308, 313, 317, 1271, 899, 1176, 799, 1338, 411]. Biomathematics [1027, 1397]. Biometric [884, 1162, 369, 834, 839, 15, 16, 17, 156, 157, 158, 325, 326, 327, 328, 478, 479, 480, 481, 605, 606, 607, 724, 725, 726, 727, 864, 865, 866, 986, 1126, 1127, 1275, 1272, 1273, 1274, 630, 362, 1048, 107, 275]. Biometrical [708, 301, 509, 1190]. Biometrics [885, 1160, 1045, 1060, 928, 1339, 1064, 1197, 935, 834, 1380, 955, 956, 1225, 278, 1093, 1229, 1240, 1241, 973, 1412, 985, 495, 496, 625, 626, 627, 628, 741, 742, 743, 744, 878, 879, 880, 1004, 1005, 1006, 1007, 1143, 1144, 1145, 1146, 1290, 1291, 1292, 655]. Biometrika [647]. Biometrischen [386]. Biometry [1306, 268, 510, 365]. Biometry. [66]. Biostatistical [226]. Biostatistics [1215, 712, 1313]. Biosynthesis [109]. Birds [1078]. Birth [450, 282, 133, 1329, 965, 684]. birth-death [1329, 965]. Bittinger [310]. Bivariate [792, 660, 1092, 1110, 472]. Blackith [549]. Blaxter [432]. Blending [1163]. Bliss [1307]. Block [159, 162, 989, 36, 183, 643, 3, 204, 913, 419, 593, 714, 932, 960]. Blocking [305]. Blocks [71, 209, 821, 568]. Blood [784]. Blowflies [178]. Blum [347]. Bodies [131]. Body [229, 285]. Bonferroni [843]. Book [1259, 1260, 1261, 1262, 149, 153, 154, 155, 301, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 319, 302, 320, 321, 322, 341, 323, 324, 477, 723, 1121, 1122, 1123, 1124, 1125, 1268, 1279, 1269, 1270, 1271, 1293, 176, 177, 1150, 1295, 347, 348, 498, 499, 500, 634, 635, 747, 883, 884, 1012, 1299, 1300, 1154, 1301, 503, 504, 755, 355, 1157, 184, 1309, 1310, 1311, 1313, 50, 51, 361, 1316, 509, 510, 511, 513, 514, 515, 647, 648, 649, 650, 651, 761, 762, 763, 764, 765, 766, 767, 768]. Book [769, 770, 771, 772, 189, 366, 516, 517, 518, 519, 520, 652, 773, 774, 775, 776, 898, 899, 900, 901, 902, 903, 904, 1024, 1025, 1026, 1027, 1028, 1029, 1030, 1031, 1170, 1171, 61, 62, 63, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 369, 370, 371, 372, 373, 374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384, 385, 386, 64, 65, 388, 389, 390, 66, 203, 1174, 1326, 526, 1176, 205, 1179, 789, 529, 530, 1047, 1328, 399, 533, 534, 535, 536, 537, 661, 662, 663, 793, 794, 916, 917, 918, 1050, 1051, 665, 666, 667, 797, 798, 799, 920, 921, 922]. Book [1054, 1055, 1056, 1187, 76, 77, 78, 79, 217, 1190, 1191, 219, 546, 674, 220, 221, 408, 1338, 548, 411, 549, 82, 83, 84, 85, 553, 1195, 89, 1341, 1342, 1343, 230, 1200, 1345, 416, 684, 1352, 241, 421, 422, 558, 242, 1206, 689, 1356, 110, 434, 435, 566, 1359, 113, 114, 115, 1210, 1361, 365, 512, 1211, 1362, 1363, 117, 1215, 1367, 1370, 1372, 1086, 1376, 265, 266, 445, 840, 698, 699, 268, 270, 447, 1224, 1386, 279, 280, 449, 586, 704, 845, 846, 847, 848, 1230, 962, 963, 1231, 1232, 1393, 1396, 1234, 1397, 455, 588, 706]. Book [849, 850, 707, 457, 458, 1244, 462, 712, 1404, 1405, 1406, 463, 464, 135]. Books [12, 13, 14, 148, 150, 151, 152, 300, 318]. Bookstein [1396]. Bossert [313]. both [641, 672]. Boullion [279]. Boundaries [215]. Box [661, 1370]. Boyce [520]. Bradford [934]. Bradley [885, 1032, 1074, 963]. Branching [792, 1013, 676]. Breast [1220]. Breeding [925, 447]. Bremner [537]. British [197, 561, 978]. Brodeau [1171]. Bruce [270]. Brunk [537]. Bryan [1023, 1090]. Buckland [373]. Buffon [248, 429]. Bühlman [464]. Building [548, 1370]. Bunker [195]. Burdette [314]. Bureaux [900]. Burns [315]. Business [789].

C [309, 319, 1269, 884, 1012, 1300, 44, 766, 189, 775, 898, 900, 903, 1027, 192, 202, 369, 382, 205, 789, 661, 82, 84, 1195, 1345, 684, 241, 558, 566, 115, 840, 847, 353]. Cain [704]. Calcul [425, 570]. Calculating [1258, 1159, 570]. Calculation [781, 69, 425, 671]. Calculations [382]. Calculus [845, 205, 1206]. Calibration [63, 1239, 1403]. Calibration-Statistical [63]. Cam [399]. Cameron [388]. Campbell [847]. Canadian [127]. Cancer [770, 387, 86, 96, 1220, 1242, 978]. Cancers [234]. Candidates [370]. Canonical [507, 453]. Capture [1262, 360, 243, 560, 471, 531]. Capture-Recapture [1262, 360, 560, 471]. Carboxyhemoglobin [1354]. carcass [358]. Carcinogenesis [1315, 575]. Carcinogenic [1060].

Carcinogenicity [1022]. Carcinomas [387]. Care [515, 1024]. caries [1382]. Carl [789, 303]. Carlo [181, 637, 820]. Carlson [635]. Carol [1195]. Carroll [1299]. cas [695]. Case [790, 1193, 829, 695, 696, 851, 987, 1317, 659, 1045]. Case-Control [1193, 696, 851, 1317]. Casebook [1051]. Cases [155, 539, 557]. Caste [1405]. Catchability [185, 505]. Categorical [18, 608, 527, 393, 6, 235, 417, 818, 1067, 1072, 1404, 1073, 971]. Categorized [675]. category [890]. Catered [261]. Causal [1168]. Causality [1167, 1332]. Causes [154, 385, 257, 1243]. Cavalli [277]. Cell [1149, 1164, 415, 428, 1397, 137]. Cells [86]. Cellular [716, 1295, 223]. Censored [330, 1278, 638, 188, 802, 937, 1213, 1378, 590, 592, 1339, 1340, 1248]. Censored-Survival [937]. Censoring [1350, 1080]. censorship [1327]. Censuses [1226, 1, 1225]. Centre [199]. Century [652]. Cerebello [858]. Cerebello-Thalamic [858]. Cerebral [858]. Cerebrov [424]. Certain [1334, 1203, 685]. Chain [653, 207, 394, 589]. Chain-Binomial [653, 394]. Chains [959, 1401]. Change [614, 2, 90, 418, 250, 1396]. Change-Over [614, 2, 418, 250]. Changes [282]. Chapman [1225, 471]. Character [87]. Characteristic [372]. Characteristics [375]. Characterization [446]. Characters [456]. Charles [689]. Chatterjee [1362]. Chaudhary [1190]. Chemistry [1271]. Chemosterilants [678]. Chemotherapeutic [595]. Chester [1307]. Chi [203, 831, 838]. Chi-Square [831, 838]. Chi-Squared [203]. Chiang [1300, 56]. Childhood [130]. Children [804, 576, 1236]. Choice [136, 932]. Chris [348]. Christiansen [1191]. chromiques [773]. chromiques-statisques [773]. Chromosomes [610, 54, 508, 923, 1333, 379]. cibernetice [149]. Cigarette [1354, 978]. Circadian [969, 1107]. Circular [1152]. Clarification [1208]. Class [754, 215, 678, 672, 101, 951, 859]. Classical [383]. Classification [610, 54, 508, 532, 1052, 538, 1053, 664, 1333, 822, 431, 433, 578, 272, 288, 1264, 451, 465, 861, 981]. Classifications [441, 669]. Classified [541, 641, 894]. classifying [345]. Claude [773]. Cliff [650]. Clines [1154]. Clinic [1410]. Clinical [314, 614, 871, 1418, 1015, 1036, 524, 912, 2, 1188, 1336, 552, 1201, 1068, 1351, 1366, 1216, 837, 1374, 1379, 1389, 1096, 1392, 1235, 1105, 1075, 1215]. Clinician [351]. Clinton [1195]. Clopper [855]. Clopper-Pearson [855]. Closed [1077]. Clow [845]. Clumped [1369]. Cluster [728, 1324, 249, 252, 277, 701, 1398, 1114, 761]. Clustered [1369]. Clustering [233, 816, 820, 688, 276, 557, 849]. Coal [561]. Cochran [746, 366, 1054, 1224, 1099]. Cody [920]. Coefficient [754, 1034, 210]. Coefficients [21, 611, 58, 781, 570, 836, 4, 108, 371]. Cohen [408, 1372, 135]. Cohort [1196]. Coincidence [523]. Colgan [1301]. Collapsing [180]. Collected [1107]. Collection [1376]. Colleges [197]. Colonies [1046]. Colony [428]. Column [783, 264]. Combination [729, 631, 1113]. Combinations [140, 654, 217, 924]. Combined [1255]. Combining [870, 750, 890, 267]. Come [596]. Comment [118]. Comments [1043]. Commingled [941]. Committee [322]. Common [37, 1335].

Commonwealth [900]. Communicable [1166, 274, 703]. Communities [1295, 920]. Community [780]. Compact [823]. Comparability [260]. Comparative [1419, 1011, 637, 914, 528, 658, 1069, 1239, 714, 1411, 358, 755]. Compared [730, 294]. Comparing [1258, 1159, 897, 58, 1173, 1327, 245, 687, 258, 979, 1110, 1075]. Comparison [730, 990, 36, 748, 181, 1017, 896, 57, 58, 59, 1325, 547, 812, 933, 554, 1199, 237, 820, 1360, 1083, 571, 1085, 949, 1088, 1245, 1251, 469, 598, 470, 1257, 1115, 1116, 931, 1064, 1276]. Comparisons [484, 1152, 350, 42, 639, 779, 885, 888, 905, 811, 1074, 1094]. Compartment [868, 1242, 935, 936]. Compartmental [988, 251, 432, 1355, 1082, 910]. Competing [48, 56, 1059, 409, 232, 103, 257, 120, 1219, 1243, 1296, 788]. Competition [97, 564, 254, 1357]. Complete [643, 71, 821, 703, 593, 960]. completely [641]. Complex [1293, 776, 93, 818, 1253]. Component [33, 40, 896, 1163, 224, 259, 269, 842, 278, 55, 1184]. Components [485, 1129, 36, 637, 1169, 402, 542, 74, 556, 272, 123, 124, 452, 288, 134, 970, 1021, 671, 1073, 1227, 591, 1241]. Composite [953]. Composition [642, 358]. Compound [1256]. Compounded [527]. Comprehensive [1412, 1413]. Compteur [523]. Computation [645, 705, 718, 108, 84]. Computational [292]. Computations [1419, 658]. Computer [307, 477, 789, 530, 1183, 1363, 980, 860, 411]. Computing [1014, 1021, 925, 948]. Concentrations [52]. Concept [252]. Conception [684]. Concepts [63, 76, 704]. Conceptual [270]. Concerning [1022]. Concomitant [719, 668, 929]. Concordance [1414]. Concurrence [284]. Conditional [48, 528, 832]. Conditions [100, 467]. Conference [512, 1232, 884, 798]. Confernece [902]. Confidence [637, 1022, 65, 924, 1196, 1079, 832, 1383, 855, 977, 1255]. Confounded [118, 289]. Congeneric [1239]. Conglomerates [131]. Connected [379]. Consecutive [633]. Consent [1068]. Consequences [403]. Conservatism [760, 574]. constants [935, 936]. Constrained [1201]. Constraints [868, 538, 806]. Constructing [80]. Construction [887, 381, 852]. Consultation [226]. Consulting [1177, 1112]. Consumption [229]. Contagion [696]. Contiguous [691]. Contingency [717, 161, 330, 482, 1298, 180, 188, 392, 1331, 405, 675, 102, 262, 838, 291, 1264, 641, 894, 465, 976]. Continuity [245, 439]. Continuous [20, 60, 534, 575, 851, 644, 1401]. Continuous-Carcinogenesis [575]. continuous-time [1401]. Contrast [860]. Contrasting [1095]. Contrasts [1019, 843, 1094]. Contributions [1268, 916]. Control [730, 38, 506, 762, 1166, 194, 381, 1193, 244, 571, 696, 274, 703, 851, 854, 294, 470, 1317]. Controlled [871, 912, 837]. Controls [851]. Convergence [85]. Convolution [456]. Cook [597]. Cooperative [1310]. Copernicus [764]. Cormack [320]. Cornish [600]. Coronary [1415]. Correction [159, 160, 161, 162, 163, 482, 483, 484, 485, 870, 1276, 1277, 1278, 1417, 1418, 1420, 37, 1339, 420, 245]. Corrections [18, 19, 20, 21, 329, 330, 331, 332, 608, 609, 610, 611, 612, 613, 614, 615, 728, 729, 730, 867, 868, 869, 871, 987, 988, 989, 990, 991, 1128, 1129, 1130, 1131, 1417, 1419, 1421, 1160, 1045, 1064, 935, 439, 955, 956, 1093, 1229, 1240, 1241, 973]. Correlated [357, 90, 1203, 822, 97, 972, 926, 1061, 951, 960, 472]. Correlation [21, 611, 507, 371, 401, 804, 238, 836, 576, 4, 644, 672, 1347]. correlations [1385]. Corsten [884]. Cortical [1172, 858]. Cost [861]. Coulter [523]. Council [377]. Counter [523]. Counts [1161, 1209, 1383]. Couple [128]. Course [113, 268]. Covariables [1400]. Covariance [162, 1129, 638, 368, 204, 209, 74, 1189, 1192, 1227, 591, 1241, 976, 859]. Covariances [758, 659, 1045, 551, 1104, 705, 987]. Covariates [1377, 1296]. Coverage [780]. Cowden [242]. Cox [1263, 1136, 1141, 1285]. Criteria [276]. Criterion [283, 943]. Critical [184, 425]. Critique [425, 788]. Crop [346]. Crop-Rotation [346]. Cross [1260, 1277, 541, 5, 932, 1379, 641]. Cross-Classified [541, 641]. Cross-Over [1379]. Cross-Tabulated [1260]. Cross-validatory [932]. Crosses [1018, 423]. Crossover [1418, 1205, 1105]. Croxton [242]. Cubic [813]. Culling [1129, 591, 1241]. Cultures [137]. Curve [1314, 67, 540, 1076, 1387, 1327]. Curves [19, 502, 906, 7, 829, 946, 1116, 1075, 371]. Cushing [1231]. Cutsem [1171]. Cycle [228]. Cyclic [388].

D [315, 1124, 1271, 355, 184, 1310, 650, 517, 901, 903, 1027, 195, 198, 378, 530, 537, 662, 1190, 674, 220, 1338, 1342, 416, 242, 1356, 1367, 445, 1225, 845, 1397]. D’Action [438]. Dagnelie [89]. Dam [1250]. Damped [656]. Daniel [566, 1313, 898, 1234]. Dans [523, 695]. Darroch [381]. Data [143, 296, 18, 332, 608, 1278, 882, 636, 638, 43, 639, 640, 754, 1158, 756, 760, 1022, 1025, 907, 779, 1037, 1038, 1168, 1169, 1043, 784, 785, 1175, 1178, 527, 787, 206, 393, 398, 6, 1183, 216, 1334, 675, 409, 1062, 1341, 90, 681, 1197, 234, 235, 417, 818, 1067, 1202, 237, 937, 1350, 1072, 1205, 243, 426, 560, 561, 251, 253, 691, 256, 1358, 1209, 569, 1364, 1365, 1213, 944, 1369, 1370, 574, 1373, 1086, 1220, 1378, 950, 272, 124, 961, 129, 1235, 286, 288, 1246, 1247, 1249, 851, 1410, 978, 1110, 1011, 891, 892, 641, 894]. data [1327, 1339, 1340, 1198, 1073, 1385, 1101, 1102, 1248, 971, 1254, 1341, 1102, 1260, 536, 793, 1352, 566, 1404, 762, 120]. Date [1237]. Daten [971]. David [388, 312, 1268]. Davies [411]. Dawkins [382]. Death [385, 396, 450, 1237, 1243, 133, 1329, 965, 154]. Deaths [155, 168]. Decision [37, 49, 391, 400, 963, 555, 1024]. Definite [1192]. Delay [1231]. Delayed [49]. Demographic [771, 377, 833, 50, 516]. Demography [1150, 707]. Dendritic [1080]. Densités [438]. Densities [438, 397, 825]. Density [752, 906, 395, 1181, 1368, 263, 1218, 139, 895, 683]. density-dependent [895]. dental [1382]. Department [197, 199]. Dependence [751, 752, 1059, 246]. Dependency [432]. Dependent [1042, 817, 836, 282, 345, 891, 895, 910, 676]. Derivation [507]. Derived [191]. Déscendances [695]. Descendant [695]. Descendants [695]. Describing [1092, 601]. Description [1244, 557]. Descriptive [179]. Design [162, 305, 614, 1418, 747, 1312, 1036, 374, 204, 1177, 2, 77, 81, 218, 88, 228, 418, 419, 823, 826, 261, 1370, 966, 1240, 592, 1105, 980, 860, 1116, 349, 160, 1128, 776, 1012]. Designed [8, 590]. Designing [1419, 1044, 658, 1389]. Designs [719, 159, 989, 749, 183, 643, 3, 524, 782, 783, 913, 914, 209, 80, 545, 807, 1205, 564, 250, 568, 118, 264, 1392, 1233, 711, 593, 289, 853, 714, 1108, 1109, 932, 960, 472, 776, 776, 388]. Desirable [183]. Detect [1371]. Detecting [791, 1410, 867]. Detection [332, 1163, 1035, 398, 1330, 803, 87, 90, 238, 841, 964, 1233, 713, 1106, 1417]. Determination [346, 528, 286, 709]. Determinations [1069]. Determine [1278, 137, 1339, 1340]. determined [895]. Determining [1173]. Deterministic [119, 133]. Develop [8]. Development [1295, 763]. Developments [1171, 1279]. Diagnosis [391, 833, 1091]. Diagnostic [484, 350, 907, 1195, 915, 1382]. Diagonal [948]. DIALL [307]. Diallel [307, 1277, 5, 423, 1252]. Diameters [95, 420, 1092]. Diamond [920]. dichotomiques [565]. Dichotomous [1040, 1334, 431, 851, 1416, 565]. Dictionary [373]. D’Identité [570]. Difference [1303, 65]. Differences [730, 1015, 1320, 281, 294]. Different [1043, 1180, 547, 100, 250, 572, 260, 1085, 949, 839, 1087, 1116, 1100]. Differential [784, 1338]. Differential-Equation [1338]. Differentials [354, 753]. diffusion [825, 1176]. Dillon [1261, 371]. Dilution [906, 818, 829]. Dimensional [296, 1298, 681, 641]. Dimensions [681]. dinamice [149]. D’Intervalles [438]. Diploid [810]. Diploids [758]. Direct [140, 867, 791, 1226, 134, 970, 700]. Direction [407]. Directional [354, 536]. Directions [248]. Disadvantageous [132]. Disadvantages [236]. Disagreements [351]. Disconnected [1169]. Disconnectedness [842]. Discovery [1027]. Discrepancy [1378]. Discrete [1125, 1062, 94, 439, 121, 1256, 1013, 644, 555, 1261, 499]. Discriminant [1261, 923, 1349, 290, 1104, 1414, 942, 690, 832, 943, 1047]. Discriminating [290]. Discrimination [352, 137]. Discs [694]. Discussion [885]. Disease [331, 41, 179, 364, 646, 1166, 396, 404, 539, 231, 1358, 1212, 274, 1091, 703, 964, 1247, 713, 1415, 1200]. Diseased [292]. Diseases [155, 696]. Disequilibrium [987, 1408, 659, 1045]. Disinfectant [384]. Dispersion [1346, 1371]. Distance [908, 237, 94, 430, 1360, 1368, 263, 809]. Distances [190]. Distributed [57, 554, 245, 263, 133, 598]. Distribution [1278, 348, 640, 356, 65, 660, 539, 543, 802, 215, 804, 91, 415, 104, 687, 429, 694, 945, 1373, 1223, 121, 1388, 1092, 128, 131, 855, 468, 349, 68, 676, 1339, 1340, 582, 466, 203]. Distribution-Free [356, 65, 1373]. Distributions [1258, 329, 35, 182, 1159, 907, 522, 1057, 547, 1194, 941, 430, 830, 833, 567, 1083, 575, 439, 1256, 1217, 534, 534]. Divergent [406]. Diversity [1199, 1344, 1098]. Divisibility [316]. Division [377]. Do [1079, 596]. Doctorate [375]. Dodd [384]. Doll [978]. Domains [1380, 1381]. Dombal [1024]. Dominance [138]. Donald [1345]. Doolittle [69]. Dosage [829]. Dose [19, 730, 363, 1022, 1315, 657, 787, 7, 562, 946, 273, 1095, 294, 470]. Dose-Response [787]. Doses [1060]. Double [358, 427, 286, 31]. double-stage [31]. Douglas [1012, 1299]. Doves [395]. Down [187, 668, 596]. Dr [1142]. Drawing [43]. Driessche [798]. Drift [45, 46, 224]. Drug [413, 583, 285, 595]. Drugs [141, 632, 568, 271]. due [551]. Dummy [211]. Dunnett [1094]. Dutch [763]. Dyn [799]. Dynamical [421]. Dynamics [45, 46, 359, 665, 1406, 1415, 83, 1231, 463].

Early [1336]. Easy [80]. Eberhardt [1054]. Ecological [1175, 682, 1246]. Ecology [317, 1322, 558, 1376, 1405, 768, 189, 921, 1054, 241, 920]. Econometrics [917]. Economic [346, 375, 134, 970]. Economics [548]. Editing [762]. Editor [992, 1280, 1305]. Editorial [1207]. Edmund [600]. Education [197]. Educational [375]. Edward [494, 922]. Edwards [518, 277]. Effect [1417, 778, 1175, 1330, 1091, 1237, 1246]. Effects [1122, 1010, 40, 180, 505, 45, 46, 1325, 413, 88, 1348, 1210, 440, 125, 285, 708, 711, 592, 1250, 1032, 924]. Efficiency [305, 218, 940, 127, 1233, 851]. Efficient [749, 1204, 1108, 1254]. Egg [255, 1209]. Egg-Counts [1209]. Egyptian [1209]. Eighth [902]. Eilon [194]. Einführung [220]. Electrical [1172]. Electroencephalogram [413]. Elementare [220]. Elementary [382, 1345, 434, 193, 704, 588]. Elements [948, 895, 500, 365]. Elliptical [378, 131]. Elston-Stewart [247]. Elzey [117]. Emphasis [1179]. emphasizing [1094]. Empirical [833]. Endler [1154]. Engen [1343]. Engineering [1029]. Engineers [1123]. English [846]. English/English [846]. Enquiries [771]. Entre [438]. Entropy [1342, 1201]. Enumeration [559, 1361]. Environment [317]. Environmental [1354]. Environments [1042]. Enzyme [109, 137]. Epidemic [886, 1297, 207, 394]. Epidemics [653, 1013]. Epidemiologic [1039]. epidemiological [101]. Epistasis [987, 659, 1045]. Epistatic [708]. Epistemological [1197, 1198]. Equal [162, 185, 204, 955, 956, 957, 958]. Equality [836]. Equally [225]. Equation [499, 1338, 229]. Equations [1017, 507, 812, 566, 1088, 1231, 1227]. Equilibrium [1310, 1041]. Equivalence [1038, 800, 100, 827, 693, 972, 1317]. Equivalent [1095]. Erfassung [1328]. Ergebnisse [293]. Ernest [1051]. Erratum [1280, 1225]. Error [499, 748, 181, 690, 1382]. Errors [8, 1155, 70, 73, 1091, 1410]. especially [971]. Essays [308, 317]. Establish [1038]. Establishments [153]. Estabrook [902]. Estimate [160, 991, 1128, 81, 844, 709, 471, 1093, 966, 1240]. Estimated [21, 485, 611, 52, 402, 949, 4, 1255, 471, 345]. Estimates [332, 754, 505, 360, 62, 70, 398, 73, 547, 1070, 445, 125, 1237, 890, 1021]. Estimating [882, 1158, 1312, 363, 646, 1040, 218, 676, 88, 234, 935, 936, 426, 262, 131, 286, 397, 671, 690, 161, 868]. Estimation [142, 718, 1129, 633, 1010, 499, 886, 1013, 749, 187, 188, 1019, 644, 1029, 521, 1034, 522, 1319, 1039, 1321, 61, 62, 371, 787, 1180, 395, 1182, 539, 214, 215, 74, 808, 406, 407, 672, 222, 550, 1194, 87, 415, 556, 559, 828, 829, 431, 251, 253, 111, 1082, 259, 945, 1368, 120, 1380, 1381, 1221, 267, 577, 839, 952, 1087, 1223, 842, 278, 281, 1230, 129, 706, 459, 967, 1242, 969, 288, 596, 974, 599, 1256, 139, 31, 349, 1184, 932, 1382, 1227, 591, 1101, 1102, 1241, 330, 988, 1421, 1060, 769, 64]. Estimator [1304, 356, 542, 1318, 757, 1335]. Estimators [36, 896, 1181, 1189, 227, 1063, 99, 437, 1085, 263, 1218, 267, 269, 272, 124, 452, 919, 811, 683, 938, 587]. Ethology [1301]. Étude [523]. Euclidean [1053]. Evaluating [719, 1351, 1216, 915]. Evaluation [322, 52, 72, 1402, 861]. Even [945]. Everitt [1352]. Eviatar [921]. Evidence [696, 962]. Evolution [308, 920, 1316, 230]. Evolutionary [1055]. Exact [528, 1186, 711, 1255, 893, 587]. Examination [185, 1033, 370, 1091]. Examine [696]. Examining [1058]. Example [1059, 97, 1362]. examples [955, 957, 1051]. Exist [252]. Existence [853, 457]. expectations [705]. Expected [354, 92, 123, 359, 69, 46]. Experiment [187, 93, 818, 261, 1407, 672, 697]. Experimental [719, 1417, 747, 749, 776, 1330, 1205, 1233, 592, 747, 374, 77]. Experimentation [900, 801, 835, 898, 899]. Experimenters [1370]. Experiments [8, 21, 160, 307, 611, 613, 1128, 1277, 1421, 178, 346, 1155, 756, 1323, 914, 71, 1060, 1334, 81, 218, 406, 407, 673, 5, 222, 550, 228, 1067, 1203, 96, 823, 97, 685, 254, 1357, 1358, 1084, 575, 4, 590, 1247, 857, 349, 55, 966, 1240, 1254, 472, 776, 1012]. Expert [322]. Exploration [897]. Explorations [304]. Exploratory [1247, 1025]. Exponential [656, 67, 668, 569, 757, 1075]. Exponentials [140, 654]. Expressing [1178, 1364]. Expressions [1077]. Extended [643, 960, 593]. extending [1032]. Extension [185, 71, 1111, 1114, 1094]. Extensions [578]. Extraneous [817]. Eyring [1271].

F [306, 316, 723, 634, 763, 766, 518, 902, 1024, 1171, 192, 195, 202, 374, 378, 385, 388, 390, 66, 1191, 242, 566, 117, 1367, 1376, 445, 447, 1396, 706, 712, 1405]. Factor [1026, 1412, 1413]. Factor-Analysis [1412, 1413]. Factorial [615, 212, 88, 289, 55, 305]. Factorials [211]. factorielle [774]. Factors [871, 912, 250, 837]. Facts [707]. Failure [257, 1219]. False [1182]. familial [1385]. Families [238, 240, 695, 1236, 1217]. Familles [695]. Family [719, 646, 906, 377, 804, 572, 1218, 576, 1098]. Famine [763]. Favorable [1130, 45, 1016, 1160]. Feature [37]. Federal [306]. Federer [649]. Feinstein [1215]. Fenchel [1191]. Fertility [186, 377, 128]. Fertilized [456]. Fertilizer [719]. Fertilizing [1165]. Fetal [1250]. Fetus [1259]. Field [426, 900]. Financial [1275]. Findings [381]. Findley [1367]. Finetti [794]. Finite [1276, 45, 46, 805, 78, 123, 589, 1245, 359, 1329, 965, 1401]. Finney [517]. Firing [655]. First [633, 362, 1354, 268, 858]. Fish [803, 559]. Fisher [1122, 1186, 1280, 772, 200, 1326]. Fisher-Irwin [1186]. Fisheries [677, 1386]. Fisherman [968]. Fishery [287]. Fit [1319, 668, 1062, 237, 1248]. Fitness [70, 73]. Fits [544]. Fitting [140, 640, 1314, 1033, 654, 67, 1046, 215, 223, 1388, 1253, 566]. Fitzpatrick [903]. Five [362, 1305]. Five-Year [1305]. Fixation [1130, 45, 46, 1016, 1160, 359]. Fixed [814, 924, 956, 958, 1122]. Fleiss [513]. Flexible [1253]. Flieger [707]. Folded [104, 68]. Follow [748, 778]. Follow-Up [778]. Food [72, 1372]. Foods [1361]. Forderberg [303]. Forecasting [689]. Forest [512, 292]. Forests [263]. Foreword [343]. Form [1038, 859]. Formalism [1342]. Formation [939, 428]. Formats [382]. Forms [1183]. Formula [1159, 1049]. Formulation [499]. formulations [1185]. Forrest [195]. Forrester [83]. Foundations [355, 652, 546]. Four [989, 913, 1251]. Four-Parameter [1251]. Fraction [88]. fractional [55]. Framingham [1415]. Francis [98]. Free [356, 65, 1373]. Freeman [117]. French [523, 525, 425, 565, 570, 695, 438]. Frequencies [1019, 190, 810, 809, 230]. Frequency [1161, 1043, 793, 1091, 669]. Friedrich [303]. Front [22, 23, 24, 25, 164, 165, 166, 167, 333, 334, 335, 336, 486, 487, 488, 489, 616, 617, 618, 619, 731, 732, 733, 734, 735, 872, 873, 874, 993, 994, 995, 996, 1132, 1133, 1134, 1135, 1281, 1282, 1283, 1284]. Fruit [900, 209]. Full [758]. Full-Sib [758]. Function [39, 352, 1033, 214, 437, 943, 323]. Functions [179, 906, 527, 401, 1104, 931, 372]. Fundamentals [510]. Further [610, 508, 660, 404, 227, 1097, 978].

G [1260, 309, 315, 302, 1293, 361, 763, 189, 366, 520, 902, 1026, 1171, 373, 381, 390, 533, 661, 662, 1187, 217, 219, 548, 411, 1370, 266, 1224, 1225, 588, 1405]. G. [740]. gain [981]. Gambling [457]. Games [723]. Gamma [432, 1011]. gamma-modeled [1011]. Gap [841, 1106]. Gardner [389]. Gart [1379]. Gauss [303]. geeigneter [971]. Gehan [314]. Gelfand [217]. Gemignani [205]. Gene [190, 810, 230, 440, 710, 809]. General [210, 1067, 1350, 242, 446, 1400, 860, 909, 910, 307]. Generalisations [964]. Generalised [1057]. Generalization [869, 1304, 946, 841, 1106, 1318, 915, 1094, 973, 975]. Generalizations [660]. Generalized [175, 818, 236, 94, 827, 128, 279, 449]. Generated [593]. Generation [1183, 676]. generation-time [676]. Generations [594, 1102]. Genes [551]. Genetic [21, 611, 749, 218, 673, 1190, 1192, 412, 1360, 1375, 4, 585, 456, 967, 134, 970, 1250, 1302, 1021, 809, 926, 1061, 692, 270]. Genetics [1316, 921, 230, 447, 708, 1100, 301, 504, 1170, 376, 797, 1056, 1232, 308]. Genetik [1393]. Genotype [1212]. Genotypes [1042]. Geographic [1154]. Geological [1026]. Geometric [1346]. George [302, 380, 553, 624, 759, 795, 796]. Georges [448]. Georgia [557]. German [465, 971]. Gertrude [1263, 1136, 1141, 1285]. Gerver [217]. Gibbons [416]. Giere [652]. Gilbert [509]. Giulio [1002]. Given [544, 412, 282]. Glenn [962]. GLIM [1309]. Glimpses [759]. Global [463]. Glossary [846]. Gnanadesikan [1086]. Goal [828]. Godambe [355]. Goel [799]. Golden [1376]. Goldstein [1261]. Gomez [1179, 1179]. Goodman [1404]. Goodness [1319, 668, 1062, 1248]. Goodness-of-Fit [1319, 1062]. Gorman [566]. Gradient [175]. Grandmaternal [970]. graph [108]. Graphical [181, 897, 1173, 1352]. Graphs [1258]. Grassland [1155]. Grazing [1323]. Green [1299, 246]. Gremy [1024]. Grid [691]. Group [642, 524, 784, 1192, 828, 568, 592, 930]. Grouped [787, 1373, 120, 1220, 843, 852, 1103]. Grouping [728, 701, 1353]. Groups [423, 1116]. Growth [19, 502, 1017, 511, 769, 1033, 656, 1046, 7, 86, 812, 826, 1076, 1387, 1395, 1242, 467, 1116, 718, 349, 895]. Grubb [227]. Grubbs [99, 587]. Grundlagen [1393]. Grundmann [770]. Guiasu [149]. Guide [390, 1359]. Guidelines [1166, 1214, 274, 703, 1409]. Gupta [533, 662]. Guttman [383].

H [307, 309, 312, 313, 1123, 1268, 1271, 634, 511, 647, 765, 766, 776, 61, 62, 63, 195, 198, 382, 384, 388, 390, 203, 533, 537, 666, 219, 82, 1200, 445, 1386, 1230, 464]. Haberman [793]. Haemopoietic [428]. Haenszel [1317, 1335, 1083]. Half [1252]. Hall [303]. Halothane [195]. Handling [133]. Hanpel [445]. Haploid [594, 359]. Hardy [893, 1041]. Harmon [375]. Harold [110]. Harris [850]. Harrison [520]. Harter [198, 61, 62]. Hartigan [849]. Hartley [647, 1268]. Harvest [127]. Harvesting [501, 777]. Hassell [1406]. Having [209, 123, 1264, 859]. Hayne [1318, 1304]. Hayne-Type [1318]. Hazard [1377, 1400]. Health [153, 154, 155, 173, 322, 515, 191, 201, 64, 674, 1062, 680, 1348, 974, 686, 1211, 51]. Healthy [292]. Heart [1415]. Heights [1092]. Heim [1293]. Hemophilia [521]. Henderson [1271]. Henry [1003, 239]. Hepatic [195]. Hepatitis [521]. Heritabilities [406, 407]. Heritability [160, 485, 1128, 1421, 1040, 402, 81, 222, 672, 966, 1101, 1102, 1240]. Heritage [764]. Hermite [945]. Heron [1365]. Herrnstein [667]. Hess [903]. Heterogeneity [1035]. Heterogeneous [639, 1407, 1254]. Heteroscedastic [550, 577]. Hexagonal [378]. Hierarchical [1071]. High [296]. High-Dimensional [296]. Higher [833]. Higher-Order [833]. Highway [381]. Hill [934, 978]. Hiorns [665]. Histories [646]. History [901]. Hit [142, 679]. Hodges [76]. Home [1037]. Homogeneity [357, 240, 827, 433, 451]. homogeneous [1329, 965, 1401]. Honor [1268, 380, 553]. Hospital [153]. Hospitals [903]. Hotelling [414]. Household [539, 964, 713]. Households [1180, 1087]. Hsiung [1056]. Huber [445]. Huizinga [526]. Hultquist [114]. Human [610, 186, 54, 508, 511, 520, 923, 1333, 238, 240, 1236, 1217, 1100, 1055]. humans [1302]. Hunger [763]. Hunter [1370, 1370]. Huzurbazar [1030]. Hybrid [410]. Hybridization [453]. Hypergeometric [567]. Hypotheses [1022, 972, 1317, 587]. Hypothesis [427].

I. [667]. Ibem [774]. ICMSF [1361]. Ideas [444]. Identification [379, 923]. identité [570]. Identities [710]. identity [570]. II [155, 308, 1295, 197, 386, 535, 956, 867, 46, 791, 807, 1336, 1061, 232, 103, 958, 1102, 592, 341]. II. [168, 407]. III [153]. Illustrated [833]. Illustration [72, 444]. Impact [1350]. Impairment [1062]. Implications [364, 926, 1061]. Importance [545, 807]. Improved [1258, 1130, 1159, 542, 815, 1070, 1230, 1016, 1160, 757]. Inbred [1421, 222]. Inbreeding [758, 781, 108]. Incidence [574, 1242, 978, 1382]. Incidences [41]. Included [423]. Including [829, 696]. Incomplete [323, 482, 36, 3, 778, 392, 913, 216, 417, 419, 821, 102, 568, 714, 932, 971, 159, 989]. incompletely [894]. incorporate [1032]. Incorporating [209]. Incorrectly [269]. Increasing [1095]. Independence [1331, 405, 959, 136, 634]. Independent [1258, 717, 608, 870, 1420, 750, 353, 1173, 393, 918, 267, 843, 472]. Index [806, 425, 1371, 129, 134]. Indexes [53, 425]. Indian [904]. Indicator [751, 906]. Indicator-Dilution [906]. Indicators [322, 1354, 252]. Indirectly [109]. Individually [111]. Individuals [1421, 540, 222]. Induced [201]. Industrial [898, 83]. Industry [548]. Inégale [695]. Inequality [585]. Infected [292]. Infectious [155, 130]. Inference [648, 537, 661, 840, 1011, 311, 755, 355, 772, 662, 416, 110]. Inferences [43, 1168, 679, 1208, 1360, 582, 1408, 892]. Inferred [1022]. Infinite [316, 1077]. Influence [72]. Influencing [87]. Informal [309]. Informatiei [149]. Information [141, 30, 174, 342, 497, 629, 745, 881, 1008, 1147, 1294, 352, 646, 1036, 395, 929, 412, 238, 971, 515]. Informed [1068]. Ing [1142]. Inheritance [364]. Initiated [234]. insbesondere [971]. Insects [1405]. Institute [370]. Institutions [1389]. integral [1384]. Integration [67]. Integrodifferential [1231]. Intensities [1059]. Intensity [116]. inter [932]. inter- [932]. Interacting [264]. Interaction [473, 330, 990, 352, 188, 933, 256, 119, 1064]. Interactions [1222]. interclass [1385]. Intermediate [176]. Internal [1390]. Internally [843]. International [884, 902, 271, 1232]. Interpolating [1049]. Interpretation [766, 1197, 974, 1407, 1198, 509]. Interrelationships [754]. Intersect [1318]. Intersecting [719]. Interval [1276, 829, 1245]. Intervals [637, 1022, 779, 65, 1196, 1079, 438, 1383, 855, 1107, 977, 924, 832]. Intervention [1175, 1246]. Interview [64]. intra [672, 859]. intra-class [672, 859]. intrablock [932]. Introduction [309, 1270, 348, 883, 765, 517, 530, 78, 241, 117, 1370, 698, 997, 310, 699, 1269, 767, 529, 422, 110, 434, 114, 265, 712]. Intuitive [219]. invariance [671]. Invariant [938, 919, 451]. Inventory [194]. Inverse [1, 397, 925, 948, 1225, 279, 449, 468, 700]. Investigating [1262]. Investigation [187, 1242]. Investigations [1068]. Invited [655, 1060, 928, 1197, 834, 1380, 278, 1412]. Involving [719, 54, 72, 569, 595, 1250, 857, 472]. Iosifescu [663]. Irradiation [1149, 396]. Irreversible [633]. Irving [848]. Irwin [1186]. Isonymy [224]. isotonic [670]. Issues [455]. Iterative [208, 547, 1021]. Ittner [1307]. IV [1028].

J [1259, 1260, 301, 314, 341, 1122, 177, 347, 1299, 1154, 1309, 511, 513, 514, 515, 650, 761, 762, 764, 768, 769, 516, 517, 519, 520, 774, 775, 776, 899, 1026, 1029, 1031, 1171, 192, 195, 202, 371, 381, 388, 390, 526, 529, 399, 535, 537, 793, 916, 667, 1054, 76, 219, 408, 1338, 82, 1195, 416, 684, 242, 110, 566, 1210, 1211, 1370, 1372, 1376, 265, 445, 1386, 845, 1230, 1231, 1397, 849, 850, 457, 295]. Jackknife [33, 657, 273]. Jacob [135]. James [498, 66, 1359, 963, 712]. January [1275]. Jared [920]. Jay [83]. Jean [773]. Jean-Claude [773]. Jensen [1017, 1088]. Jerome [44]. Jerry [499]. Jewish [648]. Jinks [301].

John [651, 1025, 388, 84, 1200, 494]. Johnson [1125, 534, 323]. Joint [71, 979]. Jointly [632]. Jolly [505, 531, 243]. Jolly-Seber [505]. Joreskog [1026]. Jr [195, 76]. Jr. [1003, 514, 515, 389, 1232]. judges [786]. Judy [64]. Juvancz [386]. Juvenile [234].

Kabe [662]. Kappa [1071]. Kappa-type [1071]. Karel [79]. Karlin [921]. kategoriale [971]. Kay [304]. Keedy [310]. Kempthorne [1232]. Kendall [373, 302, 548]. Keyfitz [707, 1150, 516]. kidney [670]. Kilpatrick [515]. Kimball [120]. Kimura [504]. Kinds [394]. Kinetic [403]. Kinetics [1397]. kinship [108]. Kirillov [217]. Kirk [77, 455]. Klassifizierung [465]. Kleijnen [775]. Klein [242]. Kline [1270]. Klovan [1026]. Knott [1114]. Knowledge [377, 31]. known [1327]. Knox [473]. Kojima [376]. Kolmogoroff [1376, 1376]. Kolmogorov [1014, 915]. Konferenz [386]. Konstantinov [217]. Kontingenztafeln [465]. Koster [82]. Kostitzin [1376]. Kotz [1125, 534, 846]. Kramer [191]. Krishnaiah [1121, 1279, 1028]. Krosigk [447]. Krotki [769]. Kruskal [390]. Ku [63]. Kushnirenko [217].

L [301, 310, 311, 323, 324, 1122, 1125, 884, 1300, 510, 513, 514, 648, 519, 776, 1027, 61, 62, 195, 198, 371, 375, 389, 1174, 1176, 399, 534, 535, 920, 1054, 76, 217, 221, 82, 1341, 266, 698, 279, 1396, 1244, 1404]. Laan [65]. Laboratory [1174]. Lachenbruck [1047]. Ladell [371]. Lag [467]. Lagged [906]. Laloire [773]. Lampkin [194]. Lancaster [765, 203]. Language [789, 1183]. Lapse [771]. Large [948, 272, 124, 1335]. Larson [110]. Latency [1242]. Latent [234, 696]. Latin [887, 80, 545, 807]. Lattice [813, 1109]. Laurence [1003]. Law [656, 438, 893]. Leach [1269]. Least [307, 251, 1390, 1021]. Lebart [1244]. LeCam [519, 535]. Lecture [1311, 1027]. Lectures [1338]. Lee [747, 747]. Lehman [76]. Lengths [95, 420]. Leslie [718, 185, 895, 677]. Leukemia [130, 557]. Level [730, 562]. Levelled [853]. Levels [730, 250, 1095, 294, 470]. Levin [1295, 1295]. Levine [1342]. Lewis [503]. Li [44, 1056]. Liability [364]. Life [1280, 767, 385, 387, 1326, 1051, 929, 815, 422, 560, 1206, 1356, 257, 1327, 1353, 1103, 1300]. Life-Time [387]. Like [682]. Likelihood [988, 522, 787, 668, 539, 547, 415, 1082, 437, 949, 272, 276, 124, 460, 1248, 518]. Limits [543, 802, 217]. Lindeman [776]. Lindley [121]. Lindsey [309]. Line [1131, 1304, 1218, 702, 1394, 1251, 1318, 892, 683, 1229]. Line-Transect [1218]. Linear [140, 8, 18, 988, 1153, 40, 636, 1161, 654, 1325, 529, 208, 403, 6, 405, 808, 928, 550, 417, 253, 1082, 1358, 267, 839, 446, 1230, 284, 1239, 461, 1243, 290, 1249, 1253, 1257, 909, 669, 924, 927, 943, 435, 840]. Lines [719, 1276, 42, 423, 430, 1245]. Linkage [987, 87, 1233, 1408, 659, 1045]. Linked [551]. Literature [648]. Litters [1044]. Liver [415]. Load [270]. locally [68]. Locating [378]. Location [65, 445]. Loci [810]. Locomotion [1210]. Locus [1157, 87, 1233, 589, 594, 138]. Log [1153, 1019, 363, 1325, 208, 405, 669, 91, 1358, 1253, 889]. Log-Dose [363]. Log-Frequencies [1019]. Log-Linear [1153, 208, 1358, 1253, 669].

Log-Zero-Poisson [91]. Logarithmic [855]. Logic [360, 434]. Logical [1197, 1198]. Logistic [352, 1156, 1017, 802, 812, 561, 687, 569, 437, 947, 1088, 1251, 582]. Logistic-Exponential [569]. Logit [213, 946, 1317]. Lognormal [640]. Logrank [574]. Loi [438]. Long [1155, 1015, 696]. Long-Term [1155, 1015]. Longitudinal [1235, 1415]. Loss [46, 290, 359]. Lotka [1376, 718]. Lucas [1003]. luck [968]. Lukacs [372, 85, 699]. Lung [978]. Lymphoma [47].

M [1259, 1261, 1262, 1263, 305, 310, 311, 315, 302, 320, 1270, 1285, 499, 504, 763, 366, 191, 371, 373, 379, 388, 1176, 205, 537, 663, 1051, 665, 920, 1054, 217, 1191, 674, 548, 82, 1195, 1342, 1345, 684, 98, 1376, 265, 449, 1231, 1397, 458, 1406]. Maccacaro [1002]. MacDonald [315, 1397]. Machines [848]. MaClean [776]. Made [1152, 1051]. Mailloux [82]. Main [88]. Mainland [1345, 1345]. Maintenance [1341]. Majority [1071]. Making [1024, 1168, 1360]. Male [678]. Malignant [168, 396]. Mammary [387]. Man [1259, 526]. Management [789, 854]. Managerial [689]. MANOVA [444]. Mantel [1090, 1317, 1335, 1208, 1083]. Mantel-Bryan [1023, 1090]. Manual [347, 377, 797]. Many [989, 913]. Mardia [536]. Marginal [161, 988, 1298, 827, 1082, 262]. Mark [1341]. Mark-Recapture [1341]. Marker [87, 1233]. Markov [207, 959, 589, 1401, 601, 458]. Markovian [1222, 32, 1097]. Marks [769]. Marolla [763]. Marriott [766]. Martin [920]. Maruyama [1170]. Mary [1136, 1141]. Matched [1039, 1193, 831, 111, 258, 581]. Matching [1081, 580, 955, 956, 957, 958, 609]. Matérn [379]. Maternal [967, 134, 970]. Mathematic [1150]. Mathematical [304, 1124, 1295, 883, 892, 184, 506, 651, 519, 1027, 198, 199, 384, 399, 535, 665, 241, 244, 255, 434, 699, 962, 858, 1299, 511, 376, 798, 548, 684, 464]. Mathematics [723, 767, 422, 454, 79, 1386, 310, 1270, 1356, 1211, 845]. mathematische [220, 1393]. Mather [301, 365]. Mating [178, 642, 758, 810, 138]. Matras [516]. Matrices [897, 1192, 449, 1116, 279]. Matrix [869, 777, 925, 1189, 677, 948, 133, 852, 895, 973, 975]. Matter [22, 23, 24, 25, 164, 165, 166, 167, 333, 334, 335, 336, 486, 487, 488, 489, 616, 617, 618, 619, 731, 732, 733, 734, 735, 872, 873, 874, 993, 994, 995, 996, 1132, 1133, 1134, 1135, 1281, 1282, 1283, 1284]. Matthews [1356]. Maurer [530]. Maurice [196]. Max [447]. Maximal [664]. Maximally [919]. Maximally-invariant [919]. Maximum [522, 787, 539, 547, 1342, 415, 1201, 437, 949, 272, 124, 981]. Maximums [956, 958]. Maxwell [1050]. Maynard [768]. McDonald [1122]. McNemar [39, 827, 1379]. Mean [882, 1303, 1178, 552, 1364, 123, 709, 31, 69, 705]. Means [728, 730, 908, 1189, 1346, 701, 294]. Measure [236, 1360]. Measurement [63, 1344, 1067, 1072, 1396]. Measurements [1418, 1194, 424, 112, 436, 1239, 1105, 1111, 1416]. Measures [351, 1199, 1346, 237, 1212, 1398, 1098]. Measuring [786, 109, 557, 1191].

Medical [315, 321, 1311, 765, 1024, 1182, 801, 1183, 1345, 835, 1214, 586, 1237, 1409, 1195]. Medicine [663]. Medieval [648]. Mehrdimensionaler [465]. Memoriam [1285, 795, 796]. Mendel [499]. Mendelschen [293]. Menken [684]. Mental [191]. Meritocracy [667]. Messy [1249]. Metastases [939]. Method [140, 163, 728, 870, 1131, 175, 750, 1158, 187, 652, 1172, 67, 912, 71, 668, 80, 925, 225, 815, 96, 243, 249, 252, 1368, 573, 578, 277, 701, 126, 702, 1390, 1395, 1114, 889, 905, 1184, 683, 1229, 715, 671]. methodes [774, 89, 773]. Methodology [473, 1188, 1065, 1067, 834, 1399]. Methods [173, 321, 746, 43, 361, 513, 772, 516, 904, 908, 58, 200, 369, 207, 1333, 1190, 93, 418, 1348, 1360, 1361, 946, 959, 276, 1096, 1235, 707, 464, 292, 1115, 69, 955, 956, 957, 958, 176, 366, 196, 588, 1086]. Methyl [72]. Metrics [688]. metropolitan [557]. Mice [387]. Micro [1361]. Micro-Organisms [1361]. Microbial [459]. Migrating [559]. Migration [440, 133]. Milestone [882]. Miller [1195, 1345]. Milton [84, 319]. Miners [561]. Minimum [352, 1321, 544, 838]. MINQUE [1337]. misallocation [832]. Misclassification [1010]. Misclassifications [286]. Missing [5, 253, 256, 441, 1235, 410, 1277]. Misspecification [446]. Mist [1078]. Mitome [305]. Mixed [393, 74, 814, 235, 138, 890, 1185, 811, 1227, 700, 608]. Mixing [316, 1194]. Mixture [344, 547, 1194, 1407]. Mixtures [1163, 522, 830]. Mobility [375]. Model [868, 988, 990, 1148, 1149, 1151, 36, 886, 751, 506, 186, 1161, 52, 364, 653, 784, 1046, 400, 668, 540, 74, 808, 223, 548, 86, 228, 933, 90, 679, 232, 417, 1203, 937, 939, 1076, 685, 103, 428, 251, 432, 253, 255, 256, 1082, 569, 572, 1369, 119, 1370, 947, 1222, 952, 269, 446, 272, 282, 589, 1400, 1242, 133, 1250, 1251, 467, 601, 1264, 1032, 910, 531, 670, 1185, 811, 1064, 935, 936, 938, 1073, 1074, 1227]. modeled [1011]. Modeles [774]. Modeling [723]. Modelle [1328]. Modelling [1365]. Models [8, 719, 18, 608, 1125, 632, 1297, 1153, 891, 754, 47, 511, 1027, 777, 1322, 655, 1325, 393, 207, 394, 208, 799, 1056, 6, 405, 1187, 1338, 677, 1195, 814, 235, 1204, 684, 244, 431, 1355, 1358, 1375, 1377, 266, 839, 123, 1231, 1236, 1243, 1249, 1251, 1253, 137, 718, 1296, 1302, 892, 788, 669, 825, 1100, 976, 1343, 435, 1230, 768, 120, 280]. Moderate [1104]. Modern [310, 1125, 1263]. Modification [843]. Modified [1023, 1057, 818, 1090, 1256]. Modified- [818]. Moment [945, 1388]. Moments [283, 1329]. Monitoring [32, 1015, 1097]. Monographs [1154, 917]. Monotone [829]. Monte [181, 637, 820]. Montgomery [1012]. Morbidity [191]. Morineau [1244]. Morphometrics [1076, 549]. Mortality [168, 613, 1300, 409, 1063, 697]. most [68]. Mosteller [306, 192, 195, 202, 390]. Mourning [395]. Mouse [47]. Movements [803]. Moves [1394]. MR [1160, 1339]. Muelen [309]. Multi [482, 34, 392, 681, 1410, 825]. multi-allele [825]. Multi-Clinic [1410]. Multi-Dimensional [681]. Multi-Variate [34]. Multi-Way [482, 392]. Multiclinic [756]. Multicollinearities [1058]. Multicompartment [716, 266]. Multicomponent [1163]. Multidimensional [180, 565, 465]. multidimensionnel [565]. Multifactorial [1302, 364, 692, 1384].

Multinomial [532, 1052, 215, 237, 573, 460, 469, 101]. multinormal [951]. Multiple [716, 990, 37, 748, 181, 639, 766, 58, 59, 933, 1071, 822, 940, 100, 252, 1366, 1226, 583, 1094, 1391, 453, 710, 1243, 851, 1, 1064, 1225]. Multiple-Decision [37]. Multiple-Stage [583]. multiplicative [976]. Multiplicity [452]. Multiresponse [332, 398]. multitrait [1227]. Multivariate [143, 329, 175, 498, 1299, 636, 182, 533, 534, 213, 1193, 413, 414, 90, 1352, 105, 688, 431, 253, 1208, 1086, 841, 955, 956, 957, 958, 129, 1235, 850, 1106, 443, 1384, 466, 1028, 662, 1050, 549]. Multiway [41]. Murray [1338]. Must [214]. Mutants [132]. Mutation [544, 138]. Mutually [887]. Myers [1123]. Myslivec [1142].

N [323, 1125, 500, 509, 510, 648, 516, 652, 904, 534, 217, 1341, 113, 1211, 1376, 845, 1244, 80]. Nagylaki [1157]. Narendra [799]. Nassar [597]. Nathan [1150, 707]. National [195, 196]. Natural [1316, 1191]. Nature [1197, 1198]. Nature-Nurture [1197, 1198]. Naus [762]. near [657, 273]. Necrosis [195]. Needs [317]. Negative [804, 1085, 1388, 397]. Negatives [1182]. Nei [1360]. Nelder [1309, 84]. Nelson [689]. Neoplasms [168]. Nested [272, 930]. Nets [1078, 378]. Neue [293]. Neuron [655]. Neurons [1187]. neutral [825]. Nevo [921]. News [26, 27, 28, 29, 169, 170, 171, 172, 337, 338, 339, 340, 490, 491, 492, 493, 620, 621, 622, 623, 736, 737, 738, 739, 875, 876, 877, 998, 999, 1000, 1001, 1137, 1138, 1139, 1140, 1286, 1287, 1288, 1289]. Neyman [519, 399, 535, 764, 660]. Niche [1372]. Nineteenth [652]. Nira [799]. No [199, 885, 352, 1160, 1339, 1064, 935, 955, 956, 1093, 1240, 973, 467]. Non [305, 870, 1269, 1420, 750, 353, 1192, 93, 95, 420, 829, 248, 429, 260, 264, 441, 1095, 1114]. Non-Additivity [305, 441]. Non-Comparability [260]. Non-Independent [870, 1420, 750, 353]. Non-Interacting [264]. Non-Monotone [829]. Non-Parametric [93, 829, 1095, 1114, 1269]. Non-Positive [1192]. Non-Random [95, 420, 248, 429]. Nonadditivity [1164]. None [418]. None-Parametric [418]. Nonlinear [389, 1338]. Nonnormal [1204]. Nonparametric [329, 182, 105, 1208, 1213, 1368, 1234, 892, 311, 416]. Normal [319, 37, 753, 906, 522, 61, 543, 215, 547, 1194, 1347, 687, 830, 129, 592, 1384, 466]. Normality [824, 950]. Note [473, 992, 1280, 37, 41, 887, 54, 404, 81, 420, 831, 120, 838, 1088, 1226, 1228, 1109, 139, 910, 1401, 160, 161, 162, 1420, 178, 501, 39, 351, 40, 181, 42, 353, 356, 357, 363, 507, 56, 521, 57, 204, 70, 71, 401, 538, 211, 404, 213, 542, 215, 80, 545, 224, 225, 227, 551, 88, 228, 231, 92, 418, 236, 237, 240, 419, 423, 97, 100, 104, 246, 106, 248, 429, 430, 564, 250, 432, 255, 256, 567, 572, 118, 573, 262, 574, 264, 269, 446, 272, 121, 122, 4, 277, 584, 585, 452, 283, 284, 460]. Note [289, 597, 471, 291]. Notes [1311, 1027]. Nuclei [415]. Null [427, 1079]. Null-Hypothesis [427]. Number [142, 168, 484, 350, 1180, 539, 679, 1087]. Numbers [785, 912, 804, 576]. Numerator [925, 948]. Numerical [902, 67, 811, 666]. Nurture [1197, 1198]. Nutrients [719]. Nutrition [1044].

O [313, 1268, 647, 765, 203, 922, 1232, 1405, 463]. Obituary [494, 624, 740, 1002, 1003, 1141, 1142, 44, 239, 98, 448, 600, 295]. Observable [387]. Observation [1164]. Observational [609, 1168, 580, 581, 1410]. Observations [1128, 49, 57, 1323, 225, 554, 245, 590, 598, 345, 410, 960, 966, 1103, 1240, 1101, 766, 1086]. Observed [415, 1327]. Observer [1072]. Observers [1071]. Observing [564]. Obstetric [234]. Obtaining [1196]. Occasions [612, 584]. Occupancy [869, 680, 686, 973, 975]. odds [889, 1335]. Odeh [1124]. Odell [279]. Offspring [160, 1128, 81, 672, 966, 1101, 1240]. Ogawa [776]. Ohta [504]. Omitted [290]. One [870, 37, 750, 1157, 1164, 657, 214, 802, 407, 562, 273, 138, 915, 672, 938, 1073]. One-and [1157]. One-Sided [870, 750, 802, 915]. one-way [938, 1073]. Onset [1351, 1216]. Open [1023, 1090]. Optima [346]. Optimal [112, 954, 1391, 854]. Optimum [501, 1172, 106, 571, 116, 930]. Oral [173]. Ord [650]. Order [319, 633, 367, 62, 1325, 80, 926, 1061, 833, 951, 122, 1032, 312, 61, 62]. Ordered [469, 1264, 669]. Organisms [632, 1361]. Organization [173, 201]. Organs [214]. Oriented [477, 1363]. Origin [130]. Orthogonal [887, 545, 807]. Oster [1405]. Other [900]. Outlier [332, 398, 909]. Outliers [867, 791, 841, 1106]. Output [1107]. Overall [708]. Overlapping [594]. Owen [1124, 901, 198]. Oxygen [229].

P [304, 321, 477, 1121, 1279, 348, 1299, 1301, 503, 355, 189, 775, 1028, 192, 195, 202, 65, 533, 661, 662, 663, 1363, 1370, 445, 279, 1397, 1406]. Painting [96]. Pair [784, 1325, 1193, 1081, 831, 1032]. Pair-Matched [1193, 831]. Pair-Matching [1081]. Paired [1152, 1325, 885, 888, 905, 1074]. Pairs [1039]. Palm [1293]. Panel [1348]. Paper [655, 1060, 928, 1197, 834, 120, 1380, 278, 597, 1412]. Papers [302, 495, 496, 625, 626, 627, 628, 741, 742, 743, 744, 878, 879, 880, 1004, 1005, 1006, 1007, 1143, 1144, 1145, 1146, 1268, 1290, 1291, 1292, 199, 380, 553, 1214, 1409, 374]. Parabolas [599]. Parallel [1251]. Parallelism [284]. Parameter [499, 49, 1319, 559, 969, 1251, 757, 1009, 1029]. Parameters [988, 749, 751, 1181, 214, 218, 547, 550, 687, 1082, 1085, 949, 839, 952, 1223, 282, 582]. Parametric [41, 1304, 93, 418, 821, 829, 1095, 1114, 1269]. Parametrization [437]. Parent [695, 672, 695]. Parental [423]. Parents [160, 1128, 81, 966, 1101, 1240]. Park [1180, 1087]. Part [1295, 775]. Partial [1277, 371, 5, 129]. Partially [80, 675, 641]. Particle [139]. Partition [1042]. Parts [168]. Passage [1354]. patient [1254]. Patients [1374, 1389, 903]. Patil [189]. Patten [558]. Pattern [691, 1371, 137, 1113, 1115, 715, 1295]. Patterns [183, 908, 1319, 601]. Paul [1051]. Pauline [798]. Peanut [950]. Pearce [900]. Pearson [323, 855]. Peason [647]. Peck [82]. Pedersen [770]. Pedley [1210]. Pegels [789]. People [317]. percent [955, 956, 957, 958]. Perennial [900, 1323]. Perimeter [1009]. Period [614, 1418, 2, 418, 1105]. periodic [410]. Periodicities [1271]. Periods [234, 696]. permutational [101]. Personal [759]. Personnel [153]. Perspectives [1306]. Pest [38, 506, 244, 854]. Pesticides [1174, 1391]. Peter [1047, 917, 666, 280]. Petrie [1311]. Petrov [918]. Ph.D [375]. Pharmacokinetics [961]. Phase [1336, 467]. Phénomène [523]. Phenomenon [523]. Phenotypic [485, 1042, 402]. Phillips [514]. Phosphenes [1172]. Physical [1310]. Physicians [978]. Physics [315]. Physiological [184]. Pielou [189, 241]. Pieters [390]. pioneer [1263]. Pirchner [447]. Placed [694]. Plackett [698]. Plane [415, 95, 420, 131]. Planes [78]. Planning [322, 377, 314]. Plans [1336]. Plant [749, 1035, 95, 420, 254, 1368]. Plants [900, 1323]. Plateau [540]. Plot [93]. Plots [296, 681]. Plotting [1172]. Pocock [912]. Point [503, 908, 1319, 378, 430, 945]. Poisson [91, 1419, 349, 640, 57, 658, 554, 1205, 245, 567, 121, 468, 598]. Poisson-Distributed [57, 554, 245, 598]. Poland [1310]. Policy [32, 1097]. Pollak [1232]. Pollard [511]. Pollen [1383]. Pollution [651, 206, 803, 1348]. Polynomial [1321, 225, 1084, 1116, 410]. Pooling [1065]. POPAN [1341]. POPAN-2. [1341]. Popular [1081]. Population [313, 38, 501, 1154, 504, 1158, 1017, 777, 1170, 376, 377, 526, 400, 1056, 540, 805, 812, 1200, 96, 559, 573, 1231, 594, 852, 467, 471, 718, 890, 359, 895, 397, 683, 555, 805, 61, 769, 921, 447, 707]. Populations [344, 1295, 37, 749, 45, 46, 642, 1165, 1316, 1035, 190, 810, 230, 95, 420, 244, 431, 123, 1394, 1233, 456, 133, 469, 1110, 359, 1217, 700, 1254, 1316, 511, 520, 62, 665, 1191, 1359]. Positions [1172]. Positive [1192, 138]. Possible [195]. Post [521]. Postelnicu [884]. Postoperative [195]. Potency [363, 829, 599, 1255]. Potentials [438]. Potentiels [438]. Poultry [255]. Power [39, 658, 401, 1371, 290, 135, 1419]. powerful [68]. Practical [226, 249]. Practice [66, 377]. Pratt [651]. Pre [315]. Pre-Medical [315]. Precision [281, 63]. Predator [119, 1222, 1406]. Predator-Prey [1222, 1406]. Predicting [1066]. Prediction [808, 925, 1065, 692, 1230]. Predictive [484, 350, 664, 1336]. Preference [1113]. Preferences [572]. preliminary [942]. Prenatal [396]. Presence [538, 937, 1219, 786, 1327, 1382]. Preservation [316]. Preservative [1151]. Press [498]. Pressat [516]. Prevalence [1358, 1247]. prèvisions [773]. Prey [119, 1222, 1406]. Price [1362]. Primer [313, 850, 184]. Principal [444]. Principle [934]. Principles [515, 66]. Pringle [449]. Prior [141, 1398, 31]. Priori [1337]. Probabilistic [678]. Probabilite [438]. Probabilities [161, 319, 48, 1329, 1192, 1063, 262, 965]. Probability [324, 331, 1125, 1130, 347, 348, 883, 45, 1031, 521, 906, 399, 535, 794, 404, 76, 1336, 231, 431, 110, 438, 128, 1016, 1160, 359, 832, 519, 901, 546, 1123, 347, 348, 648, 903, 378, 699, 963]. Probit [34, 946]. Problem [869, 352, 54, 787, 1052, 800, 680, 826, 106, 248, 693, 1221, 596, 686, 973, 975]. Probleme [971]. Problems [341, 610, 33, 177, 184, 1023, 1315, 508, 1170, 1324, 379, 798, 234, 249, 429, 839, 1090, 961, 287, 1249, 466, 971, 305, 651]. Procedure [37, 1023, 1034, 208, 668, 803, 819, 1351, 822, 1366, 1216, 1090, 843, 930, 555, 1094, 861, 981]. Procedures [484, 990, 33, 633, 350, 748, 181, 1179, 77, 414, 933, 820, 688, 583, 1185, 1064, 63].

Proceedings [1293, 884, 519, 902, 399, 535, 512, 1232]. Process [792, 75, 450, 130, 676]. Processes [503, 1176, 663, 458, 133, 1013, 1329, 965, 500]. Processing [525]. Product [543]. Production [255]. Products [1066]. Prof. [1142]. Professional [759]. Prognostic [871, 912, 837]. Program [307, 703, 860]. Programmed [117]. Programmes [191, 926, 1061, 322]. Programming [789, 411, 529, 530]. Programs [723, 1166, 274, 854, 980, 860]. Projection [175, 460, 852]. Projective [78]. Proliferation [223]. Properties [356, 907, 210, 683, 269, 1098]. Proportion [1194]. Proportionality [56, 806]. Proportions [1173, 657, 817, 258, 273, 460, 972, 1257, 890, 513]. Psychiatric [191, 391]. Psychology [226]. Public [322]. Published [495, 496, 625, 626, 627, 628, 741, 742, 743, 744, 878, 879, 880, 1004, 1005, 1006, 1007, 1143, 1144, 1145, 1146, 1290, 1291, 1292]. Puri [311].

Q [667]. Quadratic [1018, 363, 542, 556, 288, 919, 938]. Quadrats [691]. Quadric [656]. Qualitative [403, 215, 1378, 1404, 692, 1384]. Quality [32, 762, 381, 1097]. Quantal [344, 1020, 52, 1077, 562]. Quantit [1020]. Quantitative [1166, 797, 1054, 1190, 411, 87, 274, 703, 1232, 1233, 456, 1302, 1301]. Quasi [717, 608, 1298, 393, 1331]. Quasi-Independence [1331]. Quasi-Independent [717, 608, 393]. Quasi-Symmetry [1298]. Query [1014, 1161, 1023, 58, 59, 1040, 1178, 1044, 1180, 427, 1364, 261, 441, 1087, 1090, 285, 1246, 1407]. Question [130]. Queue [408]. Quick [1110]. Quotient [543].

R [1261, 307, 319, 320, 723, 1121, 1123, 1124, 1279, 1293, 347, 44, 1309, 771, 772, 516, 652, 776, 1026, 1028, 1171, 196, 369, 373, 375, 390, 66, 1326, 537, 662, 665, 666, 667, 78, 1190, 674, 411, 549, 1342, 421, 1206, 689, 435, 114, 1215, 1086, 1376, 445, 840, 698, 449, 704, 847, 712]. R. [1280]. Råde [324]. Rabinovitch [648]. Race [308]. Radiation [308]. Radio [1037]. Radiography [234]. Radioimmunoassay [1403]. Radioligand [911]. Rahman [113]. Ralph [885]. Random [1035, 780, 918, 810, 224, 95, 420, 248, 429, 430, 1080, 272, 711, 138, 1327, 924, 938]. Randomised [71]. Randomization [233, 1201, 1392, 856, 1116]. Randomize [1389]. Randomized [162, 204, 940, 1221]. Randomly [1278, 430, 694, 263, 1339, 1340, 1248]. Randomness [403]. Range [1037, 61, 822]. Rank [1420, 634, 353, 1378, 319]. Ranked [367]. Ranking [821]. Rao [369, 840]. Rapid [371, 225, 1194]. Rasch [220]. Rat [72, 678, 229]. Rate [868, 1377, 1400, 718, 935, 936]. Rater [979]. Raters [979]. Rates [748, 181, 513, 1091, 282, 349, 891, 690]. Ratio [161, 991, 1320, 668, 1196, 240, 419, 262, 276, 844, 1235, 1251, 889, 1335, 1093, 1248]. Ratios [485, 402]. Rayner [268, 449]. Re [1033]. Re-Examination [1033]. reabsorption [670]. Reach [882]. Reactions [633, 271]. Reader [455]. Readings [1270]. Ready [1409]. Real [1051]. Realised [406, 407].

Reanalysis [561]. Reappraisal [1081]. Recall [771]. Recapture [1262, 613, 360, 1341, 243, 560, 125, 471, 531, 697]. Received [12, 13, 14, 148, 150, 151, 152, 300, 318]. Reciprocal [1018]. Records [1129, 1078, 122, 1237, 591, 1241, 191]. Recruitment [613, 697]. Rectangular [378, 1109]. recurrence [692]. Reddingius [457]. reducing [955, 956, 957, 958]. reduction [956, 958]. Reference [305, 1038]. Refinement [1020]. Reflections [1308]. Region [1255]. Regions [383]. Registry [770]. Regression [160, 1276, 889, 42, 754, 1314, 645, 1034, 58, 1058, 81, 928, 225, 550, 552, 1204, 97, 687, 695, 944, 1373, 1220, 267, 446, 581, 280, 284, 709, 461, 1245, 137, 1416, 358, 909, 672, 927, 410, 859, 389, 1362]. Regressions [1156, 1065]. Regressor [569]. Reigart [1195]. Rektorys [79]. Related [316, 1421, 1176, 222, 563, 280, 845]. Relatedness [570]. Relation [1161]. Relations [1022]. Relationship [396, 925, 948, 809]. Relationships [363, 1181, 1063, 1239, 710]. Relative [1010, 363, 1039, 381, 412, 829, 111, 260, 127, 599, 1255]. Relatives [987, 758, 364, 551, 659, 1045]. Release [1309, 952]. Relevant [1051]. Reliability [391, 1202, 424]. Remarks [760]. Remington [674]. Removal [163, 1158, 126]. Remove [609, 580, 581]. Removing [1164]. Renewable [954]. Rényi [546]. Repeatability [1089]. Repeated [612, 1418, 1067, 112, 436, 584, 595, 1105, 1416]. Replicate [615, 212, 1108]. Replicated [1320, 80, 1111]. Replicates [989, 913]. réponses [565]. Report [1275, 1305]. Reported [1091]. Representation [1325, 409]. Reproduction [72, 857]. Reproductive [282]. Research [770, 1039, 197, 200, 1179, 680, 96, 1214, 961, 974, 1409, 101, 686, 443, 321, 369, 66, 1050]. Residual [40, 645, 1391]. Residuals [332, 867, 791, 398, 541, 97, 859]. Resolvable [1108]. Resource [954]. Resources [400]. Respect [1337]. Respiratory [561]. Respond [1417, 1330]. Response [719, 19, 344, 1148, 751, 754, 52, 363, 1022, 1315, 1321, 787, 7, 540, 1334, 673, 417, 940, 1077, 829, 834, 569, 1216, 1084, 946, 1221, 125, 1264, 931, 1061]. Response-Time [751, 754, 569]. Responses [716, 657, 1351, 273, 857, 858, 1416, 565, 472]. Restricted [425, 1392]. Restriction [8, 53, 425]. Restrictions [537]. Results [729, 631, 770, 516, 792, 660]. Retrospective [882, 1156, 234, 563, 947, 889]. return [891]. Review [1259, 1260, 1261, 1262, 149, 153, 154, 155, 301, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 319, 302, 320, 321, 322, 341, 323, 324, 477, 723, 1121, 1122, 1123, 1124, 1125, 1268, 1279, 1269, 1270, 1271, 1293, 176, 177, 1150, 348, 498, 499, 500, 634, 635, 883, 884, 1012, 1299, 1300, 1154, 1301, 503, 504, 755, 355, 1157, 184, 1309, 1310, 1311, 1313, 50, 51, 361, 1316, 509, 510, 511, 513, 514, 515, 647, 648, 649, 650, 651, 761, 762, 763, 764, 765, 766, 767, 768, 769, 770, 771, 772]. Review [189, 366, 516, 517, 518, 519, 520, 652, 773, 774, 775, 898, 899, 900, 901, 902, 903, 904, 1024, 1025, 1026, 1027, 1028, 1029, 1030, 1031, 1170, 1171, 61, 62, 63, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 369, 370, 371, 372, 373, 374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384, 385, 386, 64, 65, 388, 389, 390, 203, 1174, 1326, 526, 1176, 205, 1179, 789, 529, 530, 1047, 1328, 399, 533, 534, 535, 536, 537, 661, 662, 663, 793, 794, 916, 917, 918, 1050, 1051, 665, 666, 667, 797, 798, 799, 920, 921, 922, 1054, 1055, 1056, 1187]. Review [76, 77, 78, 79, 217, 1190, 1191, 219, 546, 674, 220, 221, 408, 1338, 1062, 548, 411, 549, 82, 83, 84, 85, 553, 1195, 89, 1341, 1342, 1343, 230, 1200, 1345, 416, 684, 1352, 241, 421, 422, 558, 242, 1206, 689, 834, 1356, 110, 434, 435, 566, 1359, 113, 114, 115, 1210, 1361, 365, 512, 1211, 1362, 1363, 117, 1215, 1367, 1370, 1372, 1086, 1376, 265, 266, 445, 840, 698, 699, 268, 270, 447, 1224, 1386, 279, 280, 449, 586, 704, 845, 846, 847, 848, 1230, 962, 963, 1231, 1232, 1393, 1396, 1234, 1397, 455, 588, 706, 849, 850]. Review [707, 457, 458, 1244, 462, 712, 1404, 1405, 1406, 463, 464, 135, 788]. Reviews [1295, 347, 747, 776, 66]. Revised [254, 439]. Revisited [657]. Reyment [1026, 549]. RGSP [860]. Rhythm [1107]. Rhythms [969]. Ricciardi [1176]. Rice [1179]. Richard [850]. Richards [1033]. Richter [799]. Richter-Dyn [799]. Ridgman [899]. Riedel [903]. Riggs [184]. Right [1350]. Rising [390, 390]. Risk [1010, 364, 1039, 1059, 1196, 232, 103, 111, 464, 974, 1415, 788]. Risks [48, 646, 56, 1059, 409, 257, 260, 120, 1219, 1243, 1296, 692]. Robert [66, 666, 712]. Robust [33, 398, 461, 1403, 332, 445]. Robustness [269]. Roger [635, 77, 455]. Rogers [445]. Rohatgi [883]. Rohf [1106]. Rohlf [66, 712]. Roles [824]. Romier [1171]. Ronald [200]. root [69]. Roots [95, 420]. Rosen [421]. Rosenblatt [347, 347, 458]. Rotation [346]. Rotational [237]. Rothamsted [860]. Rounded [1178, 1364]. Routt [1195]. Row [783, 264]. Row-and-Column [264]. Row-Column [783]. Roy [84]. rules [345, 555, 971]. Run [682]. Run-Like [682]. Runs [292]. Russell [1167, 1332]. Russian [846, 846]. Ruymgaart [634].

S [149, 309, 315, 323, 477, 1125, 1295, 498, 184, 1316, 515, 647, 769, 652, 900, 1030, 193, 194, 390, 526, 533, 534, 665, 799, 1187, 217, 1343, 1352, 242, 435, 566, 1362, 1363, 1370, 845]. Sachs [221]. Sacrifice [1358, 1247]. Saenger [763]. Safe [1060]. Safety [72, 1354]. Salicylate [72]. saline [670]. Sampath [1187]. Sampl [319]. Sample [1258, 329, 182, 1159, 780, 1173, 1177, 528, 532, 1186, 547, 233, 1069, 436, 571, 1085, 272, 1226, 1388, 124, 286, 709, 1104, 471, 345, 1, 915, 1335, 697, 956, 958, 1225]. Sample-Based [532]. Samples [163, 753, 1034, 1041, 61, 62, 802, 816, 817, 831, 949, 950, 126, 1099, 592, 942]. Sampling [21, 611, 612, 991, 1131, 1268, 32, 175, 1304, 903, 1034, 367, 192, 202, 64, 805, 550, 813, 427, 573, 1218, 581, 4, 127, 584, 702, 844, 1097, 1098, 1394, 286, 287, 468, 1318, 349, 358, 397, 1217, 953, 700, 1093, 1229, 1224]. Samuel [921, 846]. Sandler [78]. Sapio [1206]. Scale [1062, 1210]. Scales [691]. Schaffer [307]. Scheffler [115]. Scheme [286]. Scheuer [1051]. Schistosome [1209]. Schmidt [917]. Schork [674]. Schultz [1054]. Schumaker [219]. Schweitzer [1328]. Science [885, 888, 517, 197, 1029]. Sciences [904, 1269, 197, 77, 674, 1206, 1356, 115, 455, 135]. Scientific [772, 652, 1270, 197]. Scientist [1280, 1326]. Scientists [1123, 767, 422]. score [1317]. Scott [519, 399, 535, 1114]. Screening [1182, 400, 1204, 568, 578, 583]. Scudo [1376]. Searle [435]. Season [1180, 1087]. Seber [505, 531, 706]. Second [1388]. Section [415, 384]. Sections [131]. Segregating [1233]. Selected [160, 191, 374, 81, 122, 1236, 1094, 1124, 198, 199]. Selecting [37, 668, 819, 943, 971]. Selection [1153, 354, 753, 45, 46, 53, 1165, 368, 1321, 914, 75, 806, 808, 218, 406, 407, 673, 1191, 928, 425, 685, 116, 1084, 1375, 129, 589, 134, 970, 466, 425, 926, 1061, 927, 976, 1157]. Selective [426]. Selectively [132]. Self [1165, 545, 807, 456]. Self-Fertilized [456]. Self-Fertilizing [1165]. Selfing [589]. Sellizer [769]. Selon [438]. Semi [1222, 601]. Semi-Markov [601]. Semi-Markovian [1222]. Sensitivity [484, 350, 1337]. Separating [830]. separation [809]. Sequences [40, 217]. Sequential [37, 49, 524, 1323, 1175, 1075, 837, 1246, 1, 1225, 871, 361]. serial [1347]. Serially [90]. Series [1155, 206, 1077, 251, 689, 111, 1367, 855, 462]. Serological [43]. Server [408]. Services [191]. Set [367]. Sets [1161, 100, 833, 264, 843, 136, 924]. Several [612, 730, 37, 394, 810, 250, 691, 949, 950, 584, 1236, 294, 470, 979, 931]. Sewall [230]. Sex [168, 551, 240, 572]. Sex-Linked [551]. Sexes [238]. Sforza [277]. Shafer [962]. Shape [1209, 454, 1396]. Shelby [793]. Sheps [684]. Shift [319]. Should [1214]. Shrinkage [1238]. Sib [758]. Sibs [238, 240, 672]. Sided [870, 750, 802, 915]. Sign [1111]. Signals [858]. Significance [870, 750, 1303, 1038, 227, 681, 99, 1361, 1385]. Significant [748]. Signs [351]. Simak [379]. Similarity [210]. Simon [912]. Simple [42, 507, 653, 1034, 1049, 925, 1194, 1184, 567, 65]. Simulated [96]. Simulation [775, 223, 558, 243, 1099]. Simultaneous [42, 414, 839, 1227]. Sine [228]. Singer [378]. Singh [1190]. Single [655, 1187, 408, 228, 436, 909]. Singleton [723]. Singly [632]. Singular [1314]. Sire [1402]. Sisteme [149]. Site [168]. Six [820]. Sixth [519, 399, 535]. Size [1158, 1173, 804, 415, 1069, 1209, 695, 572, 454, 709, 359, 1217, 286]. size-biased [1217]. Sizes [1258, 1159, 528, 1186, 571, 1104, 956, 958]. Skewness [941]. Sleep [601]. Slope [419, 1251]. Slope-Ratio [419]. Slopes [162, 204]. Small [329, 182, 1041, 1085, 1380, 345, 942, 1381]. Small-Sample [329, 182, 345]. Smirnov [1014, 915]. Smith [768]. Smoke [1354]. Smokers [978]. Smooth [540]. Smoothed [1314]. Sneath [666]. Snedecor [366, 746, 759, 795, 796, 380, 553]. Social [308, 317, 1269, 1405]. Society [1162, 904, 15, 16, 17, 156, 157, 158, 325, 326, 327, 328, 478, 479, 480, 481, 605, 606, 607, 724, 725, 726, 727, 864, 865, 866, 985, 986, 1126, 1127, 1275, 1272, 1273, 1274, 630, 362, 1048, 107, 275, 649]. Sociobiology [922]. Sokal [66, 712, 66, 666]. Solomon [1027]. Solution [1130, 1016, 1160]. Solutions [347]. Solving [341, 177]. Som [771]. Some [305, 178, 36, 43, 754, 183, 1308, 508, 379, 783, 792, 660, 210, 211, 811, 414, 933, 1063, 1064, 237, 439, 440, 1375, 264, 955, 587, 1394, 589, 287, 1249, 472, 788, 610, 990, 356, 957]. Sources [196]. Sowden [561]. Space [473, 233, 816, 1372]. Space-Time [233, 816]. Spaced [225, 430]. Spacing [823, 112]. Spatial [908, 1319, 691, 1371, 1115, 715, 650]. Spatially [133]. Spearman [356, 757]. Spearman-type [757]. Special [1038, 787, 951]. Speciation [1154]. Species [640, 1043, 1199, 1344, 453]. Species-Abundance [640]. Specific [501]. Specifications [281]. Specificity [484, 350]. Specified [269]. Speckman [388]. Spherical [131]. Spike [1187]. Spinivasan [1187]. Spirit [1270]. Split [93]. Spontaneous [387, 201]. Sprent [280]. Sprott [355]. SPSS [860]. Square [378, 80, 545, 807, 813, 831, 838, 69]. Squared [203]. Squares [307, 887, 645, 1053, 251, 123, 1390, 1021, 69, 705]. Stability [895, 1066, 421, 100]. Stage [583, 31, 861, 981]. Stages [116, 852]. Stand [1092, 1228]. Standard [70, 73, 229, 931]. Standardization [1398]. Standards [1066]. Statements [1275]. States [375, 385, 1238]. Statforms. [382]. statisques [773]. Statistic [1014, 682, 1083, 1099, 915, 1248]. Statistical [311, 320, 176, 746, 503, 752, 755, 355, 47, 648, 517, 908, 63, 193, 199, 369, 373, 379, 382, 384, 394, 533, 661, 662, 1333, 1197, 416, 93, 418, 1351, 110, 1360, 1214, 1216, 1086, 840, 704, 846, 588, 462, 292, 1409, 1411, 1112, 1198, 557, 565, 66, 302, 321, 322, 477, 513, 514, 515, 651, 772, 189, 366, 775, 776, 904, 200, 380, 381, 383, 1179, 537, 84, 553, 1363, 455, 135]. Statisticians [647, 370, 512]. Statistics [153, 154, 155, 341, 1123, 1268, 1269, 519, 898, 901, 367, 61, 62, 191, 196, 198, 66, 399, 534, 535, 917, 801, 414, 1071, 835, 1211, 838, 699, 122, 1263, 885, 888, 1317, 926, 1061, 1347, 951, 322, 347, 306, 312, 324, 1121, 1124, 1279, 177, 347, 883, 1311, 51, 361, 765, 904, 1030, 1031, 1171, 205, 399, 916, 76, 1345, 242, 113, 114, 117, 265, 698, 449, 586, 963, 1234, 850, 635, 649, 390, 536, 1051, 219, 674, 115, 1370, 847]. Statistik [220]. statistique [565, 1244]. statistiques [89]. Statistische [221]. Statistischen [293]. Steele [1386]. Stein [763]. Step [668, 819]. Step-Down [668]. Step-up [819]. stepwise [555]. Stereographic [460]. Steutel [316]. Stewart [247]. Still [934]. stimulus [565]. Stochastic [868, 1149, 178, 500, 1296, 186, 1322, 223, 86, 937, 939, 251, 1355, 1387, 467, 910, 670, 935, 936, 655, 503, 1170, 389, 663, 799, 1056, 1187, 85, 1343]. Stolyar [434]. Straight [719]. Strategies [1153]. Strategy [1253]. Stratification [106]. Stratified [175, 1099]. Streams [32]. Strengthening [291]. Structural [514, 1239]. Structure [520, 1344, 1092, 1228, 458, 859]. Strum [529]. Stuart [302]. Studentized [61]. Students [315, 268]. Studies [609, 631, 1156, 1037, 381, 1044, 1057, 1193, 552, 677, 1196, 1348, 424, 826, 1076, 563, 260, 947, 580, 581, 1395, 851, 889, 358, 1317, 1103, 729, 1295, 314, 1295, 1031]. Study [637, 195, 790, 243, 1359, 271, 1389, 1099, 523, 1415]. Studying [751, 572, 282, 969]. Styan [533]. Sub [1034]. Sub-Sampling [1034]. subcategory [890]. Subdivided [695]. Subdivisées [695]. Subject [1129, 387, 591, 1241]. Subjects [1417, 1330]. Subset [1389, 1414]. Successive [991, 578, 844, 1093]. Sufficient [1030]. Suggested [779]. Sum [1053, 1077]. Sums [1420, 353, 645, 918]. Super [805]. Super-Population [805]. Superparasitism [1057]. Supplementing [867, 791]. Supplements [247]. Surface [834, 1084]. Surfaces [1321, 95, 420]. Survey [1268, 655, 234, 980, 860, 127, 64, 79]. Surveys [173, 1312, 377]. Survival [1149, 1151, 638, 47, 760, 668, 214, 818, 819, 937, 1070, 1350, 1358, 1083, 1365, 574, 1373, 440, 1220, 1237, 138, 1011, 1296, 891, 1327, 1075, 1254, 1247]. Survival/ [1358]. Survival/Sacrifice [1247]. Susceptibles [394]. Susser [763]. Sustainable [954]. Symmetric [104, 599, 68]. Symmetrical [1079, 977]. Symmetry [1298, 291]. Symposium [519, 399, 535, 192, 202]. Symptoms [561]. Synapses [952]. Synchronous [137]. Synthèse [425]. Synthesis [425, 703, 137, 1184]. Synthetic [563, 1412, 1413]. System [632, 1309, 403, 1341, 421, 855]. Systematic [1410]. Systems [716, 1293, 43, 100, 1355, 1211, 127, 463, 1157, 266, 1406, 558].

T [303, 176, 504, 1157, 649, 903, 1024, 1170, 380, 533, 1191, 553, 1359, 1210, 1211, 279, 1232, 462]. Tabard [1244]. Table [161, 330, 39, 1300, 188, 815, 262, 1252, 139, 1327]. Tables [717, 482, 1124, 1420, 41, 1298, 180, 353, 1164, 647, 1038, 198, 370, 378, 385, 392, 1049, 1331, 405, 541, 1186, 929, 560, 102, 827, 833, 838, 291, 1264, 641, 894, 669, 1353, 465, 976, 66, 323]. Tabulated [1260]. Tag [613, 125, 697]. Tag-Recapture [613, 697]. Taille [695]. Tanner [1259, 1359]. Tanur [390]. Target [1394]. Targets [378]. Taste [1113]. Tautu [663]. Taxonomy [902, 666]. Teaching [324]. Technique [391, 237, 940, 105, 1081, 1208, 1221, 690, 1244]. Techniques [311, 499, 775, 790, 1352, 530, 1224]. Teissier [448]. Telemetry [1037]. Temperature [285]. temporally [472]. Teoriei [149]. Teratogenicity [857]. Term [1155, 1015]. Termination [1336, 1351, 1216]. Terms [1321, 1084, 373, 846]. Terry [1032, 1074]. Test [20, 730, 39, 748, 185, 357, 1022, 60, 528, 1186, 99, 104, 246, 247, 831, 691, 1213, 1371, 574, 841, 579, 1095, 283, 708, 461, 1106, 294, 1110, 1111, 1113, 1116, 893, 68, 557, 942, 451]. Testimating [746]. Testing [141, 1317, 1319, 1038, 909, 1041, 61, 62, 384, 1175, 204, 657, 668, 223, 1339, 1340, 828, 250, 1366, 950, 273, 959, 592, 1246, 972, 930, 1385, 1248, 162, 1278]. Testipredicting [746]. Testitesting [746]. Tests [473, 329, 870, 634, 1298, 182, 750, 779, 1320, 59, 61, 405, 810, 1062, 227, 681, 233, 816, 687, 827, 256, 257, 258, 261, 697, 1378, 1379, 964, 711, 595, 972, 136, 291, 1317, 915, 101, 587, 613]. Textbooks [917]. Thalamic [858]. Their [317, 1125, 61, 62, 1057, 545, 807, 1361]. Theoretic [252]. Theoretical [1023, 113, 1376, 1090, 1271, 1293, 504]. Théorie [425, 89]. Theories [589]. Theory [348, 634, 883, 503, 758, 56, 776, 367, 384, 399, 400, 665, 1059, 230, 232, 421, 425, 110, 257, 698, 584, 962, 1399, 854, 1412, 108, 1413, 612, 1125, 535, 464, 794]. There [1156, 252, 1214, 467]. Thermodynamic [858]. Thinking [514, 848]. Thirds [615, 212]. Thomas [1054, 1223]. Thompson [1055]. Three [163, 989, 1298, 913, 547, 1199, 126, 1115]. Three-Dimensional [1298]. Threshold [540, 1089]. Tiao [661]. Tiberiu [884]. Ties [1325]. Time [882, 751, 754, 46, 47, 387, 206, 233, 816, 1070, 1351, 251, 432, 689, 569, 1216, 1367, 282, 462, 596, 345, 1013, 359, 910, 809, 676, 1401, 473].

Time-Dependence [751]. Time-Dependency [432]. time-dependent [345, 910]. Times [1354, 1219, 965]. Timing [1391]. Tolerance [502, 383, 543, 802]. Tools [1299]. Topics [376, 1176, 563, 278, 259, 316, 280, 845, 176]. Torrens [774]. Torrens-Ibem [774]. Total [1180, 539, 1087]. Totally [118]. Toutenburg [1230]. Toxicity [1391]. Toxicological [1334, 1203, 857]. Tracts [199]. traditional [1011]. Trains [655, 1187]. Trait [1040, 1233, 786]. Traitement [773, 525]. Traits [1089, 134, 970, 1250, 1302, 692, 1384]. Transect [1131, 1304, 1218, 702, 1394, 892, 683, 1229]. Transects [682, 292]. Transformations [143, 636, 824, 809]. Transfusion [521]. Transient [1238]. Transition [825]. Transmission [331, 404, 231]. Transmitter [952]. Trap [125]. Trapped [1078]. Trapping [1078]. Treatment [730, 871, 1417, 40, 1015, 521, 912, 1330, 250, 571, 837, 1374, 281, 1095, 1096, 595, 1407, 294, 1103]. Treatments [989, 1320, 1038, 913, 211, 264, 1389]. Tree [391, 544, 1092]. Trees [900, 209, 1080, 1228, 292, 596, 1055]. Trends [1257, 443]. Triads [1152]. Trial [871, 1151, 1036, 912, 940, 837, 1235]. Trials [305, 614, 1419, 1015, 524, 528, 658, 2, 1188, 1336, 1201, 1069, 1351, 1366, 1216, 1374, 1379, 1389, 1096, 1392, 1105, 714, 977, 1411, 1075, 1418]. Triangle [1113]. Triangular [717, 1331]. Tribus [1342]. Truncated [804, 1388]. Truncation [116]. Tsokos [348]. Tübingen [1293]. tubules [670]. Tucker [239]. Tukey [192, 202, 1025, 445, 579]. Tumor [574, 1238, 1242]. Twenty [362]. Twenty-Five [362]. Twin [784]. Twin-Pair [784]. Two [1258, 161, 163, 319, 329, 614, 615, 868, 990, 1276, 1418, 1298, 182, 1156, 1157, 1159, 641, 1161, 1164, 57, 522, 1043, 1173, 212, 2, 214, 804, 547, 1194, 933, 681, 233, 816, 817, 418, 822, 423, 245, 433, 833, 256, 1360, 258, 116, 262, 440, 264, 441, 576, 959, 126, 964, 589, 594, 1245, 972, 1105, 469, 136, 599, 1108, 1110, 138, 471, 786, 915, 1185, 1064, 935, 936, 555, 1075, 451, 1102, 861, 981]. Two-Compartment [868, 935, 936]. Two-dimensional [641]. Two-Locus [1157]. Two-Period [614, 1418, 2, 418, 1105]. Two-Replicate [1108]. Two-Sampl [319]. Two-Sample [329, 182, 233, 471, 915]. two-stage [861, 981]. Two-Thirds [615, 212]. Two-Way [990, 1164, 933, 822, 433, 833, 256, 441, 1185, 1064, 451]. Tyndall [723]. Type [523, 787, 660, 1318, 757, 1071, 748, 181, 592]. Types [804, 547, 685, 440, 576, 1407].

U [500]. Udny [302]. Ultimate [1130, 45, 1016, 1160]. Unbalanced [272, 124, 711, 288, 859, 1101, 1102]. Unbiased [808, 288, 919, 690]. Uncontrolled [552]. Uncorrelated [437]. Underlying [1278, 269, 1339, 1340]. Undesirable [37]. Unequal [505, 695, 267, 852, 1104]. Unequally [1320, 951]. Ungarischen [386]. Uniformity [305]. unimodal [557]. United [385]. Universities [197]. Unknown [37, 1042, 390]. Unreplicated [944]. Unresolved [1324]. unvollständiger [971]. Unweighted [1189]. Up-and-Down [187]. Upton [1260]. Urn [1125]. Urquhart [845]. Usanis [307]. Use [141, 716, 614, 991, 1270, 175, 751, 646, 1036, 61, 62, 370, 912, 528, 2, 1186, 93, 418, 1203, 108, 569, 947, 581, 122, 844, 714, 136, 1416, 601, 1093]. Used [925, 568, 788]. Useful [719, 1351, 1216, 1092, 1395]. Uses [1297, 191]. Using [988, 1420, 33, 882, 746, 657, 668, 237, 1070, 691, 1082, 1365, 950, 124, 858, 1278, 31, 353, 1339, 1340, 414, 555]. Utilizing [1180, 1087].

V [883, 755, 355, 776, 904, 1029, 1030, 536, 918, 1376, 963]. Vaccinations [155]. Václav [1142]. Valand [1208]. Valid [934]. Validation [109]. validatory [932]. Valleron [1397]. Value [484, 350, 92, 1079]. Valued [450]. Values [1277, 925, 5, 426, 441]. Vandam [195]. Variability [1316, 673, 236]. Variable [937, 688, 1414, 359, 644]. Variable-Subset [1414]. Variables [20, 612, 751, 357, 60, 668, 928, 819, 94, 1070, 569, 584, 525, 926, 1061, 927, 555, 943, 918]. Variance [21, 485, 611, 728, 1129, 175, 33, 36, 37, 637, 356, 896, 897, 776, 1169, 1042, 64, 402, 542, 74, 1189, 814, 817, 236, 556, 433, 259, 269, 272, 842, 4, 701, 123, 124, 278, 452, 288, 1252, 856, 471, 1021, 1184, 1335, 671, 1073, 1227, 591, 1241, 485, 1122]. Variance-Covariance [1189]. Variances [36, 639, 753, 357, 1189, 822, 267, 577, 272, 124, 967, 451, 705]. Variate [34, 290]. Variates [543, 136, 951]. Variation [1154, 572, 467]. Variety [564]. Various [62, 1181]. Vascular [424]. Vector [450, 525]. Vector-Valued [450]. vectorielles [525]. Vectors [1314, 417]. ventes [773]. Versuche [293]. versus [700]. via [854]. Viability [784, 459]. Victoria [798]. Viewpoint [805, 834]. Vincent [1054]. Virological [679]. Visibility [1312]. Visual [1303]. Vital [154, 51, 191]. vitro [109]. vivo [109]. Vol [399, 535, 794, 917, 421, 558]. Volterra [1376]. Volume [30, 153, 154, 155, 174, 308, 317, 342, 497, 629, 745, 881, 1008, 1124, 1147, 1294, 1154, 1316, 647, 61, 62, 197, 198, 230]. Vorgeschichte [293]. Vortäge [386]. vs [232].

W [1261, 306, 313, 314, 316, 723, 503, 1313, 759, 649, 651, 769, 189, 366, 518, 899, 1025, 192, 194, 195, 202, 373, 380, 388, 390, 530, 1328, 533, 916, 795, 796, 665, 408, 83, 553, 566, 115, 1370, 445, 1224, 1234, 462]. W. [295]. Wadley [98]. Wald [1099]. Wallace [308, 317, 270]. Wallis [306]. Walpole [1123]. Walsh [494]. Walter [797, 1027]. Wanderungsbewegungen [1328]. Want [1079]. Warner [1221]. Water [32, 206, 1097]. Waterfowl [127]. Waters [189]. Wave [228]. Way [482, 990, 1164, 392, 933, 822, 433, 833, 256, 441, 272, 288, 1185, 1064, 938, 1073, 451]. Wayne [747]. Weather [1365]. Weber [1393]. Webs [1372]. Weibull [575, 1254]. Weight [229]. Weighted [1217, 1021]. Weights [1337, 932]. Weinberg [893, 1041]. Weiner [526]. Wen [1056]. Wen-Hsiung [1056]. Westfall [652]. Westphal [1195]. Wetherill [187, 361, 588]. where [423]. Which [843, 1389, 437, 467]. White [229]. Whitlock [1200]. Who [322, 153, 154, 155, 168]. Wickman [378]. Wiener [438]. Wijngaaden [82]. wildlife [683, 1217]. Wilhelm [707]. Wilks [193]. Will [1417, 1330]. Williams [1031, 1095]. Wilson [313, 922, 1405]. Winberg [219]. Winsor [192, 202]. Winter [763]. withdrawals [1353]. Within [1325, 1165, 1316, 1032]. Within-Pair [1325, 1032]. without [410, 1388, 138]. Wolock [388]. Wood [1151, 566]. Work [287, 443]. Workbook [747]. Workers [200]. Works [1376]. World [173, 201, 153, 154, 155]. Wright [1316, 230].

X [1049, 353].

Yates [374]. Year [1305]. Years [362]. Yeast [1046]. Yield [1181, 954]. Yield-Density [1181]. Youden [295]. Yule [302].

Z [763]. Zero [730, 657, 804, 91, 273, 294, 470]. Zero-Truncated [804]. Ziegler [1376, 916]. Zuordnungsregeln [971]. zur [1328, 293, 465]. Zyskind [624].

References

Chapman:1952:IMS

[1] Douglas G. Chapman. Inverse, multiple and sequential sample censuses. Biometrics, 8(4):286–306, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See erratum [1225].

Grizzle:1965:TPC

[2] James E. Grizzle. The two-period change-over design and its use in clinical trials. Biometrics, 21(2):467–480, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528104. See corrections [614].

Das:1966:IBD

[3] M. N. Das and G. A. Kulkarni. Incomplete block designs for bio-assays. Biometrics, 22(4):706–729, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528070. See correction [159].

Scheinberg:1966:NSV

[4] Eliyahu Scheinberg. 220. Note: The sampling variance of the correlation coefficients estimated in genetic experiments. Biometrics, 22(1): 187–191, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528225. See

corrections [21].

Hinkelmann:1968:MVP

[5] Klaus Hinkelmann. Missing values in partial diallel cross experiments. Biometrics, 24(4):903–913, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528879. See correction [1277].

Grizzle:1969:ACD

[6] James E. Grizzle, C. Frank Starmer, and Gary G. Koch. Analysis of categorical data by linear models. Biometrics, 25(3):489–504, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528901. See corrections [18].

Grizzle:1969:AGD

[7] James E. Grizzle and David M. Allen. Analysis of growth and dose response curves. Biometrics, 25(2):357–381, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528795. See corrections [19].

Anderson:1970:REL

[8] Virgil L. Anderson. Restriction errors for linear models (an aid to develop models for designed experiments). Biometrics, 26(2):255–268, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529073.

Anonymous:1970:Aa

[9] Anonymous. Abstracts. Biometrics, 26(1):166–169, March 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529062.

Anonymous:1970:Ab

[10] Anonymous. Abstracts. Biometrics, 26(2):356, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529088.

Anonymous:1970:Ac

[11] Anonymous. Abstracts. Biometrics, 26(3):598–608, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529124.

Anonymous:1970:BRa

[12] Anonymous. [Books received]. Biometrics, 26(1):164–165, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529061.

Anonymous:1970:BRb

[13] Anonymous. [Books received]. Biometrics, 26(3):596–597, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529123.

Anonymous:1970:BRc

[14] Anonymous. [Books received]. Biometrics, 26(4):872–873, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528750.

Anonymous:1970:BSa

[15] Anonymous. The Biometric Society. Biometrics, 26(1):170–174, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529063.

Anonymous:1970:BSb

[16] Anonymous. The Biometric Society. Biometrics, 26(2):357–363, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529089.

Anonymous:1970:BSc

[17] Anonymous. The Biometric Society. Biometrics, 26(3):609–610, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529125.

Anonymous:1970:CAC

[18] Anonymous. Corrections: Analysis of categorical data by linear models. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528736. See [6].

Anonymous:1970:CAG

[19] Anonymous. Corrections: Analysis of growth and dose response curves. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528735. See [7].

Anonymous:1970:CNT

[20] Anonymous. Corrections: A new test of association for continuous variables. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528734. See [60].

Anonymous:1970:CSV

[21] Anonymous. Corrections: The sampling variance of the correlation coefficients estimated in genetic experiments. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528737. See [4].

Anonymous:1970:FMa

[22] Anonymous. Front matter. Biometrics, 26(1):i–ii, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529038.

Anonymous:1970:FMb

[23] Anonymous. Front matter. Biometrics, 26(2):i–ii, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529066.

Anonymous:1970:FMc

[24] Anonymous. Front matter. Biometrics, 26(3):??, September 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529092.

Anonymous:1970:FMd

[25] Anonymous. Front matter. Biometrics, 26(4):i–ii, December 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528709.

Anonymous:1970:NAa

[26] Anonymous. News and announcements. Biometrics, 26(1):175–179, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529064.

Anonymous:1970:NAb

[27] Anonymous. News and announcements. Biometrics, 26(2):364, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529090.

Anonymous:1970:NAc

[28] Anonymous. News and announcements. Biometrics, 26(3):611–612, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529126.

Anonymous:1970:NAd

[29] Anonymous. News and announcements. Biometrics, 26(4):883–885, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528752.

Anonymous:1970:VI

[30] Anonymous. Volume information. Biometrics, 26(4):i–xxxiii, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528708.

Arnold:1970:DSE

[31] J. C. Arnold and H. A. Al-Bayyati. On double-stage estimation of the mean using prior knowledge. Biometrics, 26(4):787–800, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528724.

Arnold:1970:MSP

[32] J. C. Arnold. A Markovian sampling policy applied to water quality monitoring of streams. Biometrics, 26(4):739–747, December 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528720.

Arvesen:1970:RPV

[33] James N. Arvesen and Thomas H. Schmitz. Robust procedures for variance component problems using the jackknife. Biometrics, 26(4):677–686, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528715.

Ashford:1970:MVP

[34] J. R. Ashford and R. R. Sowden. Multi-variate probit analysis. Biometrics, 26(3):535–546, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529107.

Bartlett:1970:AD

[35] M. S. Bartlett. Age distributions. Biometrics, 26(3):377–385, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529094.

Basson:1970:CVS

[36] Rodney P. Basson. A comparison of variances of some estimators in the balanced incomplete block (BIB) variance components model. Biometrics, 26(4):657–669, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528713.

Bechhofer:1970:CNU

[37] Robert E. Bechhofer. Correction note: An undesirable feature of a sequential multiple-decision procedure for selecting the best one of several normal populations with a common unknown variance. Biometrics, 26(2):347–349, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529085.

Becker:1970:CPP

[38] Niels G. Becker. Control of a pest population. Biometrics, 26(3):365–375, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529093.

Bennett:1970:NMT

[39] B. M. Bennett and R. E. Underwood. 283. Note: On McNemar’s test for the 2 × 2 table and its power function. Biometrics, 26(2):339–343, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529083.

Berenblut:1970:NTS

[40] I. I. Berenblut. 277. Note: Treatment sequences balanced for the linear component of residual effects. Biometrics, 26(1):154–156, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529055.

Berry:1970:NPA

[41] G. Berry. 287 note: Parametric analysis of disease incidences in multiway tables. Biometrics, 26(3):572–579, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529111.

Brown:1970:NSC

[42] Barry W. Brown. 274. Note: Simple comparisons of simultaneous regression lines. Biometrics, 26(1):143–144, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529052.

Brown:1970:SMD

[43] Barry W. Brown, Henry Davis, and Harold Goodman. On some methods of drawing inferences about antigen systems from serological data. Biometrics, 26(4):701–711, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528717.

Calvin:1970:OJC

[44] Lyle D. Calvin. Obituary: Jerome C. R. Li 1914–1969. Biometrics, 26(1): 161, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529057.

Carr:1970:ESDa

[45] R. N. Carr and R. F. Nassar. Effects of selection and drift on the dynamics of finite populations. i. ultimate probability of fixation of a favorable allele. Biometrics, 26(1):41–49, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529043.

Carr:1970:ESDb

[46] R. N. Carr and R. F. Nassar. Effects of selection and drift on the dynamics of finite populations II. Expected time to fixation or loss of an allele. Biometrics, 26(2):221–227, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529070.

Chang:1970:SMA

[47] Potter C. Chang. Statistical models for animal survival time in mouse lymphoma. Biometrics, 26(4):749–766, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528721.

Chiang:1970:CRC

[48] Chin Long Chiang. Competing risks and conditional probabilities. Biometrics, 26(4):767–776, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528722.

Choi:1970:SDB

[49] S. C. Choi and V. A. Clark. Sequential decision for a binomial parameter with delayed observations. Biometrics, 26(3):411–420, September 1970.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529098.

Cornell:1970:BRD

[50] Richard G. Cornell. Book review: Demographic Analysis, by B. Benjamin. Biometrics, 26(1):163–164, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529059.

Cornell:1970:BRH

[51] Richard G. Cornell. Book review: Health and Vital Statistics, by B. Benjamin. Biometrics, 26(1):164, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529060.

Cornell:1970:EQR

[52] Richard G. Cornell and Norman J. Petersen. Evaluation of a quantal response model with estimated concentrations. Biometrics, 26(4):713– 722, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528718.

Cunningham:1970:RSI

[53] E. P. Cunningham, R. A. Moen, and T. Gjedrem. Restriction of selection indexes. Biometrics, 26(1):67–74, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529045.

Curnow:1970:NCP

[54] R. N. Curnow. 285 note: A classification problem involving human chromosomes. Biometrics, 26(3):559–566, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529109.

Das:1970:CAF

[55] M. N. Das and R. C. Jain. On component analysis of factorial and fractional factorial experiments. Biometrics, 26(4):823–833, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528727.

David:1970:NCP

[56] H. A. David. 282. Note: On Chiang’s proportionality assumption in the theory of competing risks. Biometrics, 26(2):336–339, June 1970.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529082.

Detre:1970:NCT

[57] Katherine Detre and Colin White. 294. Note: The comparison of two Poisson-distributed observations. Biometrics, 26(4):851–854, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528732.

Duncan:1970:QMC

[58] David B. Duncan. 273. query: Multiple comparison methods for comparing regression coefficients. Biometrics, 26(1):141–143, March 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529051.

Dunnett:1970:QMC

[59] C. W. Dunnett. 272. query: Multiple comparison tests. Biometrics, 26(1):139–141, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529050.

Elston:1970:NTA

[60] R. C. Elston and John Stewart. A new test of association for continuous variables. Biometrics, 26(2):305–314, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529077. See corrections [20].

F:1970:BROa

[61] W. T. F. Book review: Order Statistics and Their Use in Testing and Estimation: Volume 1. Tests Based on Range and Studentized Range of Samples from a Normal Population by H. L. Harter. Biometrics, 26(4):866–867, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528742.

F:1970:BROb

[62] W. T. F. Book review: Order Statistics and Their Use in Testing and Estimation: Volume 2. Estimates Based on Order Statistics of Samples from Various Populations, by H. L. Harter. Biometrics, 26(4):867–868, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528743.

F:1970:BRP

[63] W. T. F. Book review: Precision Measurement and Calibration-Statistical Concepts and Procedures, by H. H. Ku. Biometrics, 26(4):868– 870, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528744.

Federer:1970:BRE

[64] W. T. Federer. Book review: Estimation and Sampling Variance in the Health Interview Survey, by Judy A. Bean. Biometrics, 26(4):864, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528739.

Federer:1970:BRS

[65] W. T. Federer. Book review: Simple Distribution-Free Confidence Intervals for a Difference in Location, by P. Van Der Laan. Biometrics, 26(4): 871, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528748.

Fienberg:1970:BRB

[66] Stephen E. Fienberg. Book reviews: Biometry. The Principles and Practice of Statistics in Biological Research, by Robert R. Sokal, F. James Rohlf; Statistical Tables, by F. James Rohlf, Robert R. Sokal. Biometrics, 26(2):351–355, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529087.

Foss:1970:MEC

[67] Stephen D. Foss. A method of exponential curve fitting by numerical integration. Biometrics, 26(4):815–821, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528726.

Gart:1970:LMP

[68] John J. Gart. A locally most powerful test for the symmetric folded binomial distribution. Biometrics, 26(1):129–138, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529049.

Gaylor:1970:CEM

[69] D. W. Gaylor, H. L. Lucas, and R. L. Anderson. Calculation of expected mean squares by the abbreviated Doolittle and square root methods. Biometrics, 26(4):641–655, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528712.

Gill:1970:NSE

[70] J. L. Gill. 276. Note: On standard errors for estimates of fitness. Biometrics, 26(1):149–153, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529054.

Gomes:1970:NEM

[71] F. Pimentel Gomes. 281. Note: An extension of the method of joint analysis of experiments in complete randomised blocks. Biometrics, 26(2): 332–336, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529081.

Gross:1970:ESF

[72] M. Adrian Gross, O. Garth Fitzhugh, and Nathan Mantel. Evaluation of safety for food additives: an illustration involving the influence of methyl salicylate on rat reproduction. Biometrics, 26(2):181–194, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529067.

Hamdan:1970:SEE

[73] M. A. Hamdan. [On standard errors for estimates of fitness]. Biometrics, 26(3):586–587, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529114.

Harvey:1970:EVC

[74] Walter R. Harvey. Estimation of variance and covariance components in the mixed model. Biometrics, 26(3):485–504, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529104.

Harville:1970:ASP

[75] David A. Harville. Approximating the selection process. Biometrics, 26(1):51–66, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529044.

Hedayat:1970:BRB

[76] A. Hedayat. Book review: Basic Concepts of Probability and Statistics, by J. L. Hodges, Jr., E. L. Lehman. Biometrics, 26(3):589–590, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529117.

Hedayat:1970:BRE

[77] A. Hedayat. Book review: Experimental Design: Procedures for the Behavioral Sciences, by Roger E. Kirk. Biometrics, 26(3):590–593, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529119.

Hedayat:1970:BRI

[78] A. Hedayat. Book review: An Introduction to Finite Projective Planes, by A. A. Albert, R. Sandler. Biometrics, 26(1):162–163, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529058.

Hedayat:1970:BRS

[79] A. Hedayat. Book review: Survey of Applicable Mathematics, by Karel Rektorys. Biometrics, 26(3):594–596, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529121.

Hedayat:1970:NEM

[80] A. Hedayat and W. T. Federer. 279. Note: An easy method of constructing partially replicated latin square designs of order n for all n ¿ 2. Biometrics, 26(2):327–330, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529079.

Hill:1970:NDE

[81] William G. Hill. 286 note: Design of experiments to estimate heritability by regression of offspring on selected parents. Biometrics, 26(3):566–571, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529110. See

correction [160].

Innis:1970:BRA

[82] George Innis. Book review: Algøl 68, by A. Van Wijngaaden, B. M. Mailloux, J. E. L. Peck, C. H. A. Koster. Biometrics, 26(4):871–872, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528749.

Innis:1970:BRI

[83] George Innis. Book review: Industrial Dynamics, by Jay W. Forrester. Biometrics, 26(4):865, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528741.

Innis:1970:BRS

[84] George Innis. Book review: Statistical Computation, by Roy C. Milton, John A. Nelder. Biometrics, 26(4):871, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528747.

Isaacson:1970:BRS

[85] Dean Isaacson. Book review: Stochastic Convergence, by E. Lukacs. Biometrics, 26(4):870, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528746.

Iyer:1970:SMG

[86] K. S. S. Iyer and V. N. Saksena. A stochastic model for the growth of cells in cancer. Biometrics, 26(3):401–410, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529097.

Jayakar:1970:DEL

[87] S. D. Jayakar. On the detection and estimation of linkage between a locus influencing a quantitative character and a marker locus. Biometrics, 26(3):451–464, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529101.

John:1970:NFF

[88] Peter W. M. John. 278. Note: A fraction of the 4^{3} factorial design for estimating main effects. Biometrics, 26(1):157–160, March 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529056.

Jolicoeur:1970:BRT

[89] Pierre Jolicoeur. Book review: Théorie et méthodes statistiques. Applications agronomiques, by Pierre Dagnelie. Biometrics, 26(4):864–865, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528740.

Jones:1970:CDM

[90] Richard H. Jones, David H. Crowell, and Linda E. Kapuniai. Change detection model for serially correlated multivariate data. Biometrics, 26(2): 269–280, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529074.

Katti:1970:LZP

[91] S. K. Katti and A. Vijaya Rao. The log-zero-poisson distribution. Biometrics, 26(4):801–813, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528725.

Kharshikar:1970:NEV

[92] A. V. Kharshikar. 284. Note: On the expected value of S^{2}/X̄. Biometrics, 26(2):343–346, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529084.

Koch:1970:UNP

[93] Gary G. Koch. The use of non-parametric methods in the statistical analysis of a complex split plot experiment. Biometrics, 26(1):105–128, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529048.

Kurczynski:1970:GDD

[94] T. W. Kurczynski. Generalized distance and discrete variables. Biometrics, 26(3):525–534, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529106.

Lang:1970:LDP

[95] A. R. G. Lang and F. M. Melhuish. Lengths and diameters of plant roots in non-random populations by analysis of plane surfaces. Biometrics, 26(3):421–431, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529099. See correction [420].

Lee:1970:SPM

[96] P. N. Lee. The simulated population method of analysis of animal painting experiments in cancer research. Biometrics, 26(4):777–785, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528723.

Machin:1970:NRC

[97] D. Machin. 290. Note: Regression with correlated residuals: an example from competition experiments. Biometrics, 26(4):835–840, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528728.

Maloney:1970:OFM

[98] Clifford J. Maloney. Obituary: Francis M. Wadley, 1892–1969. Biometrics, 26(2):350, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529086.

Maloney:1970:STG

[99] Clifford J. Maloney and Suresh C. Rastogi. Significance test for Grubbs’s estimators. Biometrics, 26(4):671–676, December 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528714.

Mandel:1970:NED

[100] S. P. H. Mandel. 291. Note: The equivalence of different sets of stability conditions for multiple allelic systems. Biometrics, 26(4):840–845, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528729.

Mantel:1970:CPM

[101] Nathan Mantel and John C. Bailar III. A class of permutational and multinomial tests arising in epidemiological research. Biometrics, 26 (4):687–700, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528716.

Mantel:1970:ICT

[102] Nathan Mantel. Incomplete contingency tables. Biometrics, 26(2):291– 304, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529076.

Mantel:1970:MMI

[103] Nathan Mantel, John C. Bailar, and Iii. Model I or model II in competing risk analysis. Biometrics, 26(4):861–863, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528738.

Mantel:1970:NAT

[104] Nathan Mantel. 293. Note: An alternative test for the symmetric folded binomial distribution. Biometrics, 26(4):848–851, December 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528731.

Mantel:1970:TNM

[105] Nathan Mantel and Ranchhodbhai S. Valand. A technique of nonparametric multivariate analysis. Biometrics, 26(3):547–558, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529108.

Marriott:1970:NPO

[106] F. H. C. Marriott. 292. Note: A problem of optimum stratification. Biometrics, 26(4):845–847, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528730.

Martin:1970:BS

[107] L. Martin. The Biometric Society. Biometrics, 26(4):874–882, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528751.

Maruyama:1970:UGT

[108] Takeo Maruyama and Norikazu Yasuda. Use of graph theory in computation of inbreeding and kinship coefficients. Biometrics, 26(2):209–219, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529069.

McHugh:1970:VVE

[109] Richard B. McHugh, John F. Van Pilsum, and Franklin W. Briese. Validation of an in vitro enzyme assay indirectly measuring an in vivo biosynthesis. Biometrics, 26(2):195–207, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529068.

Meeden:1970:BRI

[110] Glen Meeden. Book review: Introduction to Probability Theory and Statistical Inference, by Harold J. Larson. Biometrics, 26(4):870, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528745.

Miettinen:1970:ERR

[111] Olli S. Miettinen. Estimation of relative risk from individually matched series. Biometrics, 26(1):75–86, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529046.

Morrison:1970:OSR

[112] Donald F. Morrison. The optimal spacing of repeated measurements. Biometrics, 26(2):281–290, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529075.

Mullen:1970:BRC

[113] Kenneth Mullen. Book review: A Course in Theoretical Statistics, by N. A. Rahman. Biometrics, 26(3):593–594, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529120.

Mullen:1970:BRI

[114] Kenneth Mullen. Book review: Introduction to Statistics, by R. A. Hultquist. Biometrics, 26(3):590, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529118.

Mullen:1970:BRS

[115] Kenneth Mullen. Book review: Statistics for the Biological Sciences, by W. C. Scheffler. Biometrics, 26(3):596, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529122.

Namkoong:1970:OAS

[116] Gene Namkoong. Optimum allocation of selection intensity in two stages of truncation selection. Biometrics, 26(3):465–476, September 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529102.

Niven:1970:BRP

[117] B. Susan Niven. Book review: A Programmed Introduction to Statistics, by Freeman F. Elzey. Biometrics, 26(3):588–589, September 1970. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529116.

Pearce:1970:NCT

[118] S. C. Pearce. 280. Note: A comment on totally confounded designs. Biometrics, 26(2):330–332, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529080.

Pearce:1970:NDM

[119] C. Pearce. A new deterministic model for the interaction between predator and prey. Biometrics, 26(3):387–392, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529095.

Pike:1970:NNK

[120] M. C. Pike. 288 note: a note on Kimball’s paper ‘Models for the Estimation of Competing Risks from Grouped Data’. Biometrics, 26(3):579–581, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529112.

Sankaran:1970:NDP

[121] Munuswamy Sankaran. 275. Note: The discrete Poisson–Lindley distribution. Biometrics, 26(1):145–149, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529053.

Schaeffer:1970:NUO

[122] L. R. Schaeffer, L. D. Van Vleck, and J. A. Velasco. 295. Note: The use of order statistics with selected records. Biometrics, 26(4):854–859, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528733.

Searle:1970:EMS

[123] S. R. Searle and R. F. Fawcett. Expected mean squares in variance components models having finite populations. Biometrics, 26(2):243– 254, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529072.

Searle:1970:LSV

[124] S. R. Searle. Large sample variances of maximum likelihood estimators of variance components using unbalanced data. Biometrics, 26(3):505–524, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529105.

Seber:1970:ETR

[125] G. A. F. Seber. The effects of trap response on tag recapture estimates. Biometrics, 26(1):13–22, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529040.

Seber:1970:RMT

[126] G. A. F. Seber and J. F. Whale. The removal method for two and three samples. Biometrics, 26(3):393–400, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529096. See correction [163].

Sen:1970:RES

[127] A. R. Sen. Relative efficiency of sampling systems in the Canadian Waterfowl Harvest Survey, 1967–68. Biometrics, 26(2):315–326, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529078.

Singh:1970:GPD

[128] S. N. Singh and B. N. Bhattacharya. A generalized probability distribution for couple fertility. Biometrics, 26(1):33–40, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529042.

Smith:1970:SIE

[129] W. B. Smith and R. C. Pfaffenberger. Selection index estimation from partial multivariate normal data. Biometrics, 26(4):625–639, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528711.

Spiers:1970:QIP

[130] Philip S. Spiers and Dana Quade. On the question of an infectious process in the origin of childhood leukemia. Biometrics, 26(4):723–737, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528719.

Tallis:1970:EDS

[131] G. M. Tallis. Estimating the distribution of spherical and elliptical bodies in conglomerates from plane sections. Biometrics, 26(1):87–103, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529047.

Thomson:1970:SDM

[132] Glenys J. Thomson. Selectively disadvantageous mutants. Biometrics, 26(2):229–241, June 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529071.

Usher:1970:DMM

[133] M. B. Usher and M. H. Williamson. A deterministic matrix model for handling the birth, death, and migration processes of spatially distributed populations. Biometrics, 26(1):1–12, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529039.

VanVleck:1970:ISD

[134] L. D. Van Vleck. Index selection for direct and maternal genetic components of economic traits. Biometrics, 26(3):477–483, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529103.

Wassertheil:1970:BRS

[135] Sylvia Wassertheil. Book review: Statistical Power Analysis for the Behavioral Sciences, by Jacob Cohen. Biometrics, 26(3):588, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529115.

Williams:1970:CUT

[136] J. S. Williams. The choice and use of tests for the independence of two sets of variates. Biometrics, 26(4):613–624, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528710.

Williams:1970:DBR

[137] D. A. Williams. Discrimination between regression models to determine the pattern of enzyme synthesis in synchronous cell cultures. Biometrics, 26(1):23–32, March 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529041.

Winston:1970:SMU

[138] Jerome A. Winston. Survival of a mutation under mixed positive assorative and random mating. i. one locus, two alleles without dominance. Biometrics, 26(3):433–450, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529100.

Wyshak:1970:NNT

[139] Grace Wyshak and Katherine Detre. 289 note: a new table for the estimation of particle density. Biometrics, 26(3):582–585, September 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529113.

Agha:1971:DMF

[140] M. Agha. A direct method for fitting linear combinations of exponentials. Biometrics, 27(2):399–413, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529004.

Allen:1971:UPI

[141] David M. Allen. The use of prior information in the testing of new drugs. Biometrics, 27(3):699–708, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528607.

Alling:1971:EHN

[142] David W. Alling. Estimation of hit number. Biometrics, 27(3):605–613, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528599.

Andrews:1971:TMD

[143] D. F. Andrews, R. Gnanadesikan, and J. L. Warner. Transformations of multivariate data. Biometrics, 27(4):825–840, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528821.

Anonymous:1971:Aa

[144] Anonymous. Abstracts. Biometrics, 27(1):243–263, March 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528952.

Anonymous:1971:Ab

[145] Anonymous. Abstracts. Biometrics, 27(2):475–493, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529034.

Anonymous:1971:Ac

[146] Anonymous. Abstracts. Biometrics, 27(3):762–773, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528623.

Anonymous:1971:Ad

[147] Anonymous. Abstracts. Biometrics, 27(4):1114, December 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528864.

Anonymous:1971:BRb

[148] Anonymous. [Books received]. Biometrics, 27(1):241–242, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528951.

Anonymous:1971:BRA

[149] Anonymous. Book review: Aplicatii ale Teoriei Informatiei. Sisteme dinamice. Sisteme cibernetice, by S. Guiasu. Biometrics, 27(2):467–468, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529021.

Anonymous:1971:BRc

[150] Anonymous. [Books received]. Biometrics, 27(2):473–474, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529033.

Anonymous:1971:BRd

[151] Anonymous. [Books received]. Biometrics, 27(3):760–761, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528622.

Anonymous:1971:BRe

[152] Anonymous. [Books received]. Biometrics, 27(4):1112–1113, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528863.

Anonymous:1971:BRWa

[153] Anonymous. Book review: World Health Statistics Annual, Volume III 1966: Health Personnel and Hospital Establishments by WHO. Biometrics, 27(2):472–473, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529030.

Anonymous:1971:BRWb

[154] Anonymous. Book review: World Health Statistics Annual, Volume I, 1967: Vital Statistics and Causes of Death, by WHO. Biometrics, 27(2): 473, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529031.

Anonymous:1971:BRWc

[155] Anonymous. Book review: World Health Statistics Annual 1967, Volume II, Infectious Diseases: Cases, Deaths and Vaccinations by WHO. Biometrics, 27(2):473, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529032.

Anonymous:1971:BSa

[156] Anonymous. The Biometric Society. Biometrics, 27(2):494–498, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529035.

Anonymous:1971:BSb

[157] Anonymous. The Biometric Society. Biometrics, 27(3):774–779, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528624.

Anonymous:1971:BSc

[158] Anonymous. The Biometric Society. Biometrics, 27(4):1115–1119, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528865.

Anonymous:1971:CIB

[159] Anonymous. Correction: Incomplete block designs for bio-assays. Biometrics, 27(4):1103, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528848. See [3].

Anonymous:1971:CND

[160] Anonymous. Correction: 286. Note: Design of experiments to estimate heritability by regression of offspring on selected parents. Biometrics, 27(4):1103, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528850. See [81].

Anonymous:1971:CNE

[161] Anonymous. Correction: 299. Note: Estimating the ratio of two marginal probabilities in a contingency table. Biometrics, 27(4):1104, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528851. See [262].

Anonymous:1971:CNT

[162] Anonymous. Correction: 300. Note: Testing for equal slopes in a randomized block design with covariance. Biometrics, 27(4):1103, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528849. See [204].

Anonymous:1971:CRM

[163] Anonymous. Correction: The removal method for two and three samples. Biometrics, 27(4):1104, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528852. See [126].

Anonymous:1971:FMa

[164] Anonymous. Front matter. Biometrics, 27(1):i–ii, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528927.

Anonymous:1971:FMb

[165] Anonymous. Front matter. Biometrics, 27(2):i–ii, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528994.

Anonymous:1971:FMc

[166] Anonymous. Front matter. Biometrics, 27(3):i–ii, September 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528591.

Anonymous:1971:FMd

[167] Anonymous. Front matter. Biometrics, 27(4):i–ii, December 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528819.

Anonymous:1971:MMN

[168] Anonymous. Mortality from malignant neoplasms 1955–1965, number of deaths by site, sex and age. parts I and II. by WHO. Biometrics, 27(2):

472, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529029.

Anonymous:1971:NAa

[169] Anonymous. News and announcements. Biometrics, 27(1):270–274, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528954.

Anonymous:1971:NAb

[170] Anonymous. News and announcements. Biometrics, 27(2):499, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529036.

Anonymous:1971:NAc

[171] Anonymous. News and announcements. Biometrics, 27(3):780–782, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528625.

Anonymous:1971:NAd

[172] Anonymous. News and announcements. Biometrics, 27(4):1120–1121, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528866.

Anonymous:1971:OHS

[173] Anonymous. Oral health surveys: Basic methods by world health organization. Biometrics, 27(4):1111, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528861.

Anonymous:1971:VI

[174] Anonymous. Volume information. Biometrics, 27(4):i–xxxv, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528818.

Arvanitis:1971:UGV

[175] L. G. Arvanitis and Biyi Afonja. Use of the generalized variance and the gradient projection method in multivariate stratified sampling. Biometrics, 27(1):119–127, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528931.

Ashton:1971:BRT

[176] Gordon C. Ashton. Book review: Topics in Intermediate Statistical Methods, by T. A. Bancroft. Biometrics, 27(1):236–237, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528947.

Banerjee:1971:BRS

[177] K. S. Banerjee. Book review: Solving Problems in Statistics, by J. Aitchison. Biometrics, 27(1):236, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528946.

Bartlett:1971:NSA

[178] M. S. Bartlett, Jennifer M. Brennan, and J. N. Pollock. 313. Note: Stochastic analysis of some experiments on the mating of blowflies. Biometrics, 27(3):725–730, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528610.

Best:1971:DFD

[179] William R. Best, Jack M. Becktel, and Arthur F. Johnson. Descriptive functions in disease. Biometrics, 27(4):895–901, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528826.

Bishop:1971:ECM

[180] Yvonne M. M. Bishop. Effects of collapsing multidimensional contingency tables. Biometrics, 27(3):545–562, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528596.

Boardman:1971:NGM

[181] Thomas J. Boardman and Donald R. Moffitt. 316. Note: Graphical Monte Carlo Type I error rates for multiple comparison procedures. Biometrics, 27(3):738–744, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528613.

Bradley:1971:ASS

[182] Ralph A. Bradley, Kantilal M. Patel, and Dennis D. Wackerly. Approximate small-sample distributions for multivariate two-sample nonparametric tests. Biometrics, 27(3):515–530, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528593. See corrections [329].

Calinski:1971:SDP

[183] T. Caliński. On some desirable patterns in block designs. Biometrics, 27(2):275–292, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528995.

Carmer:1971:BRM

[184] S. G. Carmer. Book review: The Mathematical Approach to Physiological Problems: a Critical Primer, by D. S. Riggs. Biometrics, 27(2):469– 470, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529026.

Carothers:1971:EEL

[185] A. D. Carothers. An examination and extension of Leslie’s test of equal catchability. Biometrics, 27(3):615–630, September 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528600.

Chiang:1971:SMH

[186] Chin Long Chiang. A stochastic model of human fertility. Biometrics, 27(2):345–356, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528999.

Choi:1971:IWM

[187] S. C. Choi. An investigation of Wetherill’s method of estimation for the up-and-down experiment. Biometrics, 27(4):961–970, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528831.

Cohen:1971:EIC

[188] Joel E. Cohen. Estimation and interaction in a censored 2×2×2 contingency table. Biometrics, 27(2):379–386, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529002. See corrections [330].

Dagnelie:1971:BRS

[189] P. Dagnelie. Book review: Statistical Ecology, by G. P. Patil, E. C. Pielou, W. E. Waters. Biometrics, 27(4):1106–1108, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528856.

Edwards:1971:DBP

[190] A. W. F. Edwards. Distances between populations on the basis of gene frequencies. Biometrics, 27(4):873–881, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528824.

F:1971:BRAa

[191] W. T. F. Book review: Applications of Mental Health Statistics: Uses in Mental Health Programmes of Statistics Derived from Psychiatric Services and Selected Vital and Morbidity Records, by M. Kramer. Biometrics, 27(2):468–469, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529023.

F:1971:BRAb

[192] W. T. F. Book review: Acceptance Sampling — a Symposium, by J. W. Tukey, F. Mosteller, C. P. Winsor. Biometrics, 27(4):1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528858.

F:1971:BRE

[193] W. T. F. Book review: Elementary Statistical Analysis, by S. S. Wilks. Biometrics, 27(4):1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528859.

F:1971:BRI

[194] W. T. F. Book review: Inventory Control Abstracts 1953–1965, by S. Eilon, W. Lampkin. Biometrics, 27(2):467, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529020.

F:1971:BRNa

[195] W. T. F. Book review: The National Halothane Study. A Study of the Possible Association between Halothane Anesthesia and Postoperative Hepatic Necrosis, by J. P. Bunker, W. H. Forrest, Jr., F. Mosteller, L. D. Vandam. Biometrics, 27(2):466, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529017.

F:1971:BRNb

[196] W. T. F. Book review: National Accounts Statistics: Sources and Methods, by R. Maurice. Biometrics, 27(2):469, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529024.

F:1971:BRSa

[197] W. T. F. Book review: Scientific Research in British Universities and Colleges 1969–1970. Volume II: Biological Sciences, by Department of Education and Science. Biometrics, 27(2):467, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529019.

F:1971:BRSb

[198] W. T. F. Book review: Selected Tables in Mathematical Statistics Volume I, by H. L. Harter, D. B. Owen. Biometrics, 27(2):468, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529022.

F:1971:BRSc

[199] W. T. F. Book review: Selected Statistical Papers 1 (No. 26) and 2 (No. 27), by Statistical Department of the Mathematical Centre Tracts. Biometrics, 27(2):472, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529028.

F:1971:BRSd

[200] W. T. F. Book review: Statistical Methods for Research Workers, by Ronald A. Fisher. Biometrics, 27(4):1106, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528855.

F:1971:BRSe

[201] W. T. F. Book review: Spontaneous and Induced Abortion, by World Health Organization. Biometrics, 27(4):1111, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528862.

F:1971:RAS

[202] W. T. F. Review: Acceptance Sampling — a Symposium, by J. W. Tukey, F. Mosteller, and C. P. Winsor. Biometrics, 27(4):1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528858.

Fienberg:1971:BRC

[203] S. E. Fienberg. Book review: The Chi-Squared Distribution, by H. O. Lancaster. Biometrics, 27(1):238–241, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528950.

Fleiss:1971:NTE

[204] Joseph L. Fleiss. 300. Note: Testing for equal slopes in a randomized block design with covariance. Biometrics, 27(1):225–229, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528943. See correction [162].

Fraser:1971:BRC

[205] W. Fraser. Book review: Calculus and Statistics, by M. C. Gemignani. Biometrics, 27(1):238, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528949.

Fuller:1971:TSA

[206] F. C. Fuller, Jr. and Chris P. Tsokos. Time series analysis of water pollution data. Biometrics, 27(4):1017–1034, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528835.

Gani:1971:MCM

[207] J. Gani and D. Jerwood. Markov chain methods in chain binomial epidemic models. Biometrics, 27(3):591–603, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528598.

Gokhale:1971:IPA

[208] D. V. Gokhale. An iterative procedure for analysing log-linear models. Biometrics, 27(3):681–687, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528605.

Goodchild:1971:DHF

[209] N. A. Goodchild. Designs having fruit trees as blocks and incorporating analysis of covariance. Biometrics, 27(2):313–323, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528997.

Gower:1971:GCS

[210] J. C. Gower. A general coefficient of similarity and some of its properties. Biometrics, 27(4):857–871, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528823.

Gravett:1971:NSA

[211] Isabelle M. Gravett. 306. Note: Some aspects of 2^{n} factorials with dummy treatments. Biometrics, 27(2):436–440, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529009.

Gravett:1971:TTR

[212] Isabelle M. Gravett. A two-thirds replicate of a 3^{4} factorial. Biometrics, 27(4):1043–1055, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528837. See corrections [615].

Grizzle:1971:NML

[213] James E. Grizzle. 317. Note: Multivariate logit analysis. Biometrics, 27(4):1057–1062, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528838.

Gross:1971:ESP

[214] Alan J. Gross, Virginia A. Clark, and Vinnie Liu. Estimation of survival parameters when one of two organs must function for survival. Biometrics, 27(2):369–377, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529001.

Hamdan:1971:NEC

[215] M. A. Hamdan. 311. Note: Estimation of class boundaries in fitting a normal distribution to a qualitative multinomial distribution. Biometrics, 27(2):457–459, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529014.

Hartley:1971:AID

[216] H. O. Hartley and R. R. Hocking. The analysis of incomplete data. Biometrics, 27(4):783–823, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528820.

Hedayat:1971:BRS

[217] A. Hedayat. Book review: Sequences, Combinations, Limits, by S. I. Gelfand, M. L. Gerver, A. A. Kirillov, N. N. Konstantinov, A. G. Kushnirenko. Biometrics, 27(1):237–238, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528948.

Hill:1971:DES

[218] William G. Hill. Design and efficiency of selection experiments for estimating genetic parameters. Biometrics, 27(2):293–311, June 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528996.

Hines:1971:BRS

[219] W. G. S. Hines. Book review: Statistics: an Intuitive Approach, by G. H. Winberg, J. A. Schumaker. Biometrics, 27(4):1110–1111, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528860.

Hinkelmann:1971:BRE

[220] Klaus Hinkelmann. Book review: Elementare Einführung in die mathematische Statistik, by D. Rasch. Biometrics, 27(2):469, June 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529025.

Hinkelmann:1971:BRS

[221] K. Hinkelmann. Book review: Statistische Auswertungsmethoden, by L. Sachs. Biometrics, 27(2):470–471, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529027.

Hinkelmann:1971:EHE

[222] Klaus Hinkelmann. Estimation of heritability from experiments with inbred and related individuals. Biometrics, 27(1):183–190, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528936. See corrections [1421].

Hoel:1971:SFT

[223] David G. Hoel and Toby J. Mitchell. The simulation, fitting, and testing of a stochastic cellular proliferation model. Biometrics, 27(1):191–199, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528937.

Holgate:1971:NDR

[224] P. Holgate. 309. Note: Drift in the random component of isonymy. Biometrics, 27(2):448–451, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529012.

Huehn:1971:NRM

[225] M. Huehn. 322. Note: A rapid method for polynomial regression with equally spaced observations. Biometrics, 27(4):1083–1086, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528843.

Hyams:1971:PPB

[226] Lyon Hyams. The practical psychology of biostatistical consultation. Biometrics, 27(1):201–211, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528938.

Jaech:1971:NFT

[227] John L. Jaech. 325. Note: Further tests of significance for Grubb’s estimators. Biometrics, 27(4):1097–1101, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528846.

Johnson:1971:NDE

[228] A. F. Johnson. 314. Note: Design of experiments: The single cycle sine wave model. Biometrics, 27(3):730–735, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528611.

Jolicoeur:1971:AEA

[229] Pierre Jolicoeur and Alfred A. Heusner. The allometry equation in the analysis of the standard oxygen consumption and body weight of the white rat. Biometrics, 27(4):841–855, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528822.

Kempthorne:1971:BRE

[230] Oscar Kempthorne. Book review: Evolution and the Genetics of Populations. Volume 2: The Theory of Gene Frequencies, by Sewall Wright. Biometrics, 27(3):758–760, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528621.

Kerr:1971:NPD

[231] J. D. Kerr. 298. Note: The probability of disease transmission. Biometrics, 27(1):219–222, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528941. See corrections [331].

Kimball:1971:MVM

[232] A. W. Kimball. Model I vs. model II in competing risk theory. Biometrics, 27(2):462–465, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529016.

Klauber:1971:TSR

[233] Melville R. Klauber. Two-sample randomization tests for space-time clustering. Biometrics, 27(1):129–142, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528932.

Kneale:1971:PAE

[234] G. W. Kneale. Problems arising in estimating from retrospective survey data the latent periods of juvenile cancers initiated by obstetric radiography. Biometrics, 27(3):563–590, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528597.

Koch:1971:ACD

[235] Gary G. Koch and Donald W. Reinfurt. The analysis of categorical data from mixed models. Biometrics, 27(1):157–173, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528934.

Kowal:1971:NDG

[236] Robert R. Kowal. 296. Note: Disadvantages of the generalized variance as a measure of variability. Biometrics, 27(1):213–216, March 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528939.

Krzanowski:1971:NCS

[237] W. J. Krzanowski. 318. Note: A comparison of some distance measures applicable to multinomial data, using a rotational fit technique. Biometrics, 27(4):1062–1068, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528839.

Ku:1971:AIA

[238] H. H. Ku. Analysis of information — an alternative approach to the detection of a correlation between the sexes of adjacent sibs in human families. Biometrics, 27(1):175–182, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528935.

Kuehl:1971:OHT

[239] Robert O. Kuehl. Obituary: Henry Tucker 1923–1971. Biometrics, 27(4): 1102, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528847.

Kullback:1971:NHS

[240] S. Kullback. 310. Note: The homogeneity of the sex ratio of adjacent sibs in human families. Biometrics, 27(2):452–457, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529013.

Levin:1971:BRI

[241] Simon A. Levin and Daniel L. Solomon. Book review: An Introduction to Mathematical Ecology, by E. C. Pielou. Biometrics, 27(3):751–754, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528618.

Malik:1971:BRA

[242] Henrick J. Malik. Book review: Applied General Statistics, by F. E. Croxton, D. J. Cowden, S. Klein. Biometrics, 27(2):466–467, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529018.

Manly:1971:SSJ

[243] B. F. J. Manly. A simulation study of Jolly’s method for analysing capture–recapture data. Biometrics, 27(2):415–424, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529005.

Mann:1971:MMC

[244] Stuart H. Mann. Mathematical models for the control of pest populations. Biometrics, 27(2):357–368, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529000.

Mantel:1971:CCC

[245] Nathan Mantel. The continuity correction in comparing two Poissondistributed observations. Biometrics, 27(3):747–749, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528615.

Mantel:1971:NGT

[246] Nathan Mantel. 308. Note: On Green’s test for dependence. Biometrics, 27(2):444–447, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529011.

Mantel:1971:SES

[247] Nathan Mantel and R. C. Elston. Supplements to the elston-stewart test of association. Biometrics, 27(2):460–462, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529015.

Marriott:1971:NBP

[248] F. H. C. Marriott. 302. Note: Buffon’s problem for non-random directions. Biometrics, 27(1):233–235, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528945.

Marriott:1971:PPM

[249] F. H. C. Marriott. Practical problems in a method of cluster analysis. Biometrics, 27(3):501–514, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528592.

Mason:1971:NCD

[250] J. M. Mason and K. Hinkelmann. 305. Note: Change-over designs for testing different treatment factors at several levels. Biometrics, 27(2): 430–435, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529008.

Matis:1971:SCA

[251] J. H. Matis and H. O. Hartley. Stochastic compartmental analysis: Model and least squares estimation from time series data. Biometrics, 27(1):77– 102, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528929.

Mayer:1971:MCA

[252] Lawrence S. Mayer. A method of cluster analysis when there exist multiple indicators of a theoretic concept. Biometrics, 27(1):143–155, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528933.

McDonald:1971:EMD

[253] Lyman McDonald. On the estimation of missing data in the multivariate linear model. Biometrics, 27(3):535–543, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528595.

McGilchrist:1971:RAP

[254] C. A. McGilchrist and B. R. Trenbath. A revised analysis of plant competition experiments. Biometrics, 27(3):659–671, September 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528603.

McNally:1971:NMM

[255] D. H. McNally. 315. Note: Mathematical model for poultry egg production. Biometrics, 27(3):735–738, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528612.

Milliken:1971:NTI

[256] George A. Milliken and Franklin A. Graybill. 321. Note: Tests for interaction in the two-way model with missing data. Biometrics, 27(4):1079– 1083, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528842.

Moeschberger:1971:LTU

[257] M. L. Moeschberger and H. A. David. Life tests under competing causes of failure and the theory of competing risks. Biometrics, 27(4):909–933, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528828.

Nam:1971:TTC

[258] Jun-Mo Nam. On two tests for comparing matched proportions. Biometrics, 27(4):945–959, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528830.

Nelder:1971:TVC

[259] J. A. Nelder. [Topics in variance component estimation]. Biometrics, 27(3):750, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528616.

Peacock:1971:NCR

[260] P. B. Peacock. The non-comparability of relative risks from different studies. Biometrics, 27(4):903–907, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528827.

Pearce:1971:QTC

[261] S. C. Pearce. 303. query: Tests not catered for by the design of an experiment. Biometrics, 27(2):425–426, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529006.

Peritz:1971:NER

[262] Eric Peritz. 299. Note: Estimating the ratio of two marginal probabilities in a contingency table. Biometrics, 27(1):223–225, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528942. See correction [161].

Pollard:1971:DED

[263] J. H. Pollard. On distance estimators of density in randomly distributed forests. Biometrics, 27(4):991–1002, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528833.

Preece:1971:NSN

[264] D. A. Preece. 304. Note: Some new balanced row-and-column designs for two non-interacting sets of treatments. Biometrics, 27(2):426–430, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529007.

Puri:1971:BRI

[265] S. C. Puri. Book review: Introduction to Statistics, by J. M. Bevan. Biometrics, 27(3):751, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528617.

Puri:1971:BRM

[266] S. C. Puri. Book review: Multicompartment Models for Biological Systems, by G. L. Atkins. Biometrics, 27(4):1105, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528853.

Rao:1971:CIE

[267] J. N. K. Rao and Kathleen Subrahmaniam. Combining independent estimators and estimation in linear regression with unequal variances. Biometrics, 27(4):971–990, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528832.

Rojas:1971:BRF

[268] Basilio A. Rojas. Book review: A First Course in Biometry for Agriculture Students, by A. A. Rayner. Biometrics, 27(4):1108–1109, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528857.

Rosenberg:1971:NRP

[269] Saul H. Rosenberg and Charles A. Rohde. 324. Note: Robustness properties of variance component estimators when the underlying model is incorrectly specified. Biometrics, 27(4):1091–1097, December 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528845.

Roux:1971:BRG

[270] C. Z. Roux. Book review: Genetic Load: Its Biological and Conceptual Aspects, by Bruce Wallace. Biometrics, 27(3):755–757, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528620.

Royall:1971:IAS

[271] B. W. Royall. International aspects of the study of adverse reactions to drugs. Biometrics, 27(3):689–698, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528606.

Rudan:1971:NLS

[272] J. W. Rudan and S. R. Searle. 323. Note: Large sample variances of maximum likelihood estimators of variance components in the 3-way nested classification, random model, with unbalanced data. Biometrics, 27(4):1087–1091, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528844.

Salsburg:1971:TDR

[273] David Salsburg. Testing dose responses on proportions near zero or one with the jackknife. Biometrics, 27(4):1035–1041, December 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528836.

Sanders:1971:QGC

[274] J. L. Sanders. Quantitative guidelines for communicable disease control programs. Biometrics, 27(4):883–893, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528825.

Schneider:1971:BS

[275] B. Schneider and H. Thoni. The Biometric Society. Biometrics, 27(1): 264–269, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528953.

Scott:1971:CMB

[276] A. J. Scott and M. J. Symons. Clustering methods based on likelihood ratio criteria. Biometrics, 27(2):387–397, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529003.

Scott:1971:NEC

[277] A. J. Scott and Michael J. Symons. 297. Note: On the Edwards and Cavalli–Sforza method of cluster analysis. Biometrics, 27(1):217–219, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528940.

Searle:1971:BIP

[278] S. R. Searle. A biometrics invited paper. topics in variance component estimation. Biometrics, 27(1):1–76, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528928.

Searle:1971:BRG

[279] S. R. Searle. Book review: Generalized Inverse Matrices, by T. L. Boullion, P. L. Odell. Biometrics, 27(4):1105–1106, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528854.

Searle:1971:BRM

[280] S. R. Searle. Book review: Models in Regression and Related Topics, by Peter Sprent. Biometrics, 27(3):754–755, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528619.

Sedransk:1971:PSE

[281] J. Sedransk. Precision specifications for the estimation of treatment differences. Biometrics, 27(3):673–680, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528604.

Sheps:1971:MSB

[282] Mindel C. Sheps and Jane A. Menken. A model for studying birth rates given time dependent changes in reproductive parameters. Biometrics, 27(2):325–343, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528998.

Siskind:1971:NMC

[283] Victor Siskind. 319. Note: Moments of the criterion of a test for association. Biometrics, 27(4):1069–1074, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528840.

Sprent:1971:NPC

[284] P. Sprent. 307. Note: Parallelism and concurrence in linear regression. Biometrics, 27(2):440–444, June 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529010.

Strauss:1971:QAD

[285] D. J. Strauss. 312. query: Analysis of drug effects on body temperature. Biometrics, 27(3):721–724, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528609.

Tenenbein:1971:DSS

[286] Aaron Tenenbein. A double sampling scheme for estimating from binomial data with misclassifications: Sample size determination. Biometrics, 27(4):935–944, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528829.

Tomlinson:1971:SSP

[287] Patrick K. Tomlinson. Some sampling problems in fishery work. Biometrics, 27(3):631–641, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528601.

Townsend:1971:BQU

[288] E. C. Townsend and S. R. Searle. Best quadratic unbiased estimation of variance components from unbalanced data in the 1-way classification. Biometrics, 27(3):643–657, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528602.

Tyagi:1971:NCA

[289] B. N. Tyagi. 301. Note: Confounded asymmetrical factorial designs. Biometrics, 27(1):229–232, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528944.

Urbakh:1971:LDA

[290] V. Yu. Urbakh. Linear discriminant analysis: Loss of discriminating power when a variate is omitted. Biometrics, 27(3):531–534, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528594.

vanBelle:1971:NST

[291] Gerald van Belle and Richard G. Cornell. 320. Note: Strengthening tests of symmetry in contingency tables. Biometrics, 27(4):1074–1078, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528841.

Vithayasai:1971:RHD

[292] C. Vithayasai. Runs of healthy and diseased trees in transects through an infected forest: Statistical and computational methods. Biometrics, 27(4):1003–1016, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528834.

Weiling:1971:NES

[293] F. Weiling. Neue ergebnisse zur statistischen vorgeschichte der mendelschen versuche. Biometrics, 27(3):709–719, September 1971. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528608.

Williams:1971:TDB

[294] D. A. Williams. A test for differences between treatment means when several dose levels are compared with a zero dose control. Biometrics, 27(1): 103–117, March 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528930. See

corrections [730].

Zelen:1971:OWJ

[295] Marvin Zelen. Obituary: W. J. Youden 1900–1971. Biometrics, 27 (3):745–746, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528614.

Andrews:1972:PHD

[296] D. F. Andrews. Plots of high-dimensional data. Biometrics, 28(1):125– 136, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528964.

Anonymous:1972:Aa

[297] Anonymous. Abstracts. Biometrics, 28(1):258–275, March 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528990.

Anonymous:1972:Ab

[298] Anonymous. Abstracts. Biometrics, 28(3):899–912, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528783.

Anonymous:1972:Ac

[299] Anonymous. Abstracts. Biometrics, 28(4):1157–1185, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528667.

Anonymous:1972:BRa

[300] Anonymous. [Books received]. Biometrics, 28(1):256–257, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528989.

Anonymous:1972:BRB

[301] Anonymous. Book review: Biometrical Genetics, by K. Mather, J. L. Jinks. Biometrics, 28(2):638–639, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556195.

Anonymous:1972:BRSa

[302] Anonymous. Book review: Statistical Papers of George Udny Yule by A. Stuart, M. G. Kendall. Biometrics, 28(2):645, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556204.

Anonymous:1972:BRC

[303] Anonymous. Book review: Carl Friedrich Gauss. A Biography, by T. Hall, Albert Forderberg. Biometrics, 28(1):248–249, March 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528976.

Anonymous:1972:BRE

[304] Anonymous. Book review: Explorations in Mathematical Anthropology by P. Kay. Biometrics, 28(2):636–637, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556193.

Anonymous:1972:BRFa

[305] Anonymous. Book review: Factorial Arrangements and Analysis of Uniformity Trials, with Reference to Efficiency and Non-Additivity in Some Blocking Design Problems by M. Mitome. Biometrics, 28(2):640, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556197.

Anonymous:1972:BRFb

[306] Anonymous. Book review: Federal Statistics, by W. A. Wallis, F. Mosteller, et al. Biometrics, 28(2):646–647, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556207.

Anonymous:1972:BRGa

[307] Anonymous. Book review: General Least Squares Analysis of Diallel Experiments. A Computer Program — DIALL by H. Schaffer, R. A. Usanis. Biometrics, 28(1):255, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528986.

Anonymous:1972:BRGb

[308] Anonymous. Book review: Genetics, Evolution, Race, Radiation Biology: Essays in Social Biology, Volume II, by B. Wallace. Biometrics, 28(2): 646, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556206.

Anonymous:1972:BRI

[309] Anonymous. Book review: Informal Introduction to ALGOL 68, by C. H. Lindsey, S. G. Van Der Muelen. Biometrics, 28(1):250–253, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528981.

Anonymous:1972:BRM

[310] Anonymous. Book review: Mathematics: a Modern Introduction, by M. L. Keedy, M. L. Bittinger. Biometrics, 28(2):638, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556194.

Anonymous:1972:BRN

[311] Anonymous. Book review: Nonparametric Techniques in Statistical Inference by M. L. Puri. Biometrics, 28(2):641–642, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556199.

Anonymous:1972:BRO

[312] Anonymous. Book review: Order Statistics by H. A. David. Biometrics, 28(2):633, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556187.

Anonymous:1972:BRPa

[313] Anonymous. Book review: A Primer of Population Biology by E. O. Wilson, W. H. Bossert. Biometrics, 28(1):255, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528987.

Anonymous:1972:BRPb

[314] Anonymous. Book review: Planning and Analysis of Clinical Studies by W. J. Burdette, E. A. Gehan. Biometrics, 28(2):631–632, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556184.

Anonymous:1972:BRPc

[315] Anonymous. Book review: Physics for Biology and Pre-Medical Students, by D. M. Burns, S. G. G. MacDonald. Biometrics, 28(2):632, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556185.

Anonymous:1972:BRPd

[316] Anonymous. Book review: Preservation of Infinite Divisibility under Mixing and Related Topics, by F. W. Steutel. Biometrics, 28(2):643– 644, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556202.

Anonymous:1972:BRPe

[317] Anonymous. Book review: People, Their Needs, Environment, Ecology: Essays in Social Biology, Volume I, by B. Wallace. Biometrics, 28(2): 645–646, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556205.

Anonymous:1972:BRd

[318] Anonymous. [Books received]. Biometrics, 28(2):649–651, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556210.

Anonymous:1972:BRR

[319] Anonymous. Book review: Rank Order Probabilities. Two-Sampl Normal Shift Alternatives, by R. C. Milton. Biometrics, 28(2):639–640, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556196.

Anonymous:1972:BRSb

[320] Anonymous. Book review: The Statistical Argument by R. M. Cormack. Biometrics, 28(1):247, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528974.

Anonymous:1972:BRSc

[321] Anonymous. Book review: Statistical Methods in Medical Research by P. Armitage. Biometrics, 28(2):629–630, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556182.

Anonymous:1972:BRSd

[322] Anonymous. Book review: Statistical Indicators for the Planning and Evaluation of Public Health Programmes by Who Expert Committee on Health Statistics. Biometrics, 28(2):649, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556209.

Anonymous:1972:BRTa

[323] Anonymous. Book review: Tables of the Incomplete Beta-Function, by E. S. Pearson, N. L. Johnson. Biometrics, 28(2):641, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556198.

Anonymous:1972:BRTb

[324] Anonymous. Book review: The Teaching of Probability and Statistics by L. Råde. Biometrics, 28(2):642–643, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556200.

Anonymous:1972:BSa

[325] Anonymous. The Biometric Society. Biometrics, 28(1):276–282, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528991.

Anonymous:1972:BSb

[326] Anonymous. The Biometric Society. Biometrics, 28(2):652–659, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556211.

Anonymous:1972:BSc

[327] Anonymous. The Biometric Society. Biometrics, 28(3):913–915, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528784.

Anonymous:1972:BSd

[328] Anonymous. The Biometric Society. Biometrics, 28(4):1186–1190, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528668.

Anonymous:1972:CAS

[329] Anonymous. Corrections: Approximate small-sample distributions for multivariate two-sample nonparametric tests. Biometrics, 28(4):1141, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528648. See [182].

Anonymous:1972:CEI

[330] Anonymous. Corrections: Estimation and interaction in a censored 2 × 2×2 contingency table. Biometrics, 28(4):1141, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528649. See [188].

Anonymous:1972:CPD

[331] Anonymous. Corrections: The probability of disease transmission. Biometrics, 28(4):1142, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528651. See [231].

Anonymous:1972:CRE

[332] Anonymous. Corrections: Robust estimates, residuals, and outlier detection with multiresponse data. Biometrics, 28(4):1142, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528650. See [398].

Anonymous:1972:FMa

[333] Anonymous. Front matter. Biometrics, 28(1):i–ii, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528956.

Anonymous:1972:FMb

[334] Anonymous. Front matter. Biometrics, 28(2):i–ii, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556147.

Anonymous:1972:FMc

[335] Anonymous. Front matter. Biometrics, 28(3):i–ii, September 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528754.

Anonymous:1972:FMd

[336] Anonymous. Front matter. Biometrics, 28(4):i–ii, December 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528628.

Anonymous:1972:NAa

[337] Anonymous. News and announcements. Biometrics, 28(1):283–284, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528992.

Anonymous:1972:NAb

[338] Anonymous. News and announcements. Biometrics, 28(2):660–662, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556212.

Anonymous:1972:NAc

[339] Anonymous. News and announcements. Biometrics, 28(3):916–919, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528785.

Anonymous:1972:NAd

[340] Anonymous. News and announcements. Biometrics, 28(4):1191–1194, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528669.

Anonymous:1972:BRSe

[341] Anonymous. Book review: Solving Problems in Statistics II, by J. Aitchison. Biometrics, 28(3):892, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528774.

Anonymous:1972:VI

[342] Anonymous. Volume information. Biometrics, 28(4):xiv–lxi, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528627.

Armitage:1972:F

[343] P. Armitage. Foreword. Biometrics, 28(1):1, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528957.

Ashford:1972:QRA

[344] J. R. Ashford and P. J. Walker. Quantal response analysis for a mixture of populations. Biometrics, 28(4):981–988, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528634.

Azen:1972:ASS

[345] S. P. Azen and A. A. Afifi. Asymptotic and small-sample behavior of estimated Bayes rules for classifying time-dependent observations. Biometrics, 28(4):989–998, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528635.

Battese:1972:DEO

[346] George E. Battese and Wayne A. Fuller. Determination of economic optima from crop-rotation experiments. Biometrics, 28(3):781–792, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528762.

Beaver:1972:BRPa

[347] Robert J. Beaver. Book reviews: Probability and Statistics, by J. R. Blum, J. I. Rosenblatt; Solutions Manual for Probability and Statistics, by J. R. Blum, J. I. Rosenblatt. Biometrics, 28(4):1146, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528655.

Beaver:1972:BRPb

[348] Robert J. Beaver. Book review: Probability Distribution: an Introduction to Probability Theory with Applications, by Chris P. Tsokos. Biometrics, 28(4):1156, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528666.

Behnken:1972:BED

[349] Donald W. Behnken and Donald G. Watts. Bayesian estimation and design of experiments for growth rates when sampling from the Poisson distribution. Biometrics, 28(4):999–1009, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528636.

Bennett:1972:CSS

[350] B. M. Bennett. On comparisons of sensitivity, specificity and predictive value of a number of diagnostic procedures. Biometrics, 28(3):793–800, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528763. See

correction [484].

Bennett:1972:NMC

[351] B. M. Bennett. 334. Note: Measures for clinicians’ disagreements over signs. Biometrics, 28(2):607–612, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556176.

Berkson:1972:MDI

[352] Joseph Berkson. Minimum discrimination information, the ‘no interaction’ problem, and the logistic function. Biometrics, 28(2):443–468, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2556159.

Brunden:1972:NAN

[353] Marshall N. Brunden. 333. Note: The analysis of non-independent 2 X 2 tables and 2 X 2 C tables using rank sums. Biometrics, 28(2): 603–607, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556175. See

correction [1420].

Burrows:1972:ESD

[354] Peter M. Burrows. Expected selection differentials for directional selection. Biometrics, 28(4):1091–1100, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528642.

Cady:1972:BRF

[355] Foster B. Cady. Book review: Foundations of Statistical Inference, by V. P. Godambe, D. A. Sprott. Biometrics, 28(4):1149–1150, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528659.

Chang:1972:NSD

[356] Potter C. Chang and Eugene A. Johnson. 341. Note: Some distributionfree properties of the asymptotic variance of the Spearman estimator in bioassays. Biometrics, 28(3):882–889, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528772.

Choi:1972:NTH

[357] S. C. Choi and R. Wette. 330. Note: A test for the homogeneity of variances among correlated variables. Biometrics, 28(2):589–592, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556172.

Conniffe:1972:DSR

[358] D. Conniffe and M. A. Moran. Double sampling with regression in comparative studies of carcass composition. Biometrics, 28(4):1011–1023, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528637.

Cook:1972:DFP

[359] R. D. Cook and R. F. Nassar. Dynamics of finite populations. I. The expected time to fixation or loss and the probability of fixation of an allele in a haploid population of variable size. Biometrics, 28(2):373–384, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556154. See comments [597].

Cormack:1972:LCR

[360] R. M. Cormack. The logic of capture-recapture estimates. Biometrics, 28(2):337–343, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556151.

Cox:1972:BRS

[361] C. Philip Cox. Book review: Sequential Methods in Statistics, by G. B. Wetherill. Biometrics, 28(2):647–649, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556208.

Cox:1972:BSF

[362] Gertrude M. Cox. The Biometric Society: The first twenty-five years (1947–1972). Biometrics, 28(2):285–311, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556148.

Cox:1972:NER

[363] C. Philip Cox. 340. Note: On estimating relative potency from quadratic log-dose response relationships. Biometrics, 28(3):875–881, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528771.

Curnow:1972:MMI

[364] R. N. Curnow. The multifactorial model for the inheritance of liability to disease and its implications for relatives at risk. Biometrics, 28(4):931– 946, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528630.

N:1972:BRE

[365] F. N. D. Book review: Elements of Biometry, by K. Mather. Biometrics, 28(4):1151, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528661.

Darroch:1972:BRS

[366] J. G. Darroch. Book review: Statistical Methods, by G. W. Snedecor, W. M. Cochran. Biometrics, 28(4):1153–1155, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528664.

Dell:1972:RSS

[367] T. R. Dell and J. L. Clutter. Ranked set sampling theory with order statistics background. Biometrics, 28(2):545–555, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556166.

Dempster:1972:CS

[368] A. P. Dempster. Covariance selection. Biometrics, 28(1):157–175, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528966.

F:1972:BRAa

[369] W. T. F. Book review: Advanced Statistical Methods in Biometric Research, by C. R. Rao. Biometrics, 28(1):253–254, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528984.

F:1972:BRAb

[370] W. T. F. Book review: Abridged Tables for Use of Examination Candidates, by The Institute of Statisticians. Biometrics, 28(2):635, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556191.

F:1972:BRCa

[371] W. T. F. Book review: Curves for the Rapid Estimation of Partial Correlation Coefficients, by J. L. Ladell, M. J. Dillon. Biometrics, 28(1): 250, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528980.

F:1972:BRCb

[372] W. T. F. Book review: Characteristic Functions, by E. Lukacs. Biometrics, 28(1):253, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528982.

F:1972:BRD

[373] W. T. F. Book review: A Dictionary of Statistical Terms, by M. G. Kendall, W. R. Buckland. Biometrics, 28(1):249, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528978.

F:1972:BRE

[374] W. T. F. Book review: Experimental Design. Selected Papers, by F. Yates. Biometrics, 28(1):256, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528988.

F:1972:BRMa

[375] W. T. F. Book review: Mobility of Ph.D.’s before and after the Doctorate (with Associated Economic and Educational Characteristics of States), by L. R. Harmon. Biometrics, 28(1):249, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528977.

F:1972:BRMb

[376] W. T. F. Book review: Mathematical Topics in Population Genetics, by K. Kojima. Biometrics, 28(1):249–250, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528979.

F:1972:BRMc

[377] W. T. F. Book review: A Manual for Surveys of Fertility and Family Planning: Knowledge, Attitudes, and Practice, by Demographic Division, the Population Council. Biometrics, 28(2):634, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556189.

F:1972:BRP

[378] W. T. F. Book review: Probability Tables for Locating Elliptical Targets with Square, Rectangular, and Hexagonal Point Nets, by D. A. Singer, F. E. Wickman. Biometrics, 28(2):643, June 1972. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556201.

F:1972:BRSa

[379] W. T. F. Book review: On Some Statistical Problems Connected with the Identification of Chromosomes, by B. Matérn, M. Simak. Biometrics, 28 (1):253, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528983.

F:1972:BRSb

[380] W. T. F. Book review: Statistical Papers in Honor of George W. Snedecor, by T. A. Bancroft. Biometrics, 28(2):630–631, June 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556183.

F:1972:BRSc

[381] W. T. F. Book review: Statistical Quality Control Studies and Findings Relative to Highway Construction: an Annotated Bibliography, by J. G. Darroch. Biometrics, 28(2):633, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556186.

F:1972:BRSd

[382] W. T. F. Book review: Statforms. Formats for Elementary Statistical Calculations, by H. C. Dawkins. Biometrics, 28(2):633–634, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556188.

F:1972:BRSe

[383] W. T. F. Book review: Statistical Tolerance Regions: Classical and Bayesian, by I. Guttman. Biometrics, 28(2):634–635, June 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556190.

F:1972:BRT

[384] W. T. F. Book review: The Theory of Disinfectant Testing with Mathematical & Statistical Section, by A. H. Dodd. Biometrics, 28(1):247–248, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528975.

F:1972:BRU

[385] W. T. F. Book review: United States Life Tables by Causes of Death: 1959-61, by F. Bayo. Biometrics, 28(1):247, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528973.

F:1972:BRV

[386] W. T. F. Book review: Vortäge der II. Ungarischen Biometrischen Konferenz, by I. Juvancz. Biometrics, 28(2):635–636, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556192.

Faulkner:1972:BOC

[387] J. E. Faulkner and R. B. McHugh. Bias in observable cancer age and life-time of mice subject to spontaneous mammary carcinomas. Biometrics, 28(2):489–498, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556162.

Federer:1972:BRC

[388] W. T. Federer. Book review: Cyclic Designs, by J. A. John, F. W. Wolock, H. A. David, J. M. Cameron, J. A. Speckman. Biometrics, 28(3):894–895, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528778.

Federer:1972:BRSa

[389] W. T. Federer. Book review: Stochastic Approximation and Nonlinear Regression, by A. E Albert, L. A. Gardner, Jr. Biometrics, 28(2):629, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2556181.

Federer:1972:BRSb

[390] W. T. Federer. Book review: Statistics: a Guide to the Unknown, by J. Tanur, F. Mosteller, W. H. Kruskal, R. F. Pieters, R. S. Rising, G. R. Rising. Biometrics, 28(4):1155, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528665.

Feldman:1972:RDT

[391] Sydney Feldman, Donald F. Klein, and Gilbert Honigfeld. The reliability of a decision tree technique applied to psychiatric diagnosis. Biometrics, 28(3):831–840, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528766.

Fienberg:1972:AIM

[392] Stephen E. Fienberg. The analysis of incomplete multi-way contingency tables. Biometrics, 28(1):177–202, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528967. See correction [482].

Gail:1972:MQI

[393] M. H. Gail. Mixed quasi-independent models for categorical data. Biometrics, 28(3):703–712, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528757. See corrections [608].

Gart:1972:SAC

[394] John J. Gart. The statistical analysis of chain-binomial epidemic models with several kinds of susceptibles. Biometrics, 28(4):921–930, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528629.

Gates:1972:EDM

[395] Charles E. Gates and William B. Smith. Estimation of density of mourning doves from aural information. Biometrics, 28(2):345–359, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556152.

Gehan:1972:RPI

[396] Edmund A. Gehan and G. W. Kneale. Relationship of prenatal irradiation to death from a malignant disease. Biometrics, 28(1):239–245, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528971.

Gerrard:1972:IBS

[397] Douglas J. Gerrard and R. D. Cook. Inverse binomial sampling as a basis for estimating negative binomial population densities. Biometrics, 28(4):971–980, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528633.

Gnanadesikan:1972:RER

[398] R. Gnanadesikan and J. R. Kettenring. Robust estimates, residuals, and outlier detection with multiresponse data. Biometrics, 28(1):81– 124, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528963. See

corrections [332].

Gokhale:1972:BRP

[399] D. V. Gokhale. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Vol. I: Theory of Statistics, by L. L. Cam, J. Neyman, E. L. Scott. Biometrics, 28(4):1147–1148, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528657.

Goldstein:1972:ARP

[400] H. Goldstein. The allocation of resources in population screening: a decision theory model. Biometrics, 28(2):499–518, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556163.

Good:1972:NCP

[401] I. J. Good. 342. Note: Correlation for power functions. Biometrics, 28(4):1127–1129, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528645. See correction, ibid. 29 (1973), 829.

Gordon:1972:VHR

[402] I. L. Gordon, D. E. Byth, and L. N. Balam. Variance of heritability ratios estimated from phenotypic variance components. Biometrics, 28(2): 401–415, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556156. See

correction [485].

Gradijan:1972:QCR

[403] Jack R. Gradijan and Per-Erik E. Bergner. Qualitative consequences of randomness in a linear kinetic system. Biometrics, 28(2):313–328, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2556149.

Griffiths:1972:NFN

[404] D. A. Griffiths. 344. Note: A further note on the probability of disease transmission. Biometrics, 28(4):1133–1139, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528647.

Grizzle:1972:LLM

[405] James E. Grizzle and O. Dale Williams. Log linear models and tests of independence for contingency tables. Biometrics, 28(1):137–156, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528965.

Hill:1972:ERHa

[406] William G. Hill. Estimation of realised heritabilities from selection experiments. I. Divergent selection. Biometrics, 28(3):747–765, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528760.

Hill:1972:ERHb

[407] William G. Hill. Estimation of realised heritabilities from selection experiments. II. selection in one direction. Biometrics, 28(3):767–780, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528761.

Hodgson:1972:BRS

[408] Vincent Hodgson. Book review: The Single Server Queue, by J. W. Cohen. Biometrics, 28(3):893–894, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528776.

Hoel:1972:RMD

[409] David G. Hoel. A representation of mortality data by competing risks. Biometrics, 28(2):475–488, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2556161.

Housworth:1972:HPP

[410] William J. Housworth. Hybrid polynomial and periodic regression with and without missing observations. Biometrics, 28(4):1025–1042, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528638.

Howland:1972:BRC

[411] H. C. Howland. Book review: Computer Programming in Quantitative Biology, by R. G. Davies. Biometrics, 28(3):894, September 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528777.

Jacquard:1972:GIG

[412] Albert Jacquard. Genetic information given by a relative. Biometrics, 28(4):1101–1114, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528643.

Jenden:1972:MAA

[413] Donald J. Jenden, M. David Fairchild, Max R. Mickey, Robert W. Silverman, and Coralee Yale. A multivariate approach to the analysis of drug effects on the electroencephalogram. Biometrics, 28(1):73–80, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528962.

Jensen:1972:SSM

[414] D. R. Jensen. Some simultaneous multivariate procedures using Hotelling’s T^{2} statistics. Biometrics, 28(1):39–53, March 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528960.

Keiding:1972:MLE

[415] Niels Keiding and S. Tolver Jensen. Maximum likelihood estimation of the size distribution of liver cell nuclei from the observed distribution in a plane section. Biometrics, 28(3):813–829, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528765.

Klotz:1972:BRN

[416] J. Klotz. Book review: Nonparametric Statistical Inference, by J. D. Gibbons. Biometrics, 28(4):1148–1149, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528658.

Koch:1972:LMA

[417] Gary G. Koch, Peter B. Imrey, and Donald W. Reinfurt. Linear model analysis of categorical data with incomplete response vectors. Biometrics, 28(3):663–692, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528755.

Koch:1972:NUN

[418] Gary G. Koch. 328. Note: The use of none-parametric methods in the statistical analysis of the two-period change-over design. Biometrics, 28(2): 577–584, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556170.

Kulshreshtha:1972:NNI

[419] A. C. Kulshreshtha. 329. Note: A new incomplete block design for sloperatio assays. Biometrics, 28(2):585–589, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556171.

Lang:1972:CNL

[420] A. R. C. Lang. Correction note: Lengths and diameters of plant roots in non-random populations by analysis of plane surfaces: a correction. Biometrics, 28(2):625–627, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2556180. See [95].

Levin:1972:BRD

[421] S. A. Levin. Book review: Dynamical System Theory in Biology. Vol. 1. Stability Theory and Its Applications, by R. Rosen. Biometrics, 28 (4):1152–1153, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528663.

Levin:1972:BRI

[422] Simon Levin. Book review: Introduction to Mathematics for Life Scientists, by E. Batschelet. Biometrics, 28(3):892–893, September 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528775.

Lin:1972:NAD

[423] Chuang-Sheng Lin. 335. Note: Analysis of diallel crosses between two groups where parental lines are included. Biometrics, 28(2):612–618, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2556177.

Loewenson:1972:RMS

[424] Ruth B. Loewenson, Jacob E. Bearman, and Joseph A. Resch. Reliability of measurements for studies of cerebrov vascular atherosclerosis. Biometrics, 28(2):557–569, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556167.

Mallard:1972:TCI

[425] J. Mallard. La théorie et le calcul des index de sélection avec restriction: Synthèse critique. (French) [the theory and calculation of restricted selection indexes: Critical synthesis]. Biometrics, 28(3):713–735, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528758.

Manly:1972:ESV

[426] B. F. J. Manly. Estimating selective values from field data. Biometrics, 28(4):1115–1125, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528644.

Mantel:1972:QDS

[427] Nathan Mantel. 326. query: Double sampling and the null-hypothesis. Biometrics, 28(2):571–573, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2556168.

Maritz:1972:MHC

[428] J. S. Maritz, E. R. Stanley, G. F. Yeo, and D. Metcalf. A model of haemopoietic cell colony formation. Biometrics, 28(3):801–811, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528764.

Marriott:1972:NBP

[429] F. H. C. Marriott. 337. Note: Buffon’s problems for non-random distribution. Biometrics, 28(2):621–624, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556179.

Marriott:1972:NDD

[430] F. H. C. Marriott. 339. Note: The distance distributions of a random point from randomly spaced lines. Biometrics, 28(3):874–875, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528770.

Martin:1972:PME

[431] Donald C. Martin and Ralph A. Bradley. Probability models, estimation, and classification for multivariate dichotomous populations. Biometrics, 28(1):203–221, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528968.

Matis:1972:NGT

[432] J. M. Matis. 332. Note: Gamma time-dependency in Blaxter’s compartmental model. Biometrics, 28(2):597–602, June 1972. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2556174.

McIntyre:1972:HVT

[433] G. A. McIntyre. Homogeneity of variance in a two-way classification. Biometrics, 28(1):246, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528972.

Mendelson:1972:BRI

[434] E. Mendelson. Book review: Introduction to Elementary Mathematical Logic, by A. A. Stolyar. Biometrics, 28(2):644–645, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556203.

Milliken:1972:BRL

[435] G. A. Milliken. Book review: Linear Models, by S. R. Searle. Biometrics, 28(3):898, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528782.

Morrison:1972:ASS

[436] Donald F. Morrison. The analysis of a single sample of repeated measurements. Biometrics, 28(1):55–71, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528961.

Patton:1972:PLF

[437] R. A. Patton and G. K. Krause. A parametrization of the logistic function which admits uncorrelated maximum likelihood estimators. Biometrics, 28(2):469–474, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556160.

Pernier:1972:AAD

[438] Jacques Pernier. Ajustement automatique des densités de probabiliţe D’Intervalles entre potentiels D’Action selon la loi de Wiener. (French) [Automatic adjustment of probability densities of intervals between action potentials according to Wiener’s law]. Biometrics, 28(3):737–745, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528759.

Pirie:1972:SRC

[439] W. R. Pirie and M. A. Hamdan. Some revised continuity corrections for discrete distributions. Biometrics, 28(3):693–701, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528756.

Pollak:1972:SET

[440] Edward Pollak. Some effects of two types of migration on the survival of a gene. Biometrics, 28(2):385–400, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556155.

Preece:1972:QNA

[441] D. A. Preece. 327. query: Non-additivity in two-way classifications with missing values. Biometrics, 28(2):574–577, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2556169.

Ramsey:1972:BAB

[442] Fred L. Ramsey. A Bayesian approach to bioassay. Biometrics, 28(3):841– 858, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528767. See

correction [483].

Rao:1972:RTR

[443] C. Radhakrishna Rao. Recent trends of research work in multivariate analysis. Biometrics, 28(1):3–22, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528958.

Rempe:1972:IPI

[444] U. Rempe and E. E. Weber. An illustration of the principal ideas of MANOVA. Biometrics, 28(1):235–238, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528970.

Robertson:1972:BRR

[445] C. A. Robertson. Book review: Robust Estimates of Location, by D. F. Andrews, P. J. Bickel, F. R. Hanpel, P. J. Huber, W. H. Rogers, J. W. Tukey. Biometrics, 28(4):1144–1145, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528653.

Rosenberg:1972:NCM

[446] Saul H. Rosenberg and Paul S. Levy. 343. Note: A characterization on misspecification in the general linear regression model. Biometrics, 28(4):1129–1133, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528646.

Roux:1972:BRP

[447] C. Z. Roux. Book review: Population Genetics in Animal Breeding, by F. Pirchner, Max von Krosigk. Biometrics, 28(4):1151–1152, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528662.

Schwartz:1972:OGT

[448] D. Schwartz. Obituary: Georges Teissier 1900–1972. Biometrics, 28(3): 891, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528773.

Searle:1972:BRG

[449] S. R. Searle. Book review: Generalized Inverse Matrices with Applications to Statistics, by R. M. Pringle, A. A. Rayner. Biometrics, 28(3):895–896, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528779.

Shanbhag:1972:VVB

[450] D. N. Shanbhag. On a vector-valued birth and death process. Biometrics, 28(2):417–425, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556157.

Shukla:1972:ITH

[451] G. K. Shukla. An invariant test for the homogeneity of variances in a two-way classification. Biometrics, 28(4):1063–1072, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528640.

Sirken:1972:NVC

[452] Monroe G. Sirken. 338. Note: Variance components of multiplicity estimators. Biometrics, 28(3):869–873, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528769.

Smouse:1972:CAM

[453] Peter E. Smouse. The canonical analysis of multiple species hybridization. Biometrics, 28(2):361–371, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556153.

Sprent:1972:MSS

[454] P. Sprent. The mathematics of size and shape. Biometrics, 28(1):23– 37, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528959.

Strauss:1972:BRS

[455] D. J. Strauss. Book review: Statistical Issues: a Reader for the Behaviorial Sciences, by Roger E. Kirk. Biometrics, 28(4):1151, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528660.

Tan:1972:CAG

[456] W. Y. Tan and W. C. Chang. Convolution approach to the genetic analysis of quantitative characters of self-fertilized populations. Biometrics, 28(4):1073–1090, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528641.

Taylor:1972:BRG

[457] H. M. Taylor. Book review: Gambling for Existence, by J. Reddingius. Biometrics, 28(3):896–897, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528780.

Taylor:1972:BRM

[458] H. M. Taylor. Book review: Markov Processes. Structure and Asymptotic Behavior, by M. Rosenblatt. Biometrics, 28(3):897–898, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528781.

Thomas:1972:EMV

[459] V. J. Thomas, N. A. Doughty, R. H. Fletcher, and J. G. Robertson. Estimation of microbial viability. Biometrics, 28(4):947–958, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528631.

Thompson:1972:NLM

[460] E. A. Thompson. 336. Note: The likelihood for multinomial proportions under stereographic projection. Biometrics, 28(2):618–620, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2556178.

Thornby:1972:RTL

[461] John I. Thornby. A robust test for linear regression. Biometrics, 28(2): 533–543, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2556165.

Tong:1972:BRS

[462] H. Tong. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Biometrics, 28(4):1143–1144, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528652.

Vaduva:1972:BRG

[463] I. Văduva. Book review: Global Systems Dynamics, by E. O. Attinger. Biometrics, 28(4):1145–1146, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528654.

Vaduva:1972:BRM

[464] I. Văduva. Book review: Mathematical Methods in Risk Theory, by H. Bühlman. Biometrics, 28(4):1146–1147, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528656.

Victor:1972:KMK

[465] N. Victor. Zur Klassifizierung Mehrdimensionaler Kontingenztafeln. (German) [Toward the classification of multidimensional contingency tables]. Biometrics, 28(2):427–441, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2556158.

Wang:1972:SPU

[466] Ying Y. Wang. Selection problems under multivariate normal distribution. Biometrics, 28(1):223–233, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528969.

Waugh:1972:ALP

[467] W. A. O’N. Waugh. The apparent ‘lag phase’ in a stochastic population model in which there is no variation in the conditions of growth. Biometrics, 28(2):329–336, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556150.

Weiler:1972:ISP

[468] H. Weiler. Inverse sampling of a Poisson distribution. Biometrics, 28(4):959–970, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528632.

Weisberg:1972:BCT

[469] Herbert I. Weisberg. Bayesian comparison of two ordered multinomial populations. Biometrics, 28(3):859–867, September 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528768.

Williams:1972:CSD

[470] D. A. Williams. The comparison of several dose levels with a zero dose control. Biometrics, 28(2):519–531, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2556164. See corrections [730].

Wittes:1972:NBE

[471] Janet T. Wittes. 331. Note: On the bias and estimated variance of Chapman’s two-sample capture-recapture population estimate. Biometrics, 28(2):592–597, June 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2556173.

Wu:1972:SDA

[472] Sing-Chou Chou Wu, James S. Williams, and Paul W. Mielke, Jr. Some designs and analyses for temporally independent experiments involving correlated bivariate responses. Biometrics, 28(4):1043–1061, December 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528639.

Abe:1973:NMK

[473] O. Abe. A note on the methodology of Knox’s tests of “Time and space interaction”. Biometrics, 29(1):67–77, March 1973. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529677.

Anonymous:1973:Aa

[474] Anonymous. Abstracts. Biometrics, 29(2):419–420, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529415.

Anonymous:1973:Ab

[475] Anonymous. Abstracts. Biometrics, 29(3):609–613, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529192.

Anonymous:1973:Ac

[476] Anonymous. Abstracts. Biometrics, 29(4):836–847, December 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529160.

Anonymous:1973:BRS

[477] Anonymous. Book review: Statistical Analysis: a Computer Oriented Approach by A. A. Afifi, S. P. Azen. Biometrics, 29(3):603, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529182.

Anonymous:1973:BSa

[478] Anonymous. The Biometric Society. Biometrics, 29(1):233–235, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529700.

Anonymous:1973:BSb

[479] Anonymous. The Biometric Society. Biometrics, 29(2):421–425, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529416.

Anonymous:1973:BSc

[480] Anonymous. The Biometric Society. Biometrics, 29(3):614–619, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529193.

Anonymous:1973:BSd

[481] Anonymous. The Biometric Society. Biometrics, 29(4):848–849, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529161.

Anonymous:1973:CAI

[482] Anonymous. Correction: The analysis of incomplete multi-way contingency tables. Biometrics, 29(4):829, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529150. See [392].

Anonymous:1973:CBA

[483] Anonymous. Correction: A Bayesian approach to bioassay. Biometrics, 29(4):830, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529152. See [442].

Anonymous:1973:CCS

[484] Anonymous. Correction: On comparisons of sensitivity, specificity and predictive value of a number of diagnostic procedures. Biometrics, 29(4): 829, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529149. See

[350].

Anonymous:1973:CVH

[485] Anonymous. Correction: Variance of heritability ratios estimated from phenotypic variance components. Biometrics, 29(4):830, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529151. See [402].

Anonymous:1973:FMa

[486] Anonymous. Front matter. Biometrics, 29(1):i–ii, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529670.

Anonymous:1973:FMb

[487] Anonymous. Front matter. Biometrics, 29(2):i–ii, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529388.

Anonymous:1973:FMc

[488] Anonymous. Front matter. Biometrics, 29(3):i–ii, September 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529165.

Anonymous:1973:FMd

[489] Anonymous. Front matter. Biometrics, 29(4):i–ii, December 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529129.

Anonymous:1973:NAa

[490] Anonymous. News and announcements. Biometrics, 29(1):236–239, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529701.

Anonymous:1973:NAb

[491] Anonymous. News and announcements. Biometrics, 29(2):426–427, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529417.

Anonymous:1973:NAc

[492] Anonymous. News and announcements. Biometrics, 29(3):620–621, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529194.

Anonymous:1973:NAd

[493] Anonymous. News and announcements. Biometrics, 29(4):850–851, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529162.

Anonymous:1973:OJE

[494] Anonymous. Obituary: John Edward Walsh 1919–1972. Biometrics, 29 (1):221, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529687.

Anonymous:1973:PPBa

[495] Anonymous. Papers to be published in Biometrics. Biometrics, 29(3): 622, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529195.

Anonymous:1973:PPBb

[496] Anonymous. Papers to be published in Biometrics. Biometrics, 29(4): 852, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529163.

Anonymous:1973:VI

[497] Anonymous. Volume information. Biometrics, 29(4):x–xlvii, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529128.

Beaver:1973:BRA

[498] Robert J. Beaver. Book review: Applied Multivariate Analysis, by S. James Press. Biometrics, 29(4):833–834, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529157.

Beaver:1973:BRD

[499] Robert J. Beaver. Book review: Discrete Techniques of Parameter Estimation. The Equation Error Formulation, by Jerry M. Mendel. Biometrics, 29(3):607, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529190.

Beaver:1973:BRE

[500] Robert J. Beaver. Book review: Elements of Applied Stochastic Processes, by U. N. Bhat. Biometrics, 29(2):415, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529409.

Beddington:1973:NOA

[501] J. R. Beddington and D. B. Taylor. 356. Note: Optimum age specific harvesting of a population. Biometrics, 29(4):801–809, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529145.

Bowden:1973:TBG

[502] David C. Bowden and R. Kirk Steinhorst. Tolerance bands for growth curves. Biometrics, 29(2):361–371, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529398.

Brillinger:1973:BRS

[503] David R. Brillinger. Book review: Stochastic Point Processes: Statistical Analysis, Theory and Applications, by P. A. W. Lewis. Biometrics, 29(2): 416–418, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529412.

Butcher:1973:BRT

[504] Donald F. Butcher. Book review: Theoretical Aspects of Population Genetics, by M. Kimura, T. Ohta. Biometrics, 29(3):606–607, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529188.

Carothers:1973:EUC

[505] A. D. Carothers. The effects of unequal catchability on jolly-seber estimates. Biometrics, 29(1):79–100, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529678.

Chatterjee:1973:MMP

[506] Samprit Chatterjee. A mathematical model for pest control. Biometrics, 29(4):727–734, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529138.

Cramer:1973:NSD

[507] Elliot M. Cramer. 345. Note: A simple derivation of the canonical correlation equations. Biometrics, 29(2):379–380, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529400.

Curnow:1973:SFP

[508] R. N. Curnow and M. F. Franklin. Some further problems in the classification of human chromosomes. Biometrics, 29(3):429–440, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529166. See corrections [610].

D:1973:BRB

[509] F. N. D. Book review: Biometrical Interpretation, by N. Gilbert. Biometrics, 29(3):606, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529187.

D:1973:BRF

[510] F. N. D. Book review: Fundamentals of Biometry, by L. N. Balaam. Biometrics, 29(3):603, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529183.

D:1973:BRM

[511] F. N. D. Book review: Mathematical Models for the Growth of Human Populations, by J. H. Pollard. Biometrics, 29(4):833, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529156.

N:1973:BRP

[512] F. N. D. Book review: Proceedings of the 3rd Conference of Forest Statisticians. Biometrics, 29(4):834, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529158.

D:1973:BRSa

[513] F. N. D. Book review: Statistical Methods for Rates and Proportions, by J. L. Fleiss. Biometrics, 29(3):606, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529186.

D:1973:BRSb

[514] F. N. D. Book review: Statistical Thinking: a Structural Approach, by J. L. Phillips, Jr. Biometrics, 29(3):608, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529191.

D:1973:BRSc

[515] F. N. D. Book review: Statistical Principles in Health Care Information, by S. J. Kilpatrick, Jr. Biometrics, 29(4):831–832, December 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529154.

David:1973:BRD

[516] F. N. David. Book review: Demographic Analysis: Methods, Results, Applications, by R. Pressat, J. Matras, N. Keyfitz. Biometrics, 29(1): 230–231, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529697.

David:1973:BRI

[517] F. N. David. Book review: An Introduction to Statistical Science in Agriculture, by D. J. Finney. Biometrics, 29(1):230, March 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529694.

David:1973:BRL

[518] F. N. David. Book review: Likelihood, by A. W. F. Edwards. Biometrics, 29(1):228–230, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529693.

David:1973:BRP

[519] F. N. David. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability, by L. LeCam, J. Neyman, E. L. Scott. Biometrics, 29(2):416, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529411.

David:1973:BRS

[520] F. N. David. Book review: The Structure of Human Populations, by G. A. Harrison, A. J. Boyce. Biometrics, 29(1):230, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529695.

Davis:1973:NEP

[521] C. E. Davis, James E. Grizzle, and James A. Bryan, II. 347. Note: Estimation of the probability of post transfusion hepatitis in hemophilia treatment. Biometrics, 29(2):386–392, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529402.

Dick:1973:MLE

[522] Nathan P. Dick and David C. Bowden. Maximum likelihood estimation for mixtures of two normal distributions. Biometrics, 29(4):781–790, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529143.

Didry:1973:EDP

[523] J. R. Didry, A. German, and J. Maccario. Étude du phénomène de coincidence dans un compteur de type Coulter. (French) [Study of the coincidence phenomenon in a Coulter type counter]. Biometrics, 29(2): 281–292, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529392.

Elfring:1973:GSD

[524] G. L. Elfring and J. R. Schultz. Group sequential designs for clinical trials. Biometrics, 29(3):471–477, September 1973. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529170.

Escoufier:1973:TVV

[525] Yves Escoufier. Le traitement des variables vectorielles. (French) [Processing vector variables]. Biometrics, 29(4):751–760, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529140.

Fix:1973:BRA

[526] Alan G. Fix. Book review: The Assessment of Population Affinities in Man, by J. S. Weiner, J. Huizinga. Biometrics, 29(1):231–232, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529698.

Forthofer:1973:ACF

[527] Ronald N. Forthofer and Gary G. Koch. An analysis for compounded functions of categorical data. Biometrics, 29(1):143–157, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529683.

Gail:1973:DSS

[528] Mitchell Gail and John J. Gart. The determination of sample sizes for use with the exact conditional test in 2 × 2 comparative trials. Biometrics, 29(3):441–448, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529167.

Garber:1973:BRI

[529] M. J. Garber. Book review: Introduction to Linear Programming, by J. E. Strum. Biometrics, 29(2):418, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529414.

Garber:1973:BRP

[530] M. J. Garber. Book review: Programming, An Introduction to Computer Techniques, by W. D. Maurer. Biometrics, 29(2):418, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529413.

Gilbert:1973:ABJ

[531] Richard O. Gilbert. Approximations of the bias in the Jolly–Seber capture–recapture model. Biometrics, 29(3):501–526, September 1973.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529173.

Glick:1973:SBM

[532] Ned Glick. Sample-based multinomial classification. Biometrics, 29(2): 241–256, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529389.

Gokhale:1973:BRB

[533] D. V. Gokhale. Book review: A Bibliography of Multivariate Statistical Analysis, by T. W. Anderson, S. Das Gupta, G. P. H. Styan. Biometrics, 29(2):414, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529407.

Gokhale:1973:BRD

[534] D. V. Gokhale. Book review: Distributions in Statistics: Continuous Multivariate Distributions, by N. L. Johnson, S. Kotz. Biometrics, 29 (1):230, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529696.

Gokhale:1973:BRP

[535] D. V. Gokhale. Book review: Proceedings of the Sixth Berkeley Symposium on Mathematical Statistics and Probability. Vol. II: Probability Theory, by L. LeCam, J. Neyman, E. L. Scott. Biometrics, 29(1):227– 228, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529692.

Gokhale:1973:BRSa

[536] D. V. Gokhale. Book review: Statistics of Directional Data, by K. V. Mardia. Biometrics, 29(3):607, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529189.

Gokhale:1973:BRSb

[537] D. V. Gokhale. Book review: Statistical Inference under Restrictions, by R. E. Barlow, D. J. Bartholomew, J. M. Bremner, H. D. Brunk. Biometrics, 29(4):831, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529153.

Gordon:1973:NCP

[538] A. D. Gordon. 359. Note: Classification in the presence of constraints. Biometrics, 29(4):821–827, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529148.

Griffiths:1973:MLE

[539] D. A. Griffiths. Maximum likelihood estimation for the beta-binomial distribution and an application to the household distribution of the total number of cases of a disease. Biometrics, 29(4):637–648, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529131.

Gurnow:1973:SPR

[540] R. N. Gurnow. A smooth population response curve based on an abrupt threshold and plateau model for individuals. Biometrics, 29(1):1–10, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529671.

Haberman:1973:ARC

[541] Shelby J. Haberman. The analysis of residuals in cross-classified tables. Biometrics, 29(1):205–220, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529686.

Haile:1973:NIQ

[542] Alemayehu Haile and J. T. Webster. 351. Note: An improved quadratic estimator for variance components. Biometrics, 29(2):408–413, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529406.

Hall:1973:TLD

[543] I. J. Hall and C. B. Sampson. Tolerance limits for the distribution of the product and quotient of normal variates. Biometrics, 29(1):109–119, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529680.

Hartigan:1973:MMF

[544] J. A. Hartigan. Minimum mutation fits to a given tree. Biometrics, 29(1): 53–65, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529676.

Hedayat:1973:NSO

[545] A. Hedayat. 348. Note: Self orthogonal latin square designs and their importance. Biometrics, 29(2):393–396, June 1973. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529403.

Hines:1973:BRF

[546] W. G. S. Hines. Book review: Foundations of Probability, by A. Rényi. Biometrics, 29(4):834–835, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529159.

Hosmer:1973:CIM

[547] David W. Hosmer, Jr. A comparison of iterative maximum likelihood estimates of the parameters of a mixture of two normal distributions under three different types of sample. Biometrics, 29(4):761–770, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529141.

Houck:1973:BRM

[548] J. P. Houck. Book review: Mathematical Model Building in Economics and Industry, by M. G. Kendall. Biometrics, 29(2):415–416, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529410.

Huszar:1973:BRM

[549] C. K. Huszar. Book review: Multivariate Morphometrics, by R. E. Blackith, R. A. Reyment. Biometrics, 29(3):603–604, September 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529184.

Jacquez:1973:SEE

[550] John A. Jacquez and Marija Norusis. Sampling experiments on the estimation of parameters in heteroscedastic linear regression. Biometrics, 29(4):771–780, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529142.

James:1973:NCB

[551] J. W. James. 353. Note: Covariances between relatives due to sex-linked genes. Biometrics, 29(3):584–588, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529178.

James:1973:RTM

[552] Kenneth E. James. Regression toward the mean in uncontrolled clinical studies. Biometrics, 29(1):121–130, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529681.

Johnson:1973:BRS

[553] J. P. Johnson. Book review: Statistical Papers in Honor of George W. Snedecor, by T. A. Bancroft. Biometrics, 29(1):227, March 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529691.

Kalbfleisch:1973:CPD

[554] J. G. Kalbfleisch and D. A. Sprott. The comparison of Poissondistributed observations. Biometrics, 29(1):223–224, March 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529688.

Lachin:1973:SPT

[555] John M. Lachin. On a stepwise procedure for two population Bayes decision rules using discrete variables. Biometrics, 29(3):551–564, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529175.

LaMotte:1973:QEV

[556] L. R. LaMotte. Quadratic estimation of variance components. Biometrics, 29(2):311–330, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529395.

Larsen:1973:STM

[557] Richard J. Larsen, Charles L. Holmes, and Clark W. Heath, Jr. A statistical test for measuring unimodal clustering: a description of the test and of its application to cases of acute leukemia in metropolitan Atlanta, Georgia. Biometrics, 29(2):301–309, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529394.

Levin:1973:BRS

[558] S. A. Levin and C. A. S. Hall. Book review: Systems Analysis and Simulation on Ecology, Vol. I, by B. C. Patten. Biometrics, 29(4):832– 833, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529155.

Lord:1973:PPE

[559] Gary E. Lord. Population and parameter estimation in the acoustic enumeration of a migrating fish population. Biometrics, 29(4):713–725, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529137.

Manly:1973:ALT

[560] B. F. J. Manly and G. A. F. Seber. Animal life tables from capturerecapture data. Biometrics, 29(3):487–500, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529172.

Mantel:1973:LRA

[561] Nathan Mantel and Charles Brown. A logistic reanalysis of Ashford and Sowden’s data on respiratory symptoms in British coal miners. Biometrics, 29(4):649–665, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529132.

Mantel:1973:QBO

[562] Nathan Mantel. Quantal bioassay with one animal at a dose level. Biometrics, 29(1):225–226, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529690.

Mantel:1973:SRS

[563] Nathan Mantel. Synthetic retrospective studies and related topics. Biometrics, 29(3):479–486, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529171.

Martin:1973:NBD

[564] Frank B. Martin. 349. Note: Beehive designs for observing variety competition. Biometrics, 29(2):397–402, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529404.

Meste:1973:ASR

[565] M. Meste. Sur l’analyse statistique des réponses dichotomiques à un stimulus multidimensionnel. (French) [On the statistical analysis of dichotomous responses to a multidimensional stimulus]. Biometrics, 29(4):735– 749, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529139.

Mitchell:1973:BRF

[566] Toby J. Mitchell. Book review: Fitting Equations to Data, by C. Daniel, F. S. Wood, J. W. Gorman. Biometrics, 29(3):604–605, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529185.

Molenaar:1973:NSA

[567] W. Molenaar. 350. Note: Simple approximations to the Poisson, binomial, and hypergeometric distributions. Biometrics, 29(2):403–407, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529405.

Moran:1973:AIB

[568] M. A. Moran. The analysis of incomplete blocks designs as used in the group screening of drugs. Biometrics, 29(1):131–142, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529682.

Myers:1973:LEM

[569] Max H. Myers, Benjamin F. Hankey, and Nathan Mantel. A logisticexponential model for use with response-time data involving regressor variables. Biometrics, 29(2):257–269, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529390.

Nadot:1973:AIA

[570] R. Nadot and G. Vaysseix. Apparentement et identité. Algorithme du calcul des coefficients D’Identité. (French) [Relatedness and identity. Algorithm for calculating identity coefficients]. Biometrics, 29(2):347–359, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529397.

Nam:1973:OSS

[571] Jun-Mo Nam. Optimum sample sizes for the comparison of the control and treatment. Biometrics, 29(1):101–108, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529679.

Pathak:1973:NMS

[572] K. B. Pathak. 354. Note: On a model for studying variation in the family size under different sex preferences. Biometrics, 29(3):589–595, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529179.

Pederson:1973:NAM

[573] D. G. Pederson. 358. Note: An approximate method of sampling a multinomial population. Biometrics, 29(4):814–821, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529147.

Peto:1973:NCA

[574] Richard Peto and M. C. Pike. 352. Note: Conservatism of the approximation ∑(O − E)^{2}/E in the logrank test for survival data or tumor incidence data. Biometrics, 29(3):579–584, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529177.

Peto:1973:WDC

[575] Richard Peto and Peter Lee. Weibull distributions for continuouscarcinogenesis experiments. Biometrics, 29(3):457–470, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529169.

Rao:1973:CBN

[576] B. Raja Rao, Sati Mazumdar, Jerome H. Waller, and C. C. Li. Correlation between the numbers of two types of children in a family. Biometrics, 29(2):271–279, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529391.

Rao:1973:EHV

[577] J. N. K. Rao. On the estimation of heteroscedastic variances. Biometrics, 29(1):11–24, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529672.

Reading:1973:ESS

[578] J. C. Reading and M. R. Klauber. Extensions of the successive screening method of classification. Biometrics, 29(4):791–800, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529144.

Rojas:1973:TTA

[579] Basilio A. Rojas. On Tukey’s test of additivity. Biometrics, 29(1):45– 52, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529675.

Rubin:1973:MRB

[580] Donald B. Rubin. Matching to remove bias in observational studies. Biometrics, 29(1):159–183, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2529684. See correction [609].

Rubin:1973:UMS

[581] Donald B. Rubin. The use of matched sampling and regression adjustment to remove bias in observational studies. Biometrics, 29(1):185–203, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529685.

Schafer:1973:IPL

[582] R. E. Schafer and T. S. Sheffield. Inferences on the parameters of the logistic distribution. Biometrics, 29(3):449–455, September 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529168.

Schultz:1973:MSP

[583] J. R. Schultz, F. R. Nichol, G. L. Elfring, and S. D. Weed. Multiple-stage procedures for drug screening. Biometrics, 29(2):293–300, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529393.

Sen:1973:NTA

[584] A. R. Sen. 346. Note: Theory and application of sampling on repeated occasions with several auxiliary variables. Biometrics, 29(2):381–385, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529401. See corrections [612].

Seneta:1973:NGI

[585] E. Seneta. 357. Note: On a genetic inequality. Biometrics, 29(4):810–814, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529146.

Sevacherian:1973:BRB

[586] Vahram Sevacherian. Book review: Basic Medical Statistics, by Anita K. Bahn. Biometrics, 29(2):414, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529408.

Shukla:1973:SET

[587] G. K. Shukla. Some exact tests of hypotheses about Grubbs’s estimators. Biometrics, 29(2):373–377, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529399.

Strauss:1973:BRE

[588] D. J. Strauss. Book review: Elementary Statistical Methods, by G. Barrie. Wetherill. Biometrics, 29(1):232, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529699.

Tan:1973:ASF

[589] W. Y. Tan. Applications of some finite Markov chain theories to two locus selfing model with selection. Biometrics, 29(2):331–346, June 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529396.

Taylor:1973:ADE

[590] J. Taylor. The analysis of designed experiments with censored observations. Biometrics, 29(1):35–43, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529674.

Thompson:1973:EVC

[591] R. Thompson. The estimation of variance and covariance components with an application when records are subject to culling. Biometrics, 29(3):527–550, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529174. See corrections [1129, 1241].

Tiku:1973:TGE

[592] M. L. Tiku. Testing group effects from Type II censored normal samples in experimental design. Biometrics, 29(1):25–33, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529673.

Trail:1973:ECB

[593] Stanley M. Trail and David L. Weeks. Extended complete block designs generated by BIBD. Biometrics, 29(3):565–578, September 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529176.

Trajstman:1973:TLH

[594] A. C. Trajstman. A two locus haploid population with overlapping generations. Biometrics, 29(4):701–711, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529136.

Ullman:1973:ATC

[595] Nelly S. Ullman and John A. Jacquez. Analysis of tests of chemotherapeutic agents involving repeated drug treatment. Biometrics, 29(4):677– 693, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529134.

Wagner:1973:WTD

[596] Stephen S. Wagner and Stuart A. Altmann. What time do the baboons come down from the trees? (an estimation problem). Biometrics, 29(4):623–635, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529130.

Watterson:1973:NRP

[597] G. A. Watterson. 355. Note: On a recent paper by Cook and Nassar. Biometrics, 29(3):595–600, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529180. See [359].

White:1973:CPD

[598] C. White and K. Detre. [The comparison of Poisson-distributed observations]. Biometrics, 29(1):225, March 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529689.

Williams:1973:ERP

[599] D. A. Williams. The estimation of relative potency from two parabolas in symmetric bioassays. Biometrics, 29(4):695–700, December 1973. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529135.

Williams:1973:OEA

[600] E. J. Williams. Obituary: Edmund Alfred Cornish 1909–1973. Biometrics, 29(3):601–602, September 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529181.

Yang:1973:USM

[601] Mark C. K. Yang and Carolyn J. Hursch. The use of a semi-Markov model for describing sleep patterns. Biometrics, 29(4):667–676, December 1973. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529133.

Anonymous:1974:Aa

[602] Anonymous. Abstracts. Biometrics, 30(1):215–216, March 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529638.

Anonymous:1974:Ab

[603] Anonymous. Abstracts. Biometrics, 30(2):374–388, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529665.

Anonymous:1974:Ac

[604] Anonymous. Abstracts. Biometrics, 30(3):559–566, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529215.

Anonymous:1974:BSa

[605] Anonymous. The Biometric Society. Biometrics, 30(1):217–220, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529639.

Anonymous:1974:BSb

[606] Anonymous. The Biometric Society. Biometrics, 30(3):567–568, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529216.

Anonymous:1974:BSc

[607] Anonymous. The Biometric Society. Biometrics, 30(4):731–734, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529252.

Anonymous:1974:CMQ

[608] Anonymous. Corrections: Mixed quasi-independent models for categorical data. Biometrics, 30(4):726, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529241. See [393].

Anonymous:1974:CMR

[609] Anonymous. Corrections: Matching to remove bias in observational studies. Biometrics, 30(4):728, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529245. See [580].

Anonymous:1974:CSF

[610] Anonymous. Corrections: Some further problems in the classification of human chromosomes. Biometrics, 30(4):726, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529240. See [508].

Anonymous:1974:CSV

[611] Anonymous. Corrections: The sampling variance of the correlation coefficients estimated in genetic experiments. Biometrics, 30(4):728, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529246. See [4, 21].

Anonymous:1974:CTA

[612] Anonymous. Corrections: Theory and application of sampling on repeated occasions with several auxiliary variables. Biometrics, 30(4):728, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529247. See [584].

Anonymous:1974:CTM

[613] Anonymous. Corrections: Tests for mortality and recruitment in tagrecapture experiments. Biometrics, 30(4):728, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529244. See [697].

Anonymous:1974:CTP

[614] Anonymous. Corrections: The two-period change-over design and its use in clinical trials. Biometrics, 30(4):727, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529243. See [2].

Anonymous:1974:CTT

[615] Anonymous. Corrections: A two-thirds replicate of a 3^{4} factorial. Biometrics, 30(4):727, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529242. See [212].

Anonymous:1974:FMa

[616] Anonymous. Front matter. Biometrics, 30(1):i–ii, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529612.

Anonymous:1974:FMb

[617] Anonymous. Front matter. Biometrics, 30(2):i–ii, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529643.

Anonymous:1974:FMc

[618] Anonymous. Front matter. Biometrics, 30(3):i–ii, September 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529197.

Anonymous:1974:FMd

[619] Anonymous. Front matter. Biometrics, 30(4):i–ii, December 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529221.

Anonymous:1974:NAa

[620] Anonymous. News and announcements. Biometrics, 30(1):221–223, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529640.

Anonymous:1974:NAb

[621] Anonymous. News and announcements. Biometrics, 30(2):395–396, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529667.

Anonymous:1974:NAc

[622] Anonymous. News and announcements. Biometrics, 30(3):569–571, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529217.

Anonymous:1974:NAd

[623] Anonymous. News and announcements. Biometrics, 30(4):735–736, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529253.

Anonymous:1974:OGZ

[624] Anonymous. Obituary: George Zyskind (1929–1974). Biometrics, 30(4): 725, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529239.

Anonymous:1974:PPBa

[625] Anonymous. Papers to be published in Biometrics. Biometrics, 30(1): 224, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529641.

Anonymous:1974:PPBb

[626] Anonymous. Papers to be published in Biometrics. Biometrics, 30(2): 397, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529668.

Anonymous:1974:PPBc

[627] Anonymous. Papers to be published in Biometrics. Biometrics, 30(3): 572, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529218.

Anonymous:1974:PPBd

[628] Anonymous. Papers to be published in Biometrics. Biometrics, 30(4): 737, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529254.

Anonymous:1974:VI

[629] Anonymous. Volume information. Biometrics, 30(4):vii–xxxvii, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529220.

Armitage:1974:BS

[630] Peter Armitage and Foster B. Cady. The Biometric Society. Biometrics, 30(2):389–394, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529666.

Armitage:1974:SCB

[631] P. Armitage, Judith M. Bailey, Aviva Petrie, L. Annable, and M. P. Stack-Dunne. Studies in the combination of bioassay results. Biometrics, 30(1):1–9, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529613. See corrections [729].

Ashford:1974:SMA

[632] J. R. Ashford and J. M. Cobby. A system of models for the action of drugs applied singly or jointly to biological organisms. Biometrics, 30(1): 11–31, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529614.

Bailey:1974:EPC

[633] R. Clifton Bailey, G. S. Eadie, and Frederick H. Schmidt. Estimation procedures for consecutive first order irreversible reactions. Biometrics, 30(1):67–75, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529618.

Beaver:1974:BRA

[634] R. J. Beaver. Book review: Asymptotic Theory of Rank Tests Independence, by F. H. Ruymgaart. Biometrics, 30(2):372, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529663.

Beaver:1974:BRS

[635] Robert J. Beaver. Book review: Statistics, by Roger A. Carlson. Biometrics, 30(1):211–212, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529633.

Bloomfield:1974:LTM

[636] P. Bloomfield. Linear transformations for multivariate binary data. Biometrics, 30(4):609–617, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529226.

Boardman:1974:CIV

[637] Thomas J. Boardman. Confidence intervals for variance components — a comparative Monte Carlo study. Biometrics, 30(2):251–262, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529647.

Breslow:1974:CAC

[638] N. Breslow. Covariance analysis of censored survival data. Biometrics, 30(1):89–99, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529620.

Brown:1974:AMC

[639] Morton B. Brown and Alan B. Forsythe. 372: The ANOVA and multiple comparisons for data with heterogeneous variances. Biometrics, 30(4):719–724, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529238.

Bulmer:1974:FPL

[640] M. G. Bulmer. On fitting the Poisson lognormal distribution to speciesabundance data. Biometrics, 30(1):101–110, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529621.

Chen:1974:TDC

[641] Tar Chen and Stephen E. Fienberg. Two-dimensional contingency tables with both completely and partially cross-classified data. Biometrics, 30(4):629–642, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529228.

Colgan:1974:AGC

[642] Patrick Colgan, Fred Cooke, and J. Terry Smith. 367: an analysis of group composition in assortatively mating populations. Biometrics, 30(4):693–696, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529233.

Cornell:1974:MEC

[643] John A. Cornell. More on extended complete block designs. Biometrics, 30(1):179–186, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529627.

Cox:1974:ECB

[644] N. R. Cox. Estimation of the correlation between a continuous and a discrete variable. Biometrics, 30(1):171–178, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529626.

Cramer:1974:CRS

[645] Elliot M. Cramer. 371: Computation of residual sums of squares in regression analysis. Biometrics, 30(4):717–718, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529237.

Curnow:1974:UAI

[646] R. N. Curnow. The use of additional information in estimating disease risks from family histories. Biometrics, 30(4):655–665, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529230.

D:1974:BRB

[647] F. N. D. Book review: Biometrika Tables for Statisticians. Volume 2, by E. S. Peason, H. O. Hartley. Biometrics, 30(2):372, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529662.

D:1974:BRP

[648] F. N. D. Book review: Probability and Statistical Inference in Ancient and Medieval Jewish Literature, by N. L. Rabinovitch. Biometrics, 30(3):557–558, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529212.

D:1974:BRSa

[649] F. N. D. Book review: Statistics and Society, by W. T. Federer. Biometrics, 30(3):557, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529211.

D:1974:BRSb

[650] F. N. D. Book review: Spatial Autocorrelation, by A. D. Cliff, J. K. Ord. Biometrics, 30(4):729, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529248.

D:1974:BRSc

[651] F. N. D. Book review: Statistical and Mathematical Aspects of Pollution Problems, by John W. Pratt. Biometrics, 30(4):730, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529251.

David:1974:BRF

[652] F. N. David. Book review: Foundations of Scientific Method: The Nineteenth Century, by R. N. Giere, R. S. Westfall. Biometrics, 30(2):371, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529658.

Dayananda:1974:ACB

[653] P. W. A. Dayananda. 369: an approximate chain-binomial model for simple epidemics. Biometrics, 30(4):705–708, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529235.

DellaCorte:1974:FLC

[654] M. Della Corte, L. Buricchi, and S. Romano. 363: On the fitting of linear combinations of exponentials. Biometrics, 30(2):367–369, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529657.

Fienberg:1974:BIP

[655] Stephen E. Fienberg. A Biometrics invited paper. Stochastic models for single neuron firing trains: a survey. Biometrics, 30(3):399–427, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529198.

Fletcher:1974:QLD

[656] R. Ian Fletcher. The quadric law of damped exponential growth. Biometrics, 30(1):111–124, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529622.

Frawley:1974:UJT

[657] W. H. Frawley. 364: Using the jackknife in testing dose responses in proportions near zero or one — revisited. Biometrics, 30(3):539–545, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529207.

Gail:1974:PCD

[658] Mitchell Gail. Power computations for designing comparative Poisson trials. Biometrics, 30(2):231–237, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529645. See corrections [1419].

Gallais:1974:CBA

[659] A. Gallais. Covariances between arbitrary relatives with linkage and epistasis in the case of linkage disequilibrium. Biometrics, 30(3):429– 446, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529199. See

corrections [987, 1045].

Gillings:1974:SFR

[660] Dennis B. Gillings. Some further results for bivariate generalizations of the Neyman Type A distribution. Biometrics, 30(4):619–628, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529227.

Gokhale:1974:BRB

[661] D. V. Gokhale. Book review: Bayesian Inference in Statistical Analysis, by G. E. P. Box, G. C. Tiao. Biometrics, 30(1):211, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529631.

Gokhale:1974:BRM

[662] D. V. Gokhale. Book review: Multivariate Statistical Inference, by D. G. Kabe, R. P. Gupta. Biometrics, 30(1):212, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529634.

Gokhale:1974:BRS

[663] D. V. Gokhale. Book review: Stochastic Processes and Applications in Biology and Medicine, by M. Iosifescu, P. Tautu. Biometrics, 30(2):371– 372, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529660.

Gower:1974:MPC

[664] J. C. Gower. Maximal predictive classification. Biometrics, 30(4):643– 654, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529229.

Green:1974:BRM

[665] Richard F. Green. Book review: The Mathematical Theory of the Dynamics of Biological Populations, by M. S. Bartlett, R. W. Hiorns. Biometrics, 30(3):557, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529210.

Green:1974:BRN

[666] Richard F. Green. Book review: Numerical Taxonomy, by Peter H. A. Sneath, Robert R. Sokal. Biometrics, 30(2):372–373, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529664.

Green:1974:BRQ

[667] Richard F. Green. Book review: I. Q. in the Meritocracy, by R. J. Herrnstein. Biometrics, 30(2):371, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529659.

Greenberg:1974:SCV

[668] Richard A. Greenberg, Steven Bayard, and David Byar. Selecting concomitant variables using a likelihood ratio step-down procedure and a method of testing goodness of fit in an exponential survival model. Biometrics, 30(4):601–608, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529225.

Haberman:1974:LLM

[669] Shelby J. Haberman. Log-linear models for frequency tables with ordered classifications. Biometrics, 30(4):589–600, December 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529224.

Hardy:1974:SMR

[670] Robert J. Hardy and Edward J. Weinman. A stochastic model for the reabsorption of isotonic saline by the tubules in the kidney. Biometrics, 30(1):33–40, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529615.

Henderson:1974:ICM

[671] C. R. Henderson, S. R. Searle, and L. R. Schaeffer. The invariance and calculation of Method 2 for estimating variance components. Biometrics, 30(4):583–588, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529223.

Hill:1974:EHB

[672] W. G. Hill and F. W. Nicholas. Estimation of heritability by both regression of offspring on parent and intra-class correlation of sibs in one experiment. Biometrics, 30(3):447–468, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529200.

Hill:1974:VRS

[673] William G. Hill. 362: Variability of response to selection in genetic experiments. Biometrics, 30(2):363–366, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529656.

Hines:1974:BRS

[674] W. G. S. Hines. Book review: Statistics with Applications to the Biological and Health Sciences, by R. D. Remington, M. A. Schork. Biometrics, 30(4):729–730, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529249.

Hocking:1974:APC

[675] R. R. Hocking and H. H. Oxspring. The analysis of partially categorized contingency data. Biometrics, 30(3):469–483, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529201.

Hoel:1974:EGT

[676] David G. Hoel and Kenny S. Crump. Estimating the generation-time distribution of an age-dependent branching process. Biometrics, 30(1): 125–135, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529623.

Jensen:1974:LMM

[677] A. L. Jensen. 365: Leslie matrix models for fisheries studies. Biometrics, 30(3):547–551, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529208.

Johnson:1974:PAC

[678] Mark A. Johnson. A probabilistic analysis of a class of male rat chemosterilants. Biometrics, 30(2):341–353, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529654.

Kalbfleisch:1974:IAH

[679] J. G. Kalbfleisch and D. A. Sprott. Inferences about hit number in a virological model. Biometrics, 30(1):199–208, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529629.

Kemp:1974:AHR

[680] C. D. Kemp and Nathan Mantel. 361: Approaches to a health research occupancy problem. Biometrics, 30(2):355–362, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529655.

Kemp:1974:STP

[681] C. D. Kemp, N. A. Goodchild, and K. Vijayan. 360: Significance tests in plots of multi-dimensional data in two dimensions. Biometrics, 30(1): 209–210, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529630.

Knight:1974:RLS

[682] William Knight. 366: a run-like statistic for ecological transects. Biometrics, 30(3):553–555, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529209.

Kovner:1974:PEW

[683] J. L. Kovner and S. A. Patil. Properties of estimators of wildlife population density for the line transect method. Biometrics, 30(2):225–230, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529644.

Lee:1974:BRM

[684] K. L. Lee. Book review: Mathematical Models of Conception and Birth, by M. C. Sheps, J. A. Menken. Biometrics, 30(4):730, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529250.

Manly:1974:MCT

[685] B. F. J. Manly. A model for certain types of selection experiments. Biometrics, 30(2):281–294, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529649.

Mantel:1974:AHR

[686] Nathan Mantel. Approaches to a health research occupancy problem. Biometrics, 30(3):355–362, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1974:ATC

[687] Nathan Mantel and Charles Brown. Alternative tests for comparing normal distribution parameters based on logistic regression. Biometrics, 30(3):485–497, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529202.

Maronna:1974:MCP

[688] Ricardo Maronna and Pablo M. Jacovkis. Multivariate clustering procedures with variable metrics. Biometrics, 30(3):499–505, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529203.

Maxwell:1974:BRA

[689] Floyd Maxwell. Book review: Applied Time Series Analysis for Managerial Forecasting, by Charles R. Nelson. Biometrics, 30(2):372, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529661.

McLachlan:1974:AUT

[690] G. J. McLachlan. An asymptotic unbiased technique for estimating the error rates in discriminant analysis. Biometrics, 30(2):239–249, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529646.

Mead:1974:TSP

[691] R. Mead. A test for spatial pattern at several scales using data from a grid of contiguous quadrats. Biometrics, 30(2):295–307, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529650.

Mendell:1974:MQT

[692] Nancy Role Mendell and Robert C. Elston. Multifactorial qualitative traits: genetic analysis and prediction of recurrence risks. Biometrics, 30(1):41–57, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529616.

Metzler:1974:BPE

[693] Carl M. Metzler. Bioavailability — a problem in equivalence. Biometrics, 30(2):309–317, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529651.

Mohn:1974:DRP

[694] Erik Mohn and Per Stavem. On the distribution of randomly placed discs. Biometrics, 30(1):137–156, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529624.

Ollivier:1974:RPD

[695] L. Ollivier. La regression parent–descendant dans le cas de déscendances subdivisées en familles de taille inégale. (French) [Parent–descendant regression in the case of descendants subdivided into families of unequal size]. Biometrics, 30(1):59–66, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529617.

Pike:1974:CCA

[696] M. C. Pike and P. G. Smith. A case-control approach to examine diseases for evidence of contagion, including diseases with long latent periods. Biometrics, 30(2):263–279, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529648.

Pollock:1974:TMR

[697] K. H. Pollock, D. L. Solomon, and D. S. Robson. Tests for mortality and recruitment in a K-sample tag-recapture experiment. Biometrics, 30(1): 77–87, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529619. See

corrections [613].

Robinson:1974:BRI

[698] J. Robinson. Book review: An Introduction to the Theory of Statistics, by R. L. Plackett. Biometrics, 30(1):213–214, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529637.

Robinson:1974:BRP

[699] J. Robinson. Book review: Probability and Mathematical Statistics — an Introduction, by E. Lukacs. Biometrics, 30(1):213, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529636.

Scheaffer:1974:DVI

[700] R. L. Scheaffer. On direct versus inverse sampling from mixed populations. Biometrics, 30(1):187–198, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529628.

Scott:1974:CAM

[701] A. J. Scott and M. Knott. A cluster analysis method for grouping means in the analysis of variance. Biometrics, 30(3):507–512, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529204. See corrections [728].

Sen:1974:LTS

[702] A. R. Sen, J. Tourigny, and G. E. J. Smith. On the line transect sampling method. Biometrics, 30(2):329–340, June 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529653. See corrections [1131, 1229].

Sethi:1974:QGC

[703] Suresh P. Sethi. Quantitative guidelines for communicable disease control program: a complete synthesis. Biometrics, 30(4):681–691, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529232.

Sevacherian:1974:BRE

[704] Vahram Sevacherian. Book review: Elementary Statistical Concepts, by R. B. Cain. Biometrics, 30(1):211, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529632.

Speed:1974:CEV

[705] F. M. Speed and R. R. Hocking. Computation of expectations, variances and covariances of ANOVA mean squares. Biometrics, 30(1):157–169, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529625.

Strauss:1974:BRE

[706] David J. Strauss. Book review: The Estimation of Animal Abundance, by F. F. Seber. Biometrics, 30(3):558, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529213.

Sykes:1974:BRP

[707] Z. M. Sykes. Book review: Population: Facts and Methods of Demography, by Nathan Keyfitz, Wilhelm Flieger. Biometrics, 30(1):212–213, March 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529635.

Tan:1974:AOT

[708] W. Y. Tan. 368: The approximate overall test for epistatic effects in biometrical genetics. Biometrics, 30(4):697–703, December 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529234.

Tenenbein:1974:SSD

[709] Aaron Tenenbein. 370: Sample size determination for the regression estimate of the mean. Biometrics, 30(4):709–716, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529236.

Thompson:1974:GIM

[710] E. A. Thompson. Gene identities and multiple relationships. Biometrics, 30(4):667–680, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529231.

Tietjen:1974:EAT

[711] Gary L. Tietjen. Exact and approximate tests for unbalanced random effects designs. Biometrics, 30(4):573–581, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529222.

Tracy:1974:BRI

[712] Derrick S. Tracy. Book review: Introduction to Biostatistics, by Robert R. Sokal, F. James Rohlf. Biometrics, 30(3):558, September 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529214.

Walter:1974:DHA

[713] S. D. Walter. On the detection of household aggregation of disease. Biometrics, 30(3):525–538, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529206.

Westlake:1974:UBI

[714] W. J. Westlake. The use of balanced incomplete block designs in comparative bioavailability trials. Biometrics, 30(2):319–327, June 1974. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529652.

Zahl:1974:AMA

[715] Samuel Zahl. Application of the S-method to the analysis of spatial pattern. Biometrics, 30(3):513–524, September 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529205.

Allen:1975:UAM

[716] D. J. Allen, H. J. Reimers, I. A. Feuerstein, and J. F. Mustard. The use and analysis of multiple responses in multicompartment cellular systems. Biometrics, 31(4):921–929, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529817.

Altham:1975:QIT

[717] Patricia M. E. Altham. 376: Quasi-independent triangular contingency tables. Biometrics, 31(1):233–238, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529725.

Anderson:1975:ECG

[718] David H. Anderson. Estimation and computation of the growth rate in Leslie’s and Lotka’s population models. Biometrics, 31(3):701–718, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529553.

Anderson:1975:FMI

[719] R. L. Anderson and Larry A. Nelson. A family of models involving intersecting straight lines and concomitant experimental designs useful in evaluating response to fertilizer nutrients. Biometrics, 31(2):303–318, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529422.

Anonymous:1975:Aa

[720] Anonymous. Abstracts. Biometrics, 31(1):255–258, March 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529734.

Anonymous:1975:Ab

[721] Anonymous. Abstracts. Biometrics, 31(2):590–604, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529450.

Anonymous:1975:Ac

[722] Anonymous. Abstracts. Biometrics, 31(4):1000–1018, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529836.

Anonymous:1975:BRG

[723] Anonymous. Book review: Games and Programs: Mathematics for Modeling by R. R. Singleton, W. F. Tyndall. Biometrics, 31(3):793, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529575.

Anonymous:1975:BSa

[724] Anonymous. The Biometric Society. Biometrics, 31(1):259–261, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529735.

Anonymous:1975:BSb

[725] Anonymous. The Biometric Society. Biometrics, 31(2):605–607, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529451.

Anonymous:1975:BSc

[726] Anonymous. The Biometric Society. Biometrics, 31(3):795–798, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529579.

Anonymous:1975:BSd

[727] Anonymous. The Biometric Society. Biometrics, 31(4):1020–1023, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529840.

Anonymous:1975:CCA

[728] Anonymous. Corrections: A cluster analysis method for grouping means in the analysis of variance. Biometrics, 31(4):1019, December 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529837. See [701].

Anonymous:1975:CSC

[729] Anonymous. Corrections: Studies in the combination of bioassay results. Biometrics, 31(4):1019, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529839. See [631].

Anonymous:1975:CTD

[730] Anonymous. Corrections: “A test for differences between treatment means when several dose levels are compared with a zero dose level” and “The comparison of several dose levels with a zero dose control”. Biometrics, 31(4):1019, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529838. See [294, 470].

Anonymous:1975:FMa

[731] Anonymous. Front matter. Biometrics, 31(1):i–ii, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529703.

Anonymous:1975:FMb

[732] Anonymous. Front matter. Biometrics, 31(2):i–ii, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529419.

Anonymous:1975:FMc

[733] Anonymous. Front matter. Biometrics, 31(3):i–ii, September 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529544.

Anonymous:1975:FMd

[734] Anonymous. Front matter. Biometrics, 31(4):i–ii, December 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529808.

Anonymous:1975:FMe

[735] Anonymous. Front matter. Biometrics, 31(2):i–ii, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529419.

Anonymous:1975:NAa

[736] Anonymous. News and announcements. Biometrics, 31(1):262–263, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529736.

Anonymous:1975:NAb

[737] Anonymous. News and announcements. Biometrics, 31(2):608–609, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529452.

Anonymous:1975:NAc

[738] Anonymous. News and announcements. Biometrics, 31(3):799–801, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529580.

Anonymous:1975:NAd

[739] Anonymous. News and announcements. Biometrics, 31(4):1024–1028, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529841.

Anonymous:1975:OGB

[740] Anonymous. Obituary: G. Barbensi, 1875–1974. Biometrics, 31(1):251, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529729.

Anonymous:1975:PPBa

[741] Anonymous. Papers to be published in Biometrics. Biometrics, 31(1): 264, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529737.

Anonymous:1975:PPBb

[742] Anonymous. Papers to be published in Biometrics. Biometrics, 31(2): 610, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529453.

Anonymous:1975:PPBc

[743] Anonymous. Papers to be published in Biometrics. Biometrics, 31(3): 801, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529581.

Anonymous:1975:PPBd

[744] Anonymous. Papers to be published in Biometrics. Biometrics, 31(4): 1029, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529842.

Anonymous:1975:VI

[745] Anonymous. Volume information. Biometrics, 31(4):ix–xxxvi, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529807.

Bancroft:1975:TTT

[746] T. A. Bancroft. Testimating, testipredicting, and testitesting as aids in using Snedecor and Cochran’s statistical methods. Biometrics, 31(2): 319–323, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529423.

Beaver:1975:BRE

[747] Robert J. Beaver. Book reviews: Experimental Design and Analysis, by Wayne Lee; Workbook for Experimental Design and Analysis, by Wayne Lee. Biometrics, 31(3):791, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529567.

Bernhardson:1975:TER

[748] Clemens S. Bernhardson. 375: Type I error rates when multiple comparison procedures follow a significant F test of ANOVA. Biometrics, 31(1): 229–232, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529724.

Brown:1975:EED

[749] A. H. D. Brown. Efficient experimental designs for the estimation of genetic parameters in plant populations. Biometrics, 31(1):145–160, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529716.

Brown:1975:MCN

[750] Morton B. Brown. 400: A method for combining non-independent, onesided tests of significance. Biometrics, 31(4):987–992, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529826. See correction [870].

Brown:1975:UIV

[751] Charles C. Brown. On the use of indicator variables for studying the time-dependence of parameters in a response-time model. Biometrics, 31(4):863–872, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529811.

Bulmer:1975:SAD

[752] M. G. Bulmer. The statistical analysis of density dependence. Biometrics, 31(4):901–911, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529815.

Burrows:1975:VSD

[753] Peter M. Burrows. Variances of selection differentials in normal samples. Biometrics, 31(1):125–133, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529714.

Byar:1975:SIA

[754] David P. Byar and Nathan Mantel. 393: Some interrelationships among the regression coefficient estimates arising in a class of models appropriate to response-time data. Biometrics, 31(4):943–947, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529819.

C:1975:BRC

[755] R. C. C. Book review: Comparative Statistical Inference, by V. Barnett. Biometrics, 31(2):587, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529440.

Chakravorti:1975:ADM

[756] S. R. Chakravorti and James E. Grizzle. Analysis of data from multiclinic experiments. Biometrics, 31(2):325–338, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529424.

Church:1975:IST

[757] J. D. Church and E. Benton Cobb. An improved Spearman-type estimator for the exponential parameter. Biometrics, 31(4):913–920, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529816.

Cornelius:1975:TIC

[758] P. L. Cornelius and J. W. Dudley. Theory of inbreeding and covariances between relatives under full-sib mating in diploids. Biometrics, 31(1): 169–187, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529718.

Cox:1975:PPG

[759] Gertrude M. Cox and Paul G. Homeyer. Professional and personal glimpses of George W. Snedecor. Biometrics, 31(2):265–301, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529421.

Crowley:1975:RCS

[760] J. Crowley and N. Breslow. 396: Remarks on the conservatism of Σ (0 - E)^{2} /E in survival data. Biometrics, 31(4):957–961, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529822.

D:1975:BRC

[761] F. N. D. Book review: Cluster Analysis, by E. J. Bijnen. Biometrics, 31 (2):587, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529441.

D:1975:BRD

[762] F. N. D. Book review: Data Quality Control and Editing, by J. I. Naus. Biometrics, 31(3):793, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529576.

D:1975:BRF

[763] F. N. D. Book review: Famine and Development: The Dutch Hunger Winter of 1944-45, by Z. Stein, M. Susser, G. Saenger, F. Marolla. Biometrics, 31(4):997, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529830.

D:1975:BRH

[764] F. N. D. Book review: The Heritage of Copernicus, by J. Neyman. Biometrics, 31(4):998, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529834.

D:1975:BRIa

[765] F. N. D. Book review: An Introduction to Medical Statistics, by H. O. Lancaster. Biometrics, 31(2):589, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529447.

D:1975:BRIb

[766] F. N. D. Book review: The Interpretation of Multiple Observations, by F. H. C. Marriott. Biometrics, 31(2):589, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529448.

D:1975:BRIc

[767] F. N. D. Book review: Introduction to Mathematics for Life Scientists, by E. Batschelet. Biometrics, 31(4):997, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529832.

D:1975:BRM

[768] F. N. D. Book review: Models in Ecology, by J. Maynard Smith. Biometrics, 31(2):589, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529449.

D:1975:BRP

[769] F. N. D. Book review: Population Growth Estimation, by E. S. Marks, W. Sellizer, K. J. Krotki. Biometrics, 31(3):793–794, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529578.

D:1975:BRRa

[770] F. N. D. Book review: Recent Results in Cancer Research: Cancer Registry, by E. Grundmann, E. Pedersen. Biometrics, 31(3):792, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529573.

D:1975:BRRb

[771] F. N. D. Book review: Recall Lapse in Demographic Enquiries, by R. K. Som. Biometrics, 31(4):998–999, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529835.

D:1975:BRS

[772] F. N. D. Book review: Statistical Methods & Scientific Inference, by R. A. Fisher. Biometrics, 31(2):588, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529443.

David:1975:BRMa

[773] F. N. David. Book review: Mèthodes du Traitement des chromiquesstatisques et prèvisions de ventes, by Jean-Claude Laloire. Biometrics, 31 (1):253, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529731.

David:1975:BRMb

[774] F. N. David. Book review: Modeles et methodes de l’analyse factorielle, by J. Torrens-Ibem. Biometrics, 31(1):253–254, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529733.

David:1975:BRSa

[775] F. N. David. Book review: Statistical Techniques in Simulation: Part I, by J. P. C. Kleijnen. Biometrics, 31(1):253, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529730.

David:1975:BRSb

[776] F. N. David. Book reviews: Statistical Theory in the Analysis of Experimental Designs, by J. Ogawa; Design of Experiments, by V. L. Anderson, R. A. MaClean; Analysis of Variance in Complex Experimental Designs, by H. R. Lindeman. Biometrics, 31(1):253, March 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529732.

Doubleday:1975:HMP

[777] W. G. Doubleday. Harvesting in matrix population models. Biometrics, 31(1):189–200, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529719.

Drolette:1975:EIF

[778] Margaret E. Drolette. The effect of incomplete follow-up. Biometrics, 31(1):135–144, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529715.

Duncan:1975:TIC

[779] David B. Duncan. t tests and intervals for comparisons suggested by the data. Biometrics, 31(2):339–359, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529425.

Engen:1975:CRS

[780] Steinar Engen. The coverage of a random sample from a biological community. Biometrics, 31(1):201–208, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529720.

Falco:1975:CIC

[781] Marsha J. Falco. 401: The calculation of inbreeding coefficients. Biometrics, 31(4):993, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529827.

Federer:1975:AD

[782] W. T. Federer and D. Raghavarao. On augmented designs. Biometrics, 31(1):29–35, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529707.

Federer:1975:SAR

[783] W. T. Federer, R. C. Nair, and D. Raghavarao. Some augmented row-column designs. Biometrics, 31(2):361–373, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529426.

Fienberg:1975:DVM

[784] Stephen E. Fienberg and Michael I. Waller. 392: A differential viability model for twin-pair blood group data. Biometrics, 31(3):785–790, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529566.

Finney:1975:ND

[785] D. J. Finney. Numbers and data. Biometrics, 31(2):375–386, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529427.

Fleiss:1975:MAB

[786] Joseph L. Fleiss. Measuring agreement between two judges on the presence or absence of a trait. Biometrics, 31(3):651–659, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529549.

Fryer:1975:MLE

[787] J. G. Fryer and R. J. Pethybridge. Maximum likelihood estimation for a special type of grouped data with an application to a dose-response problem. Biometrics, 31(3):633–642, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529547.

Gail:1975:RCS

[788] Mitchell Gail. A review and critique of some models used in competing risk analysis. Biometrics, 31(1):209–222, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529721.

Ganeshalingam:1975:BRC

[789] S. Ganeshalingam. Book review: A Computer Programming Language — with Business and Management Applications, by C. Carl Pegels. Biometrics, 31(3):792–793, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529574.

Gardner:1975:CST

[790] M. J. Gardner and D. J. P. Barker. A case study in techniques of allocation. Biometrics, 31(4):931–942, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529818.

Gentleman:1975:DOI

[791] J. F. Gentleman and M. B. Wilk. Detecting outliers. II. Supplementing the direct analysis of residuals. Biometrics, 31(2):387–410, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529428. See corrections [867].

Ghai:1975:SRB

[792] G. L. Ghai and E. Pollak. 387: On some results for a bivariate branching process. Biometrics, 31(3):761–763, September 1975. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529561.

Gokhale:1975:BRA

[793] D. V. Gokhale. Book review: The Analysis of Frequency Data, by Shelby J. Haberman. Biometrics, 31(4):997, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529831.

Gokhale:1975:BRT

[794] D. V. Gokhale. Book review: Theory of Probability, Vol. I, by B. De Finetti. Biometrics, 31(4):997–998, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529833.

Graybill:1975:MGWa

[795] Franklin A. Graybill. In memoriam: George W. Snedecor, 1882–1974. Biometrics, 31(2):iii–v, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529420.

Graybill:1975:MGWb

[796] Franklin A. Graybill. In memoriam: George W. Snedecor, 1882–1974. Biometrics, 31(2):iii–v, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529420.

Green:1975:BRMa

[797] R. F. Green. Book review: Manual of Quantitative Genetics, by Walter A. Becker. Biometrics, 31(3):792, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529571.

Green:1975:BRMb

[798] R. F. Green. Book review: Mathematical Problems in Biology, Victoria Conference by Pauline Van Den Driessche. Biometrics, 31(3):792, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529572.

Green:1975:BRS

[799] Richard F. Green. Book review: Stochastic Models in Biology, by Narendra S. Goel, Nira Richter-Dyn. Biometrics, 31(2):588, June 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529444.

Hagans:1975:BPE

[800] James A. Hagans. [Bioavailability — a problem in equivalence]. Biometrics, 31(2):585, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529439.

Hagans:1975:SME

[801] James A. Hagans and Paul Meier. Statistics and medical experimentation. Biometrics, 31(4):996, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529829.

Hall:1975:OST

[802] I. J. Hall. One-sided tolerance limits for a logistic distribution based on censored samples. Biometrics, 31(4):873–880, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529812.

Hall:1975:PDP

[803] J. W. Hall, J. C. Arnold, W. T. Waller, and J. Cairns, Jr. A procedure for the detection of pollution by fish movements. Biometrics, 31(1):11– 18, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529705.

Hamdan:1975:CBN

[804] M. A. Hamdan. 388: Correlation between the numbers of two types of children when the family size distribution is zero-truncated negative binomal. Biometrics, 31(3):765–766, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529562.

Hartley:1975:SPV

[805] H. O. Hartley and R. L. Sielken, Jr. A “Super-Population viewpoint” for finite population sampling. Biometrics, 31(2):411–422, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529429.

Harville:1975:ISP

[806] D. A. Harville. 373: Index selection with proportionality constraints. Biometrics, 31(1):223–225, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529722.

Hedayat:1975:SOL

[807] A. Hedayat. 386: Self orthogonal latin square designs and their importance, II. Biometrics, 31(3):755–759, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529560.

Henderson:1975:BLU

[808] C. R. Henderson. Best linear unbiased estimation and prediction under a selection model. Biometrics, 31(2):423–447, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529430.

Heuch:1975:RBS

[809] Ivar Heuch. The relationship between separation time and genetic distance based on angular transformations of gene frequencies. Biometrics, 31(3):685–700, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529552.

Hill:1975:TAG

[810] William G. Hill. Tests for association of gene frequencies at several loci in random mating diploid populations. Biometrics, 31(4):881–888, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529813.

Hocking:1975:SAN

[811] R. R. Hocking and M. H. Kutner. Some analytical and numerical comparisons of estimators for the mixed A.O.V. model. Biometrics, 31(1): 19–28, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529706.

Jensen:1975:CLE

[812] A. L. Jensen. Comparison of logistic equations for population growth. Biometrics, 31(4):853–862, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529810.

Jessen:1975:SCL

[813] Raymond J. Jessen. Square and cubic lattice sampling. Biometrics, 31(2): 449–471, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529431.

Kempthorne:1975:FMM

[814] Oscar Kempthorne. Fixed and mixed models in the analysis of variance. Biometrics, 31(2):473–486, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529432.

Keyfitz:1975:ILT

[815] Nathan Keyfitz and James Frauenthal. An improved life table method. Biometrics, 31(4):889–899, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529814.

Klauber:1975:STC

[816] Melville R. Klauber. 380: Space-time clustering tests for more than two samples. Biometrics, 31(3):719–726, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529554.

Kleinman:1975:PEV

[817] Joel C. Kleinman. 383: Proportions with extraneous variance: Two dependent samples. Biometrics, 31(3):737–743, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529557.

Koch:1975:GMA

[818] Gary G. Koch and H. Dennis Tolley. A generalized modified-χ^{2} analysis of categorical bacteria survival data from a complex dilution experiment. Biometrics, 31(1):59–92, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529710.

Krall:1975:SPS

[819] John M. Krall, Vincent A. Uthoff, and John B. Harley. A step-up procedure for selecting variables associated with survival. Biometrics, 31(1): 49–57, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529709.

Kuiper:1975:MCC

[820] F. Kent Kuiper and Lloyd Fisher. 391: a Monte Carlo comparison of six clustering procedures. Biometrics, 31(3):777–783, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529565.

Levitt:1975:RIC

[821] David J. Levitt. 384: Ranking in incomplete or complete blocks: a parametric analysis. Biometrics, 31(3):745–748, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529558.

Levy:1975:MRP

[822] Kenneth J. Levy. 378: a multiple range procedure for correlated variances in a two-way classification. Biometrics, 31(1):243–246, March 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529727.

Lin:1975:CDS

[823] Chuang-Sheng Lin and Pamela M. Morse. A compact design for spacing experiments. Biometrics, 31(3):661–671, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529550.

Lindsey:1975:RTN

[824] J. K. Lindsey. 379: The roles of transformations to normality. Biometrics, 31(1):247–249, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529728.

Littler:1975:TDN

[825] R. A. Littler and E. D. Fackerell. Transition densities for neutral multiallele diffusion models. Biometrics, 31(1):117–123, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529713.

Machin:1975:DPG

[826] David Machin. 385: On a design problem in growth studies. Biometrics, 31(3):749–753, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529559.

Mantel:1975:EGM

[827] Nathan Mantel and Joseph L. Fleiss. 381: The equivalence of the generalized McNemar tests for marginal homogeneity in 2^{3} and 3^{2} tables. Biometrics, 31(3):727–729, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529555.

Mantel:1975:GTG

[828] Nathan Mantel, M. Sobel, and R. Elashoff. Group testing with the goal of estimation. Biometrics, 31(4):994–995, December 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529828. With a reply by Sobel and

Elashoff.

Mantel:1975:NPI

[829] Nathan Mantel and Marvin A. Schneiderman. Non-parametric interval estimation of relative potency for dilution assays, including the case of non-monotone dosage response curves. Biometrics, 31(3):619–632, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529546.

Marriott:1975:SMN

[830] F. H. C. Marriott. 389: Separating mixtures of normal distributions. Biometrics, 31(3):767–769, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529563.

Mckinlay:1975:NCS

[831] Sonja M. Mckinlay. 382: a note on the chi-square test for pair-matched samples. Biometrics, 31(3):731–735, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529556.

McLachlan:1975:CIC

[832] G. J. McLachlan. Confidence intervals for the conditional probability of misallocation in discriminant analysis. Biometrics, 31(1):161–167, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529717.

McNeil:1975:HOD

[833] D. R. McNeil and J. W. Tukey. Higher-order diagnosis of two-way tables, illustrated on two sets of demographic empirical distributions. Biometrics, 31(2):487–510, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529433.

Mead:1975:BIP

[834] R. Mead and D. J. Pike. A biometrics invited paper. a review of response surface methodology from a biometric viewpoint. Biometrics, 31(4):803– 851, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529809.

Meier:1975:SME

[835] Paul Meier. Statistics and medical experimentation. Biometrics, 31(2): 511–529, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529434.

Neill:1975:EDC

[836] John J. Neill and Olive Jean Dunn. Equality of dependent correlation coefficients. Biometrics, 31(2):531–543, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529435.

Pocock:1975:STA

[837] Stuart J. Pocock and Richard Simon. Sequential treatment assignment with balancing for prognostic factors in the controlled clinical trial. Biometrics, 31(1):103–115, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529712. See corrections [871].

Quade:1975:NMC

[838] Dana Quade and Ibrahim A. Salama. 395: a note on minimum chi-square statistics in contingency tables. Biometrics, 31(4):953–956, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529821.

Rao:1975:SEP

[839] C. Radhakrishna Rao. Simultaneous estimation of parameters in different linear models and applications to biometric problems. Biometrics, 31(2): 545–554, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529436.

Robertson:1975:BRL

[840] C. A. Robertson. Book review: Linear Statistical Inference and Its Applications, by C. R. Rao. Biometrics, 31(3):791, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529568.

Rohlf:1975:GGT

[841] F. James Rohlf. Generalization of the gap test for the detection of multivariate outliers. Biometrics, 31(1):93–101, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529711.

Schaeffer:1975:DVC

[842] L. R. Schaeffer. 398: Disconnectedness and variance component estimation. Biometrics, 31(4):969–977, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529824.

Schafer:1975:MBP

[843] William D. Schafer and George B. Macready. 374: a modification of the Bonferroni procedure on contrasts which are grouped into internally independent sets. Biometrics, 31(1):227–228, March 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529723.

Sen:1975:URE

[844] A. R. Sen, S. Sellers, and G. E. J. Smith. The use of a ratio estimate in successive sampling. Biometrics, 31(3):673–683, September 1975. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529551. See corrections [991, 1093].

Sevacherian:1975:BRM

[845] V. Sevacherian. Book review: Mathematics in Biology: Calculus and Related Topics, by D. J. Clow, N. S. Urquhart. Biometrics, 31(3):791– 792, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529569.

Sevacherian:1975:BRR

[846] V. Sevacherian. Book review: Russian–English/English–Russian Glossary of Statistical Terms, by Samuel Kotz. Biometrics, 31(2):589, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529446.

Sevacherian:1975:BRS

[847] V. Sevacherian. Book review: Statistics for Biologists, by R. C. Campbell. Biometrics, 31(2):587–588, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529442.

Sevacherian:1975:BRT

[848] V. Sevacherian. Book review: Thinking Machines, by Irving Adler. Biometrics, 31(3):792, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529570.

Strauss:1975:BRC

[849] David J. Strauss. Book review: Clustering Algorithms, by J. A. Hartigan. Biometrics, 31(3):793, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529577.

Strauss:1975:BRP

[850] David J. Strauss. Book review: A Primer of Multivariate Statistics, by Richard J. Harris. Biometrics, 31(2):588–589, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529445.

Ury:1975:ECC

[851] Hans K. Ury. Efficiency of case-control studies with multiple controls per case: Continuous or dichotomous data. Biometrics, 31(3):643–649, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529548.

Vandermeer:1975:CPP

[852] John H. Vandermeer. 377: On the construction of the population projection matrix for a population grouped in unequal stages. Biometrics, 31(1):239–242, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529726.

Veevers:1975:ELB

[853] A. Veevers and T. B. Boffey. 397: On the existence of levelled beehive designs. Biometrics, 31(4):963–967, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529823.

Vincent:1975:PMP

[854] Thomas L. Vincent. Pest management programs via optimal control theory. Biometrics, 31(1):1–10, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529704.

Wani:1975:CPS

[855] J. K. Wani and Hing Po Lo. 390: Clopper-pearson system of confidence intervals for the logarithmic series distribution. Biometrics, 31(3):771– 775, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529564.

White:1975:RAV

[856] Robert F. White. Randomization and the analysis of variance. Biometrics, 31(2):555–571, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529437.

Williams:1975:ABR

[857] D. A. Williams. 394: The analysis of binary responses from toxicological experiments involving reproduction and teratogenicity. Biometrics, 31(4):949–952, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529820.

Wing:1975:FMA

[858] Howard J. Wing and Samuel H. H. Chan. 399: a first mathematical approximation of cerebral cortical responses to cerebello-thalamic signals using a thermodynamic analogue. Biometrics, 31(4):979–985, December 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529825.

Wiorkowski:1975:URA

[859] John J. Wiorkowski. Unbalanced regression analysis with residuals having a covariance structure of intra-class form. Biometrics, 31(3):611–618, September 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529545.

Yates:1975:DCP

[860] F. Yates. The design of computer programs for survey analysis — a contrast between the Rothamsted General Survey Program (RGSP) and SPSS. Biometrics, 31(2):573–584, June 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529438.

Zielezny:1975:CET

[861] Maria Zielezny and Olive Jean Dunn. Cost evaluation of a two-stage classification procedure. Biometrics, 31(1):37–47, March 1975. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529708.

Anonymous:1976:Aa

[862] Anonymous. Abstracts. Biometrics, 32(1):202–204, March 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529352.

Anonymous:1976:Ab

[863] Anonymous. Abstracts. Biometrics, 32(3):706–713, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529772.

Anonymous:1976:BSa

[864] Anonymous. The Biometric Society. Biometrics, 32(1):205–208, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529353.

Anonymous:1976:BSb

[865] Anonymous. The Biometric Society. Biometrics, 32(3):714–716, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529773.

Anonymous:1976:BSc

[866] Anonymous. The Biometric Society. Biometrics, 32(4):958–963, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529294.

Anonymous:1976:CDO

[867] Anonymous. Corrections: Detecting outliers. II. Supplementing the direct analysis of residuals. Biometrics, 32(4):955, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529285. See [791].

Anonymous:1976:CER

[868] Anonymous. Corrections: Estimating the rate constraints in a twocompartment stochastic model. Biometrics, 32(4):954, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529281. See [936].

Anonymous:1976:CGM

[869] Anonymous. Corrections: A generalization of a matrix occupancy problem. Biometrics, 32(4):954, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529280. See [975].

Anonymous:1976:CMC

[870] Anonymous. Correction: A method for combining non-independent onesided tests of significance. Biometrics, 32(4):955, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529286. See [750].

Anonymous:1976:CST

[871] Anonymous. Corrections: Sequential treatment assignment with balancing for prognostic factors in the controlled clinical trial. Biometrics, 32(4):954–955, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529282. See [837].

Anonymous:1976:FMa

[872] Anonymous. Front matter. Biometrics, 32(1):i–ii, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529335.

Anonymous:1976:FMb

[873] Anonymous. Front matter. Biometrics, 32(3):i–iii, September 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529739.

Anonymous:1976:FMc

[874] Anonymous. Front matter. Biometrics, 32(4):i–iii, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529257.

Anonymous:1976:NAa

[875] Anonymous. News and announcements. Biometrics, 32(1):209–211, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529354.

Anonymous:1976:NAb

[876] Anonymous. News and announcements. Biometrics, 32(3):717–718, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529774.

Anonymous:1976:NAc

[877] Anonymous. News and announcements. Biometrics, 32(4):964–965, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529295.

Anonymous:1976:PPBa

[878] Anonymous. Papers to be published in Biometrics. Biometrics, 32(1): 212, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529355.

Anonymous:1976:PPBb

[879] Anonymous. Papers to be published in Biometrics. Biometrics, 32(3): 719, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529775.

Anonymous:1976:PPBc

[880] Anonymous. Papers to be published in Biometrics. Biometrics, 32(4): 966, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529296.

Anonymous:1976:VI

[881] Anonymous. Volume information. Biometrics, 32(4):xiii–xlvi, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529256.

Atwood:1976:EMT

[882] Corwin L. Atwood and Adam Taube. Estimating mean time to reach a milestone, using retrospective data. Biometrics, 32(1):159–172, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529346.

Beaver:1976:BRI

[883] Robert J. Beaver. Book review: An Introduction to Probability Theory and Mathematical Statistics, by V. K. Rohatgi. Biometrics, 32(3):702– 703, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529767.

Beaver:1976:BRP

[884] Robert J. Beaver. Book review: Proceedings of the 8th International Biometric Conference, by L. C. A. Corsten, Tiberiu Postelnicu. Biometrics, 32(3):701, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529763.

Beaver:1976:DSS

[885] R. J. Beaver, H. A. David, and Roger R. Davidson. Discussion: “Science, statistics and paired comparisons” (Biometrics 32 (1976), no. 2, 213–232) by Ralph A. Bradley. Biometrics, 32(2):233–239, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [888].

Becker:1976:EEM

[886] Niels Becker. Estimation for an epidemic model. Biometrics, 32(4):769– 777, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529263.

Boob:1976:NCM

[887] B. S. Boob and H. L. Agrawal. A note on the construction of mutually orthogonal latin squares. Biometrics, 32(1):191–193, March 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529349.

Bradley:1976:SSP

[888] Ralph A. Bradley. Science, statistics, and paired comparisons. Biometrics, 32(2):213–232, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [885].

Breslow:1976:RAL

[889] N. Breslow. Regression analysis of the log odds ratio: a method for retrospective studies. Biometrics, 32(2):409–416, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:1976:CEC

[890] G. H. Brown. Combining estimates of category and subcategory proportions in a mixed population. Biometrics, 32(2):453–457, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brownie:1976:MAA

[891] C. Brownie and D. S. Robson. Models for allowing for age-dependent survival rates for band-return data. Biometrics, 32(2):305–323, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burnham:1976:MMN

[892] K. P. Burnham and D. R. Anderson. Mathematical models for nonparametric inferences from line transect data. Biometrics, 32(2):325–336, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapco:1976:ETH

[893] W. Chapco. An exact test of the Hardy–Weinberg law. Biometrics, 32(1): 183–189, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529348.

Chen:1976:ACT

[894] Tar Chen and Stephen E. Fienberg. The analysis of contingency tables with incompletely classified data. Biometrics, 32(1):133–144, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529344.

Cooke:1976:SPG

[895] Dennis Cooke and Jesus Alberto Leon. Stability of population growth determined by 2×2 Leslie matrix with density-dependent elements. Biometrics, 32(2):435–442, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Corbeil:1976:CVC

[896] R. R. Corbeil and S. R. Searle. A comparison of variance component estimators. Biometrics, 32(4):779–791, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529264.

Corsten:1976:GEC

[897] L. C. A. Corsten and K. R. Gabriel. Graphical exploration in comparing variance matrices. Biometrics, 32(4):851–863, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529269.

David:1976:BRA

[898] F. N. David. Book review: Applications in Statistics to Industrial Experimentation, by C. Daniel. Biometrics, 32(4):956, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529289.

David:1976:BRE

[899] F. N. David. Book review: Experimentation in Biology, by W. J. Ridgman. Biometrics, 32(4):957, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529293.

David:1976:BRF

[900] F. N. David. Book review: Field Experimentation with Fruit Trees and Other Perennial Plants: Commonwealth Agricultural Bureaux, by S. C. Pearce. Biometrics, 32(4):957, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529292.

David:1976:BRH

[901] F. N. David. Book review: On the History of Statistics and Probability, by D. B. Owen. Biometrics, 32(3):701–702, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529764.

David:1976:BRPa

[902] F. N. David. Book review: Proceedings of the Eighth International Confernece on Numerical Taxonomy, by G. F. Estabrook. Biometrics, 32(4): 957, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529290.

David:1976:BRPb

[903] F. N. David. Book review: Probability Sampling of Hospitals and Patients, by I. Hess, D. C. Riedel, T. B. Fitzpatrick. Biometrics, 32(4):957, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529291.

David:1976:BRS

[904] F. N. David. Book review: Statistical Methods in Animal Sciences: Indian Society of Agricultural Statistics, by V. N. Amble. Biometrics, 32(4): 956, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529287.

Davidson:1976:BMP

[905] Roger R. Davidson and Peter H. Farquhar. A bibliography on the method of paired comparisons. Biometrics, 32(2):241–252, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davis:1976:LNF

[906] Guy C. Davis, Jr. and Michael H. Kutner. The lagged normal family of probability density functions applied to indicator-dilution curves. Biometrics, 32(3):669–675, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529755.

Dawid:1976:PDD

[907] A. P. Dawid. Properties of diagnostic data distributions. Biometrics, 32(3):647–658, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529753.

Diggle:1976:SAS

[908] Peter J. Diggle, Julian Besag, and J. Timothy Gleaves. Statistical analysis of spatial point patterns by means of distance methods. Biometrics, 32(3):659–667, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529754.

Ellenberg:1976:TSO

[909] Jonas H. Ellenberg. Testing for a single outlier from a general linear regression. Biometrics, 32(3):637–645, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529752.

Faddy:1976:NGT

[910] M. J. Faddy. A note on the general time-dependent stochastic compartmental model. Biometrics, 32(2):443–448, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finney:1976:RA

[911] D. J. Finney. Radioligand assay. Biometrics, 32(4):721–740, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529258.

Freedman:1976:UPS

[912] L. S. Freedman and Susan J. White. On the use of Pocock and Simon’s method for balancing treatment numbers over prognostic factors in the controlled clinical trial. Biometrics, 32(3):691–694, September 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529759.

Freeman:1976:IBD

[913] G. H. Freeman. Incomplete block designs for three or four replicates of many treatments. Biometrics, 32(3):519–527, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529741. See corrections [989].

Freeman:1976:SDC

[914] G. H. Freeman. On the selection of designs for comparative experiments. Biometrics, 32(1):195–199, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529350.

Gail:1976:GOS

[915] Mitchell H. Gail and Sylvan B. Green. A generalization of the one-sided two-sample Kolmogorov–Smirnov statistic for evaluating diagnostic tests. Biometrics, 32(3):561–570, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529745.

Gokhale:1976:BRC

[916] D. V. Gokhale. Book review: Contributions to Applied Statistics, by W. J. Ziegler. Biometrics, 32(3):705, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529771.

Gokhale:1976:BRE

[917] D. V. Gokhale. Book review: Econometrics (Statistics: Textbooks and Monographs, Vol. 18), by Peter Schmidt. Biometrics, 32(3):703, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529768.

Gokhale:1976:BRS

[918] D. V. Gokhale. Book review: Sums of Independent Random Variables, by V. V. Petrov. Biometrics, 32(3):702, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529766.

Goodnight:1976:MIQ

[919] James H. Goodnight. Maximally-invariant quadratic unbiased estimators. Biometrics, 32(2):477–480, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Green:1976:BRE

[920] Richard F. Green. Book review: Ecology and Evolution of Communities, by Martin L. Cody, Jared M. Diamond. Biometrics, 32(3):701, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529762.

Green:1976:BRP

[921] Richard F. Green. Book review: Population Genetics and Ecology, by Samuel Karlin, Eviatar Nevo. Biometrics, 32(3):702, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529765.

Green:1976:BRS

[922] Richard F. Green. Book review: Sociobiology, by Edward O. Wilson. Biometrics, 32(3):704–705, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529770.

Habbema:1976:DAA

[923] J. D. F. Habbema. A discriminant analysis approach to the identification of human chromosomes. Biometrics, 32(4):919–928, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529274.

Harville:1976:CIS

[924] David A. Harville. Confidence intervals and sets for linear combinations of fixed and random effects. Biometrics, 32(2):403–407, June 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henderson:1976:SMC

[925] C. R. Henderson. A simple method for computing the inverse of a numerator relationship matrix used in prediction of breeding values. Biometrics, 32(1):69–83, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529339.

Hill:1976:OSC

[926] William G. Hill. Order statistics of correlated variables and implications in genetic selection programmes. Biometrics, 32(4):889–902, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529272.

Hocking:1976:ASV

[927] R. R. Hocking. The analysis and selection of variables in linear regression. Biometrics, 32(1):1–49, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hocking:1976:BIP

[928] R. R. Hocking. A biometrics invited paper. the analysis and selection of variables in linear regression. Biometrics, 32(1):1–49, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529336.

Holford:1976:LTC

[929] Theodore R. Holford. Life tables with concomitant information. Biometrics, 32(3):587–597, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529747.

Hwang:1976:ONP

[930] F. K. Hwang. An optimum nested procedure in binomial group testing. Biometrics, 32(4):939–943, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529277.

Jensen:1976:CSR

[931] D. R. Jensen. The comparison of several response functions with a standard. Biometrics, 32(1):51–59, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529337.

Jensen:1976:CVC

[932] Louis Jensen and Mervyn Stone. Cross-validatory choice of weights for inter- and intrablock estimation in balanced incomplete block designs. Biometrics, 32(3):677–681, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529756.

Johnson:1976:SNM

[933] Dallas E. Johnson. Some new multiple comparison procedures for the two-way AOV model with interaction. Biometrics, 32(4):929–934, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529275. See cor-

rections [990, 1064].

Juvancz:1976:BHP

[934] I. Juvancz. Is the Bradford Hill Principle still valid? Biometrics, 32(1): 200–201, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529351.

Kodell:1976:CER

[935] R. L. Kodell and J. H. Matis. Corrections to: “Estimating the rate constants in a two-compartment stochastic model” (Biometrics 32 (1976), no. 2, 377–390). Biometrics, 32(4):954, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [936].

Kodell:1976:ERC

[936] R. L. Kodell and J. H. Matis. Estimating the rate constants in a twocompartment stochastic model. Biometrics, 32(2):377–390, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [868, 935].

Lagakos:1976:SMC

[937] Stephen W. Lagakos. A stochastic model for censored-survival data in the presence of an auxiliary variable. Biometrics, 32(3):551–559, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529744.

Lamotte:1976:IQE

[938] Lynn Roy Lamotte. Invariant quadratic estimators in the random, oneway ANOVA model. Biometrics, 32(4):793–804, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529265.

Liotta:1976:SMM

[939] Lance A. Liotta, Gerald M. Saidel, and Jerome Kleinerman. Stochastic model of metastases formation. Biometrics, 32(3):535–550, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529743.

Liu:1976:EMT

[940] P. T. Liu and L. P. Chow. The efficiency of the multiple trial randomized response technique. Biometrics, 32(3):607–618, September 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529749.

Maclean:1976:SCD

[941] C. J. Maclean, N. E. Morton, R. C. Elston, and S. Yee. Skewness in commingled distributions. Biometrics, 32(3):695–699, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529760.

McDonald:1976:PTD

[942] Lyman L. McDonald, Victor W. Lowe, Robert K. Smidt, and Keren A. Meister. A preliminary test for discriminant analysis based on small samples. Biometrics, 32(2):417–422, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLachlan:1976:CSV

[943] G. J. McLachlan. A criterion for selecting variables for the linear discriminant function. Biometrics, 32(3):529–534, September 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529742.

Orchard:1976:BRU

[944] T. J. Orchard. Binary regression with unreplicated data. Biometrics, 32(4):935–938, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529276.

Patel:1976:EPE

[945] Y. C. Patel. Even point estimation and moment estimation in Hermite distribution. Biometrics, 32(4):865–873, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529270.

Prentice:1976:GPL

[946] Ross L. Prentice. A generalization of the probit and logit methods for dose response curves. Biometrics, 32(4):761–768, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529262.

Prentice:1976:ULM

[947] Ross Prentice. Use of the logistic model in retrospective studies. Biometrics, 32(3):599–606, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529748.

Quaas:1976:CDE

[948] R. L. Quaas. Computing the diagonal elements and inverse of a large numerator relationship matrix. Biometrics, 32(4):949–953, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529279.

Quednau:1976:CPE

[949] H. D. Quednau. The comparison of parameters estimated from several different samples by maximum likelihood. Biometrics, 32(3):683–688, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529757.

Quesenberry:1976:TNU

[950] C. P. Quesenberry, T. B. Whitaker, and J. W. Dickens. On testing normality using several samples: an analysis of peanut aflatoxin data. Biometrics, 32(4):753–759, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529261.

Rawlings:1976:OSS

[951] J. O. Rawlings. Order statistics for a special class of unequally correlated multinormal variates. Biometrics, 32(4):875–887, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529271.

Robinson:1976:EPM

[952] J. Robinson. Estimation of parameters for a model of transmitter release at synapses. Biometrics, 32(1):61–68, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529338.

Rohde:1976:CS

[953] Charles A. Rohde. Composite sampling. Biometrics, 32(2):273–282, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rorres:1976:OSY

[954] Chris Rorres. Optimal sustainable yield of a renewable resource. Biometrics, 32(4):945–948, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529278.

Rubin:1976:CMMa

[955] D. B. Rubin. Corrections to: “Multivariate matching methods that are equal percent bias reducing. I. Some examples” (Biometrics 32 (1976), no. 1, 109–120). Biometrics, 32(4):955, December 1976. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529283. See [957].

Rubin:1976:CMMb

[956] D. B. Rubin. Corrections to: “Multivariate matching methods that are equal percent bias reducing. II. Maximums on bias reduction for fixed sample sizes” (Biometrics 32 (1976), no. 1, 121–132). Biometrics, 32(4): 955, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529284. See

[958].

Rubin:1976:MMMa

[957] Donald B. Rubin. Multivariate matching methods that are equal percent bias reducing. I. Some examples. Biometrics, 32(1):109–120, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529342. See corrections [956].

Rubin:1976:MMMb

[958] Donald B. Rubin. Multivariate matching methods that are equal percent bias reducing. II. Maximums on bias reduction for fixed sample sizes. Biometrics, 32(1):121–132, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2529343. See corrections [955].

Sandland:1976:AMT

[959] R. L. Sandland. Application of methods of testing for independence between two Markov chains. Biometrics, 32(3):629–636, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529751.

Schreckengost:1976:ECB

[960] Jack F. Schreckengost and John A. Cornell. Extended complete block designs with correlated observations. Biometrics, 32(3):505–518, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529740.

Shah:1976:DAP

[961] B. K. Shah. Data analysis problems in the area of pharmacokinetics research. Biometrics, 32(1):145–157, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529345.

Smith:1976:BRM

[962] A. F. M. Smith. Book review: A Mathematical Theory of Evidence, by Glenn Shafer. Biometrics, 32(3):703–704, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529769.

Smith:1976:BRP

[963] A. F. M. Smith. Book review: Probability; Decision; Statistics, by James V. Bradley. Biometrics, 32(4):956, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529288.

Smith:1976:GTT

[964] P. G. Smith and M. C. Pike. Generalisations of two tests for the detection of household aggregation of disease. Biometrics, 32(4):817–828, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529267.

Tan:1976:APA

[965] W. Y. Tan. On the absorption probabilities and absorption times of finite homogeneous birth-death processes. Biometrics, 32(4):745–752, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529260.

Thompson:1976:DEE

[966] Robin Thompson. Design of experiments to estimate heritability when observations are available on parents and offspring. Biometrics, 32(2): 283–304, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [1128, 1240].

Thompson:1976:EMG

[967] Robin Thompson. The estimation of maternal genetic variances. Biometrics, 32(4):903–917, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529273.

Thompson:1976:FL

[968] W. A. Thompson, Jr. Fisherman’s luck. Biometrics, 32(2):265–271, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tong:1976:PES

[969] Y. L. Tong. Parameter estimation in studying circadian rhythms. Biometrics, 32(1):85–94, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529340.

VanVleck:1976:SDM

[970] L. D. Van Vleck. Selection for direct, maternal and grandmaternal genetic components of economic traits. Biometrics, 32(1):173–181, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529347.

Victor:1976:PAG

[971] N. Victor. Probleme der Auswahl geeigneter Zuordnungsregeln bei unvollständiger Information, insbesondere für kategoriale Daten. (German) [Problems in selecting appropriate assignment rules when information is incomplete, especially for categorical data]. Biometrics, 32(3):571–585, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529746.

Vitaliano:1976:ETT

[972] Peter Paul Vitaliano. On the equivalence of two tests for testing hypotheses about two correlated proportions. Biometrics, 32(3):689–690, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529758.

Walter:1976:CGM

[973] S. D. Walter. Corrections to: “A generalization of a matrix occupancy problem” (Biometrics 32 (1976), no. 2, 471–475). Biometrics, 32(4):954, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [975].

Walter:1976:EIA

[974] S. D. Walter. The estimation and interpretation of attributable risk in health research. Biometrics, 32(4):829–849, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529268.

Walter:1976:GMO

[975] S. D. Walter. A generalization of a matrix occupancy problem. Biometrics, 32(2):471–475, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [869, 973].

Wermuth:1976:ABM

[976] Nanny Wermuth. Analogies between multiplicative models in contingency tables and covariance selection. Biometrics, 32(1):95–108, March 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529341.

Westlake:1976:SCI

[977] Wilfred J. Westlake. Symmetrical confidence intervals for bioequivalence trials. Biometrics, 32(4):741–744, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529259.

Whittemore:1976:LCI

[978] Alice Whittemore and Bernard Altshuler. Lung cancer incidence in cigarette smokers: Further analysis of Doll and Hill’s data for British physicians. Biometrics, 32(4):805–816, December 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529266.

Williams:1976:CJA

[979] George W. Williams. Comparing the joint agreement of several raters with another rater. Biometrics, 32(3):619–627, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529750.

Willmott:1976:DCP

[980] A. John Willmott and F. Yates. Design of computer programs for survey analysis. Biometrics, 32(3):700, September 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529761.

Zielezny:1976:MGT

[981] Maria A. Zielezny. Maximum gain in the two-stage classification procedure. Biometrics, 32(2):481–484, June 1976. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1977:Aa

[982] Anonymous. Abstracts. Biometrics, 33(1):276–277, March 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529330.

Anonymous:1977:Ab

[983] Anonymous. Abstracts. Biometrics, 33(2):430, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529802.

Anonymous:1977:Ac

[984] Anonymous. Abstracts. Biometrics, 33(3):577–586, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529383.

Anonymous:1977:BSa

[985] Anonymous. The Biometrics Society. Biometrics, 33(2):431–433, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529803.

Anonymous:1977:BSb

[986] Anonymous. The Biometric Society. Biometrics, 33(3):587–588, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529384.

Anonymous:1977:CCB

[987] Anonymous. Corrections: Covariances between arbitrary relatives with linkage and epistasis, in the case of linkage disequilibrium. Biometrics, 33(4):766, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2529481. See [659].

Anonymous:1977:CEL

[988] Anonymous. Corrections: Estimation of linear compartmental model parameters using marginal likelihood. Biometrics, 33(4):767, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529484. See [1082].

Anonymous:1977:CIB

[989] Anonymous. Corrections: Incomplete block designs for three or four replicates of many treatments. Biometrics, 33(4):766, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529482. See [913].

Anonymous:1977:CSN

[990] Anonymous. Corrections: Some new multiple comparison procedures for the two-way AOV model with interaction. Biometrics, 33(4):766, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529483. See [933].

Anonymous:1977:CUR

[991] Anonymous. Corrections: The use of a ratio estimate in successive sampling. Biometrics, 33(4):767, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529485. See [844].

Anonymous:1977:EN

[992] Anonymous. Editor’s note. Biometrics, 33(2):419, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529796.

Anonymous:1977:FMa

[993] Anonymous. Front matter. Biometrics, 33(1):i–272, March 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529298.

Anonymous:1977:FMb

[994] Anonymous. Front matter. Biometrics, 33(2):i–iii, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529777.

Anonymous:1977:FMc

[995] Anonymous. Front matter. Biometrics, 33(3):i–iii, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529357.

Anonymous:1977:FMd

[996] Anonymous. Front matter. Biometrics, 33(4):i–iii, December 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529456.

Anonymous:1977:I

[997] Anonymous. [Introduction]. Biometrics, 33(1):230, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529317.

Anonymous:1977:NAa

[998] Anonymous. News and announcements. Biometrics, 33(1):279–280, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529332.

Anonymous:1977:NAb

[999] Anonymous. News and announcements. Biometrics, 33(2):434–435, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529804.

Anonymous:1977:NAc

[1000] Anonymous. News and announcements. Biometrics, 33(3):589–590, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529385.

Anonymous:1977:NAd

[1001] Anonymous. News and announcements. Biometrics, 33(4):769, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529490.

Anonymous:1977:OGA

[1002] Anonymous. Obituary: Giulio Alfredo Maccacaro, 1924–1977. Biometrics, 33(4):758, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529476.

Anonymous:1977:OHL

[1003] Anonymous. Obituary: Henry Laurence Lucas, Jr. 1916–1977. Biometrics, 33(3):573, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529379.

Anonymous:1977:PPBa

[1004] Anonymous. Papers to be published in Biometrics. Biometrics, 33(1): 281, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529333.

Anonymous:1977:PPBb

[1005] Anonymous. Papers to be published in Biometrics. Biometrics, 33(2): 436, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529805.

Anonymous:1977:PPBc

[1006] Anonymous. Papers to be published in Biometrics. Biometrics, 33(3): 591, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529386.

Anonymous:1977:PPBd

[1007] Anonymous. Papers to be published in Biometrics. Biometrics, 33(4): 770, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529491.

Anonymous:1977:VI

[1008] Anonymous. Volume information. Biometrics, 33(4):xiii–xxxviii, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529455.

Armitage:1977:PP

[1009] Peter Armitage. [The perimeter of “Parameter”]. Biometrics, 33(4):765, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529480.

Barron:1977:EME

[1010] Bruce A. Barron. The effects of misclassification on the estimation of relative risk. Biometrics, 33(2):414–418, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529795.

Bartolucci:1977:CBT

[1011] Alfred A. Bartolucci and James M. Dickey. Comparative Bayesian and traditional inference for gamma-modeled survival data. Biometrics, 33(2):343–354, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529784.

Beaver:1977:BRD

[1012] Robert J. Beaver. Book review: Design and Analysis of Experiments, by Douglas C. Montgomery. Biometrics, 33(1):273–274, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529326.

Becker:1977:EDT

[1013] Niels Becker. Estimation for discrete time branching processes with application to epidemics. Biometrics, 33(3):515–522, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529366.

Campbell:1977:QCK

[1014] D. B. Campbell and C. A. Oprian. Query: On computing the Kolmogorov–Smirnov statistic. Biometrics, 33(4):750–751, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529474.

Canner:1977:MTD

[1015] Paul L. Canner. Monitoring treatment differences in long-term clinical trials. Biometrics, 33(4):603–615, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529458.

Cash:1977:ISU

[1016] W. S. Cash. An improved solution for the ultimate probability of fixation of a favorable allele. Biometrics, 33(3):528–532, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529368. See corrections [1130, 1160].

Coan:1977:JCL

[1017] Atilio L. Coan, Robert C. Francis, and A. L. Jensen. On Jensen’s comparison of logistic equations for population growth. Biometrics, 33(2): 426–427, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529799.

Cockerham:1977:QAR

[1018] C. Clark Cockerham and B. S. Weir. Quadratic analyses of reciprocal crosses. Biometrics, 33(1):187–203, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529312.

Cook:1977:ECL

[1019] G. W. Cook. Estimation of contrasts of log-frequencies. Biometrics, 33(3):548–552, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529373.

Copenhaver:1977:QAQ

[1020] T. W. Copenhaver and P. W. Mielke. Quantit analysis: a quantal assay refinement. Biometrics, 33(1):175–186, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529311.

Cornelius:1977:CIW

[1021] P. L. Cornelius and J. Byars. Computing iterative weighted least squares estimates of genetic variance components. Biometrics, 33(2):375–382, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529787.

Crump:1977:CIT

[1022] K. S. Crump, H. A. Guess, and K. L. Deal. Confidence intervals and test of hypotheses concerning dose response relations inferred from animal carcinogenicity data. Biometrics, 33(3):437–451, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529358.

Crump:1977:OQT

[1023] K. S. Crump and Nathan Mantel. Open query: Theoretical problems in the modified mantel-bryan procedure. Biometrics, 33(4):752–757, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529475.

David:1977:BRD

[1024] F. N. David. Book review: Decision Making and Medical Care, by F. T. Dombal, F. Gremy. Biometrics, 33(1):273, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529323.

David:1977:BRE

[1025] F. N. David. Book review: Exploratory Data Analysis, by John W. Tukey. Biometrics, 33(4):768, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529486.

David:1977:BRG

[1026] F. N. David. Book review: Geological Factor Analysis, by K. G. Joreskog, J. E. Klovan, R. A. Reyment. Biometrics, 33(1):273, March 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529325.

David:1977:BRL

[1027] F. N. David. Book review: Lecture Notes in Biomathematics: Mathematical Models in Biological Discovery, by D. L. Solomon, C. Walter. Biometrics, 33(4):768, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529488.

David:1977:BRM

[1028] F. N. David. Book review: Multivariate Analysis IV, by P. R. Krishnaiah. Biometrics, 33(4):768, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529489.

David:1977:BRP

[1029] F. N. David. Book review: Parameter Estimation in Engineering and Science, by J. V. Beck, K. J. Arnold. Biometrics, 33(4):768, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529487.

David:1977:BRSa

[1030] F. N. David. Book review: Sufficient Statistics, by V. S. Huzurbazar. Biometrics, 33(1):273, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529324.

David:1977:BRSb

[1031] F. N. David. Book review: Studies in Probability and Statistics, by E. J. Williams. Biometrics, 33(1):275, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529329.

Davidson:1977:EBT

[1032] Roger R. Davidson and Robert J. Beaver. On extending the Bradley– Terry model to incorporate within-pair order effects. Biometrics, 33 (4):693–702, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529467.

Davies:1977:REF

[1033] O. L. Davies and J. Y. Ku. Re-examination of the fitting of the Richards growth function. Biometrics, 33(3):546–547, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529372.

DeMets:1977:ESR

[1034] David DeMets and Max Halperin. Estimation of a simple regression coefficient in samples arising from a sub-sampling procedure. Biometrics, 33(1):47–56, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529302.

Diggle:1977:DRH

[1035] Peter J. Diggle. The detection of random heterogeneity in plant populations. Biometrics, 33(2):390–394, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529790.

Donner:1977:UAI

[1036] Allan Donner. The use of auxiliary information in the design of a clinical trial. Biometrics, 33(2):305–314, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529780.

Dunn:1977:ART

[1037] James E. Dunn and Phillip S. Gipson. Analysis of radio telemetry data in studies of home range. Biometrics, 33(1):85–101, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529305.

Dunnett:1977:STE

[1038] C. W. Dunnett and M. Gent. Significance testing to establish equivalence between treatments, with special reference to data in the form of 2 × 2 tables. Biometrics, 33(4):593–602, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529457.

Ejigou:1977:ERR

[1039] Ayenew Ejigou and Richard McHugh. Estimation of relative risk from matched pairs in epidemiologic research. Biometrics, 33(3):552–556, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529374.

Elston:1977:QEH

[1040] R. C. Elston, William G. Hill, and Charles Smith. Query: Estimating “Heritability” of a dichotomous trait. Biometrics, 33(1):231–236, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529318.

Elston:1977:THW

[1041] R. C. Elston and R. Forthofer. Testing for Hardy–Weinberg equilibrium in small samples. Biometrics, 33(3):536–542, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529370.

Emigh:1977:PPV

[1042] Ted H. Emigh. Partition of phenotypic variance under unknown dependent association of genotypes and environments. Biometrics, 33(3):505– 514, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529365.

Engen:1977:CTD

[1043] Steinar Engen. Comments on two different approaches to the analysis of species frequency data. Biometrics, 33(1):205–213, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529313.

Free:1977:QDN

[1044] S. M. Free, Jr. Query: Designing nutrition studies with litters. Biometrics, 33(3):561–563, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529376.

Gallais:1977:CCB

[1045] A. Gallais. Corrections to: “Covariances between arbitrary relatives with linkage and epistasis, in the case of linkage disequilibrium” (Biometrics 30 (1974), 429–446). Biometrics, 33(4):766, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [659].

Gani:1977:FMG

[1046] J. Gani and I. W. Saunders. Fitting a model to the growth of yeast colonies. Biometrics, 33(1):113–120, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529307.

Garber:1977:BRD

[1047] M. J. Garber. Book review: Discriminant Analysis, by Peter A. Lachenbruck. Biometrics, 33(2):429, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529801.

Gates:1977:BS

[1048] Charles E. Gates. The Biometric Society. Biometrics, 33(1):278, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529331.

Gilbert:1977:SFI

[1049] Roger I. Gilbert. A simple formula for interpolating tables of x^{2}. Biometrics, 33(2):383–385, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529788.

Gokhale:1977:BRM

[1050] D. V. Gokhale. Book review: Multivariate Analysis in Behavioral Research, by A. E. Maxwell. Biometrics, 33(3):576, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529381.

Gokhale:1977:BRS

[1051] D. V. Gokhale. Book review: Statistics Made Relevant: a Casebook of Real Life Examples, by Paul Baum, Ernest M. Scheuer. Biometrics, 33 (1):273, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529322.

Goldstein:1977:PBM

[1052] Matthew Goldstein and Edward Wolf. On the problem of bias in multinomial classification. Biometrics, 33(2):325–331, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529782.

Gordon:1977:AES

[1053] A. D. Gordon and J. T. Henderson. An algorithm for Euclidean sum of squares classification. Biometrics, 33(2):355–362, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529785.

Green:1977:BRB

[1054] Richard F. Green. Book review: A Bibliography of Quantitative Ecology, by Vincent Schultz, L. L. Eberhardt, J. M. Thomas, M. I. Cochran. Biometrics, 33(1):274, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529327.

Green:1977:BRH

[1055] Richard F. Green. Book review: Human Evolutionary Trees, by E. A. Thompson. Biometrics, 33(1):274–275, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529328.

Green:1977:BRS

[1056] Richard F. Green. Book review: Stochastic Models in Population Genetics, by Wen-Hsiung Li. Biometrics, 33(3):576, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529382.

Griffiths:1977:AMG

[1057] David Griffiths. Avoidance-modified generalised distributions and their application to studies of superparasitism. Biometrics, 33(1):103–112, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529306.

Gunst:1977:AEM

[1058] Richard F. Gunst and Robert L. Mason. Advantages of examining multicollinearities in regression analysis. Biometrics, 33(1):249–260, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529320.

Hakulinen:1977:ERD

[1059] T. Hakulinen and M. Rahiala. An example on the risk dependence and additivity of intensities in the theory of competing risks. Biometrics, 33(3):557–559, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529375.

Hartley:1977:BIP

[1060] H. O. Hartley and R. L. Sielken, Jr. A biometrics invited paper. Estimation of “Safe doses” in carcinogenic experiments. Biometrics, 33(1): 1–30, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529299.

Hill:1977:OSC

[1061] William G. Hill. Order statistics of correlated variables and implications in genetic selection programmes. II. Response to selection. Biometrics, 33(4):703–712, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529468.

Horn:1977:GFT

[1062] Susan Dadakis Horn. Goodness-of-fit tests for discrete data: a review and an application to a health impairment scale. Biometrics, 33(1):237– 247, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529319.

Johnson:1977:ARA

[1063] N. L. Johnson. Approximate relationships among some estimators of mortality probabilities. Biometrics, 33(3):542–545, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529371.

Johnson:1977:CSN

[1064] Dallas E. Johnson. Corrections to: “Some new multiple comparison procedures for the two-way AOV model with interaction” (Biometrics 32 (1976), no. 4, 929–934). Biometrics, 33(4):766, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [933].

Johnson:1977:PMR

[1065] J. P. Johnson, T. A. Bancroft, and Chien-Pai Pai Han. A pooling methodology for regressions in prediction. Biometrics, 33(1):57–67, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529303.

Kirkwood:1977:PSB

[1066] T. B. L. Kirkwood. Predicting the stability of biological standards and products. Biometrics, 33(4):736–742, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529472.

Koch:1977:GMA

[1067] Gary G. Koch, J. Richard Landis, Jean L. Freeman, Daniel H. Freeman, Jr., and Robert G. Lehnen. A general methodology for the analysis of experiments with repeated measurement of categorical data. Biometrics, 33(1):133–158, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529309.

Lachin:1977:ICC

[1068] John M. Lachin. Informed consent in clinical investigations. Biometrics, 33(4):761–762, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529478.

Lachin:1977:SSD

[1069] John M. Lachin. Sample size determinations for r × c comparative trials. Biometrics, 33(2):315–324, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529781.

Lagakos:1977:UAV

[1070] Stephen W. Lagakos. Using auxiliary variables for improved estimates of survival time. Biometrics, 33(2):399–404, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529792.

Landis:1977:AHK

[1071] J. Richard Landis and Gary G. Koch. An application of hierarchical kappa-type statistics in the assessment of majority agreement among multiple observers. Biometrics, 33(2):363–374, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529786.

Landis:1977:MOA

[1072] J. Richard Landis and Gary G. Koch. The measurement of observer agreement for categorical data. Biometrics, 33(1):159–174, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529310.

Landis:1977:OWC

[1073] J. Richard Landis and Gary G. Koch. A one-way components of variance model for categorical data. Biometrics, 33(4):671–679, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529465.

Leonard:1977:ABA

[1074] Tom Leonard. An alternative Bayesian approach to the Bradley–Terry model for paired comparisons. Biometrics, 33(1):121–132, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529308.

Louis:1977:SAC

[1075] Thomas A. Louis. Sequential allocation in clinical trials comparing two exponential survival curves. Biometrics, 33(4):627–634, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529460.

Madsen:1977:GCM

[1076] K. S. Madsen. A growth curve model for studies in morphometrics. Biometrics, 33(4):659–669, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529464.

Magnus:1977:CES

[1077] A. Magnus, P. W. Mielke, and T. W. Copenhaver. Closed expressions for the sum of an infinite series with application to quantal response assays. Biometrics, 33(1):221–223, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529315.

Manly:1977:ATR

[1078] B. F. J. Manly. The analysis of trapping records for birds trapped in mist nets. Biometrics, 33(2):404–410, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529793.

Mantel:1977:DWW

[1079] Nathan Mantel and Wilfred J. Westlake. Do we want confidence intervals symmetrical about the null value? Biometrics, 33(4):759–760, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529477.

McCabe:1977:RCD

[1080] George P. McCabe, Jr. and Myra L. Samuels. Random censoring and dendritic trees. Biometrics, 33(1):69–84, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529304.

McKinlay:1977:PMR

[1081] Sonja M. McKinlay. Pair-matching — a reappraisal of a popular technique. Biometrics, 33(4):725–735, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529471.

Minder:1977:ELC

[1082] Ch. E. Minder and I. McMillan. Estimation of linear compartmental model parameters using marginal likelihood. Biometrics, 33(2):333–341, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529783. See correc-

tions [988].

Muenz:1977:AMH

[1083] Larry R. Muenz, Sylvan B. Green, and David P. Byar. Applications of the Mantel–Haenszel statistic to the comparison of survival distributions. Biometrics, 33(4):617–626, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529459.

Park:1977:SPT

[1084] Sung H. Park. Selection of polynomial terms for response surface experiments. Biometrics, 33(1):225–229, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529316.

Pieters:1977:SSC

[1085] E. P. Pieters, C. E. Gates, J. H. Matis, and W. L. Sterling. Small sample comparison of different estimators of negative binomial parameters. Biometrics, 33(4):718–723, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529470.

Please:1977:BRM

[1086] N. W. Please. Book review: Methods for Statistical Data Analysis of Multivariate Observations, by R. Gnanadesikan. Biometrics, 33(2):429, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529800.

Robson:1977:QES

[1087] D. S. Robson and V. L. Wright. Query: Estimation of season total number of different households utilizing a park. Biometrics, 33(2):421– 425, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529798.

Routledge:1977:NJC

[1088] Garry L. Routledge and A. L. Jensen. A note on Jensen’s comparison of logistic equations. Biometrics, 33(3):574–575, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529380.

Rutledge:1977:RTT

[1089] J. J. Rutledge. Repeatability of threshold traits. Biometrics, 33(2):395– 399, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529791.

Salsburg:1977:OQT

[1090] David S. Salsburg. Open query: Theoretical problems in the modified mantel-bryan procedure. Biometrics, 33(2):419–421, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529797.

Schlesselman:1977:EED

[1091] James J. Schlesselman. The effect of errors of diagnosis and frequency of examination on reported rates of disease. Biometrics, 33(4):635–642, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529461.

Schreuder:1977:UBD

[1092] H. T. Schreuder and W. L. Hafley. A useful bivariate distribution for describing stand structure of tree heights and diameters. Biometrics, 33(3):471–478, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529361.

Sen:1977:CUR

[1093] A. R. Sen, S. Sellers, and G. E. J. Smith. Corrections to: “The use of a ratio estimate in successive sampling” (Biometrics 31 (1975), no. 3, 673–683). Biometrics, 33(4):767, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844].

Shaffer:1977:MCE

[1094] Juliet Popper Shaffer. Multiple comparisons emphasizing selected contrasts: an extension and generalization of Dunnett’s procedure. Biometrics, 33(2):293–303, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529779.

Shirley:1977:NPE

[1095] Eryl Shirley. A non-parametric equivalent of Williams’ test for contrasting increasing dose levels of a treatment. Biometrics, 33(2):386–389, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2529789.

Simon:1977:ATA

[1096] Richard Simon. Adaptive treatment assignment methods and clinical trials. Biometrics, 33(4):743–749, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529473.

Smeach:1977:FAM

[1097] Stephen C. Smeach and Robert W. Jernigan. Further aspects of a Markovian sampling policy for water quality monitoring. Biometrics, 33(1):41– 46, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529301.

Smith:1977:SPF

[1098] Woollcott Smith and J. Frederick Grassle. Sampling properties of a family of diversity measures. Biometrics, 33(2):283–292, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529778.

Somes:1977:SSW

[1099] Grant W. Somes and V. P. Bhapkar. A simulation study on a Wald statistic and Cochran’s Q statistic for stratified samples. Biometrics, 33(4):643–651, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529462.

Stene:1977:ADA

[1100] Jon Stene. Assumptions for different ascertainment models in human genetics. Biometrics, 33(3):523–527, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529367.

Thompson:1977:EHUa

[1101] Robin Thompson. The estimation of heritability with unbalanced data. I. Observations available on parents and offspring. Biometrics, 33(3):485– 495, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529363.

Thompson:1977:EHUb

[1102] Robin Thompson. The estimation of heritability with unbalanced data. II. Data available on more than two generations. Biometrics, 33(3):497– 504, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529364.

Thompson:1977:TGO

[1103] W. A. Thompson, Jr. On the treatment of grouped observations in life studies. Biometrics, 33(3):463–470, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529360.

Wahl:1977:DFW

[1104] Patricia W. Wahl and Richard A. Kronmal. Discriminant functions when covariances are unequal and sample sizes are moderate. Biometrics, 33(3):479–484, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529362.

Wallenstein:1977:ATP

[1105] Sylvan Wallenstein and Alan C. Fisher. The analysis of the two-period repeated measurements crossover design with application to clinical trials. Biometrics, 33(1):261–269, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529321. See correction [1418].

Warde:1977:RGG

[1106] William D. Warde, James M. Norton, and F. James Rohlf. On Rohf’s generalization of the gap test for the detection of multivariate outliers. Biometrics, 33(4):762–765, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529479.

Ware:1977:CRA

[1107] James H. Ware and Robert E. Bowden. Circadian rhythm analysis when output is collected at intervals. Biometrics, 33(3):566–571, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529378.

Williams:1977:ETR

[1108] E. R. Williams, H. D. Patterson, and J. A. John. Efficient two-replicate resolvable designs. Biometrics, 33(4):713–717, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529469.

Williams:1977:NRL

[1109] E. R. Williams. A note on rectangular lattice designs. Biometrics, 33(2): 410–414, June 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529794.

Williams:1977:QTC

[1110] James S. Williams. Quick test for comparing two populations with bivariate data. Biometrics, 33(1):215–219, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529314.

Woodbury:1977:EST

[1111] Max A. Woodbury, Kenneth G. Manton, and Lowell A. Woodbury. An extension of the sign test for replicated measurements. Biometrics, 33(3):453–461, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529359.

Woodward:1977:BSC

[1112] Wayne A. Woodward and William R. Schucany. Bibliography for statistical consulting. Biometrics, 33(3):564–565, September 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529377.

Woodward:1977:CPP

[1113] Wayne A. Woodward and William R. Schucany. Combination of a preference pattern with the triangle taste test. Biometrics, 33(1):31–39, March 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529300.

Worsley:1977:NPE

[1114] K. J. Worsley. A non-parametric extension of a cluster analysis method by Scott and Knott. Biometrics, 33(3):532–535, September 1977. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529369.

Zahl:1977:CTM

[1115] Samuel Zahl. A comparison of three methods for the analysis of spatial pattern. Biometrics, 33(4):681–692, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529466.

Zerbe:1977:RTC

[1116] Gary O. Zerbe and Strother H. Walker. A randomization test for comparison of groups of growth curves with different polynomial design matrices. Biometrics, 33(4):653–657, December 1977. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529463.

Anonymous:1978:Aa

[1117] Anonymous. Abstracts. Biometrics, 34(1):149–173, March 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529607.

Anonymous:1978:Ab

[1118] Anonymous. Abstracts. Biometrics, 34(2):335–336, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530032.

Anonymous:1978:Ac

[1119] Anonymous. Abstracts. Biometrics, 34(3):529–536, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530631.

Anonymous:1978:Ad

[1120] Anonymous. Abstracts. Biometrics, 34(4):725–757, December 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530407.

Anonymous:1978:BRA

[1121] Anonymous. Book review: Applications of Statistics by P. R. Krishnaiah. Biometrics, 34(3):528, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530627.

Anonymous:1978:BRF

[1122] Anonymous. Book review: Fixed Effects Analysis of Variance by L. Fisher, J. McDonald. Biometrics, 34(4):724, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530406.

Anonymous:1978:BRP

[1123] Anonymous. Book review: Probability and Statistics for Engineers and Scientists, by R. E. Walpole, R. H. Myers. Biometrics, 34(3):528, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530629.

Anonymous:1978:BRS

[1124] Anonymous. Book review: Selected Tables in Mathematical Statistics, Volume 5, by B. D. Owen, R. E. Odeh. Biometrics, 34(3):528, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530630.

Anonymous:1978:BRU

[1125] Anonymous. Book review: Urn Models and Their Application: an Approach to Modern Discrete Probability Theory by N. L. Johnson, S. Kotz.

Biometrics, 34(3):528, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530628.

Anonymous:1978:BSa

[1126] Anonymous. The Biometric Society. Biometrics, 34(2):337, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530033.

Anonymous:1978:BSb

[1127] Anonymous. The Biometric Society. Biometrics, 34(3):537, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530632.

Anonymous:1978:CDE

[1128] Anonymous. Corrections: Design of experiments to estimate heritability when observations are available on parents and offspring. Biometrics, 34 (2):329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530025. See

[966].

Anonymous:1978:CEV

[1129] Anonymous. Corrections: The estimation of variance and covariance components with an application when records are subject to culling. Biometrics, 34(2):328, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530023. See [591].

Anonymous:1978:CIS

[1130] Anonymous. Corrections: An improved solution for the ultimate probability of fixation of a favorable allele. Biometrics, 34(2):329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530026. See [1016].

Anonymous:1978:CLT

[1131] Anonymous. Corrections: On the line transect sampling method. Biometrics, 34(2):328–329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530024. See [702].

Anonymous:1978:FMa

[1132] Anonymous. Front matter. Biometrics, 34(1):??, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529583.

Anonymous:1978:FMb

[1133] Anonymous. Front matter. Biometrics, 34(2):i–iii, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530007.

Anonymous:1978:FMc

[1134] Anonymous. Front matter. Biometrics, 34(3):i–iv, September 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530596.

Anonymous:1978:FMd

[1135] Anonymous. Front matter. Biometrics, 34(4):i–iv, December 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530373.

Anonymous:1978:GMC

[1136] Anonymous. Gertrude Mary Cox — 1900–1978. Biometrics, 34(4):719– 720, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1978:NAa

[1137] Anonymous. News and announcements. Biometrics, 34(1):175–176, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529609.

Anonymous:1978:NAb

[1138] Anonymous. News and announcements. Biometrics, 34(2):338–339, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530034.

Anonymous:1978:NAc

[1139] Anonymous. News and announcements. Biometrics, 34(3):538–539, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530633.

Anonymous:1978:NAd

[1140] Anonymous. News and announcements. Biometrics, 34(4):759, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530408.

Anonymous:1978:OGM

[1141] Anonymous. Obituary: Gertrude Mary Cox 1900–1978. Biometrics, 34(4):719–720, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530397.

Anonymous:1978:OPD

[1142] Anonymous. Obituary: Prof. Dr. Ing. Václav Myslivec. Biometrics, 34 (1):145, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529603.

Anonymous:1978:PPBa

[1143] Anonymous. Papers to be published in Biometrics. Biometrics, 34(1): 177, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529610.

Anonymous:1978:PPBb

[1144] Anonymous. Papers to be published in Biometrics. Biometrics, 34(2): 340, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530035.

Anonymous:1978:PPBc

[1145] Anonymous. Papers to be published in Biometrics. Biometrics, 34(3): 540, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530634.

Anonymous:1978:PPBd

[1146] Anonymous. Papers to be published in Biometrics. Biometrics, 34(4): 761, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530409.

Anonymous:1978:VI

[1147] Anonymous. Volume information. Biometrics, 34(4):ix–xl, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530372.

Bailey:1978:MAA

[1148] R. C. Bailey, J. P. Summe, L. D. Homer, and L. E. McCracken. A model for analysis of the anesthetic response. Biometrics, 34(2):223– 232, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530012.

Bansal:1978:SMC

[1149] B. N. L. Bansal and P. C. Gupta. A stochastic model for cell survival after irradiation. Biometrics, 34(4):653–658, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530384.

Barrett:1978:BRA

[1150] John C. Barrett. Book review: Applied Mathematic Demography, by Nathan Keyfitz. Biometrics, 34(1):147, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529604.

Bartlett:1978:SMW

[1151] N. R. Bartlett. A survival model for a wood preservative trial. Biometrics, 34(4):673–679, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530388.

Bauer:1978:CTW

[1152] D. F. Bauer. Circular triads when not all paired comparisons are made. Biometrics, 34(3):458–461, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530608.

Benedetti:1978:SSL

[1153] Jacqueline K. Benedetti and Morton B. Brown. Strategies for the selection of log-linear models. Biometrics, 34(4):680–686, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530389.

Bishop:1978:BRG

[1154] J. A. Bishop. Book review: Geographic Variation, Speciation and Clines. Monographs in Population Biology, Volume 10, by J. A. Endler. Biometrics, 34(1):148, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529606.

Bjornsson:1978:ASL

[1155] Hólmgeir Björnsson. Analysis of a series of long-term grassland experiments with autocorrelated errors. Biometrics, 34(4):645–651, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530383.

Breslow:1978:TTL

[1156] N. Breslow and W. Powers. Are there two logistic regressions for retrospective studies? Biometrics, 34(1):100–105, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529594.

Cannings:1978:BRS

[1157] C. R. Cannings. Book review: Selection in One-and Two-Locus Systems, by T. Nagylaki. Biometrics, 34(3):525, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530621.

Carle:1978:NME

[1158] Frank Louis Carle and Mike R. Strub. A new method for estimating population size from removal data. Biometrics, 34(4):621–630, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530381.

Casagrande:1978:IAF

[1159] J. T. Casagrande, M. C. Pike, and P. G. Smith. An improved approximate formula for calculating sample sizes for comparing two binomial distributions. Biometrics, 34(3):483–486, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530613.

Cash:1978:CIS

[1160] W. S. Cash. Corrections to: “An improved solution for the ultimate probability of fixation of a favorable allele” [Biometrics 33 (1977), no. 3, 528–532; MR 58 #33109]. Biometrics, 34(2):329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1016].

Cohen:1978:QAL

[1161] Joel E. Cohen and Peter D’eustachio. Query: an affine linear model for the relation between two sets of frequency counts. Biometrics, 34(3):514– 521, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530617.

Cormack:1978:BS

[1162] R. M. Cormack. Biometric society. Biometrics, 34(1):174, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529608.

Cornell:1978:DAB

[1163] J. A. Cornell and J. W. Gorman. On the detection of an additive blending component in multicomponent mixtures. Biometrics, 34(2):251–263, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530015.

Cressie:1978:RNT

[1164] N. A. C. Cressie. Removing nonadditivity from two-way tables with one observation per cell. Biometrics, 34(3):505–513, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530616.

Curnow:1978:SWS

[1165] R. N. Curnow. Selection within self-fertilizing populations. Biometrics, 34(4):603–610, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530379.

Dayananda:1978:QGC

[1166] P. W. A. Dayananda. Quantitative guidelines for communicable disease control programs. Biometrics, 34(1):142, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529601.

Dolby:1978:BRC

[1167] Geoffrey R. Dolby and Oscar Kempthorne. Bertrand Russell and causality. Biometrics, 34(4):716–718, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530396.

Dupont:1978:MCI

[1168] William D. Dupont and Oscar Kempthorne. Making causal inferences from observational data. Biometrics, 34(4):713–714, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530394.

Eccleston:1978:VCD

[1169] J. A. Eccleston. Variance components and disconnected data. Biometrics, 34(3):479–482, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530612.

Edwards:1978:BRS

[1170] A. W. F. Edwards. Book review: Stochastic Problems in Population Genetics, by T. Maruyama. Biometrics, 34(3):528, September 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530626.

Elston:1978:BRR

[1171] R. C. Elston. Book review: Recent Developments in Statistics, by J. R. Barra, F. Brodeau, G. Romier, B. Van Cutsem. Biometrics, 34(4):723, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530403.

Everitt:1978:MPO

[1172] B. S. Everitt and D. N. Rushton. A method for plotting the optimum positions of an array of cortical electrical phosphenes. Biometrics, 34(3):399–410, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530602.

Feigl:1978:GAD

[1173] Polly Feigl. A graphical aid for determining sample size when comparing two independent proportions. Biometrics, 34(1):111–122, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529596.

Finney:1978:BRB

[1174] D. J. Finney. Book review: Bioassay of Pesticides in the Laboratory, by L. Bánki. Biometrics, 34(2):331, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530027.

Finney:1978:TEI

[1175] D. J. Finney, Samuel Zahl, Gerald van Belle, and Theodore Zeisig. Testing the effect of an intervention in sequential ecological data. Biometrics, 34(4):706–708, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530392.

Ford:1978:BRD

[1176] E. D. Ford. Book review: Diffusion Processes and Related Topics in Biology, by L. M. Ricciardi. Biometrics, 34(2):332, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530029.

Ford:1978:CAS

[1177] B. L. Ford and R. D. Tortora. A consulting aid to sample design. Biometrics, 34(2):299–304, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530020.

Forsythe:1978:QEM

[1178] A. B. Forsythe. Query: Expressing the mean of rounded data. Biometrics, 34(1):138, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529599.

Freeman:1978:BRS

[1179] G. H. Freeman. Book review: Statistical Procedures for Agricultural Research with Emphasis on Rice, by K. A. Gomez, A. A. Gomez. Biometrics, 34(4):721–722, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530400.

Fuller:1978:QES

[1180] W. A. Fuller and J. J. Goebel. Query: On the estimation of season total number of different households utilizing a park. Biometrics, 34(1):139– 141, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529600.

Gillis:1978:BEP

[1181] P. R. Gillis and D. A. Ratkowsky. The behaviour of estimators of the parameters of various yield-density relationships. Biometrics, 34(2):191– 198, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530009.

Goldberg:1978:EFN

[1182] J. D. Goldberg and J. T. Wittes. The estimation of false negatives in medical screening. Biometrics, 34(1):77–86, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529590.

Hanley:1978:LCG

[1183] J. A. Hanley. A language for computer generation of medical data forms. Biometrics, 34(2):288–297, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530019.

Hartley:1978:SSB

[1184] H. O. Hartley, J. N. K. Rao, and L. R. Lamotte. A simple “synthesis”- based method of variance component estimation. Biometrics, 34(2):233– 242, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530013.

Harville:1978:AFP

[1185] David A. Harville. Alternative formulations and procedures for the twoway mixed model. Biometrics, 34(3):441–453, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530606.

Haseman:1978:ESS

[1186] J. K. Haseman. Exact sample sizes for use with the fisher-irwin test for 2 × 2 tables. Biometrics, 34(1):106–109, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529595.

Hawkes:1978:BRS

[1187] A. G. Hawkes. Book review: Stochastic Models for Spike Trains of Single Neurons, by G. Sampath, S. K. Spinivasan. Biometrics, 34(3):525, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530620.

Healy:1978:NMC

[1188] M. J. R. Healy and Richard Simon. New methodology in clinical trials. Biometrics, 34(4):709–712, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530393.

Henderson:1978:VCM

[1189] C. R. Henderson. Variance-covariance matrix of estimators of variances in unweighted means ANOVA. Biometrics, 34(3):462–468, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530609.

Hill:1978:BRB

[1190] W. G. Hill. Book review: Biometrical Methods in Quantitative Genetic Analysis, by R. K. Singh, B. D. Chaudhary. Biometrics, 34(4):723, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530404.

Hill:1978:BRM

[1191] W. G. Hill. Book review: Measuring Selection in Natural Populations, by F. B. Christiansen, T. M. Fenchel. Biometrics, 34(3):527–528, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530625.

Hill:1978:PNP

[1192] W. G. Hill and R. Thompson. Probabilities of non-positive definite between-group or genetic covariance matrices. Biometrics, 34(3):429– 439, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530605.

Holford:1978:APM

[1193] T. R. Holford. The analysis of pair-matched case-control studies, a multivariate approach. Biometrics, 34(4):665–672, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530387.

James:1978:EMP

[1194] I. R. James. Estimation of the mixing proportion in a mixture of two normal distributions from simple, rapid measurements. Biometrics, 34(2): 265–275, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530016.

Johnson:1978:BRM

[1195] A. L. Johnson. Book review: Medical Diagnostic Models. A Bibliography, by M. Clinton Miller, M. C. Westphal, J. Routt Reigart, Carol Barner. Biometrics, 34(4):724, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530405.

Katz:1978:OCI

[1196] D. Katz, J. Baptista, S. P. Azen, and M. C. Pike. Obtaining confidence intervals for the risk ratio in cohort studies. Biometrics, 34(3):469–474, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530610.

Kempthorne:1978:BIP

[1197] Oscar Kempthorne. A biometrics invited paper: Logical, epistemological and statistical aspects of nature-nurture data interpretation. Biometrics, 34(1):1–23, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529584.

Kempthorne:1978:LES

[1198] Oscar Kempthorne. Logical, epistemological and statistical aspects of nature-nurture data interpretation. Biometrics, 34(1):1–23, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempton:1978:CTM

[1199] R. A. Kempton and R. W. M. Wedderburn. A comparison of three measures of species diversity. Biometrics, 34(1):25–37, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529585.

Kerridge:1978:BRP

[1200] D. Kerridge. Book review: The Population Biology of Disease, by John H. Whitlock. Biometrics, 34(4):722–723, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530402.

Klotz:1978:MEC

[1201] J. H. Klotz. Maximum entropy constrained balance randomization for clinical trials. Biometrics, 34(2):283–287, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530018.

Krippendorff:1978:RBA

[1202] Klaus Krippendorff and Joseph L. Fleiss. Reliability of binary attribute data. Biometrics, 34(1):142–144, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2529602. With a reply by Fleiss.

Kupper:1978:UCB

[1203] L. L. Kupper and J. K. Haseman. The use of a correlated binomial model for the analysis of certain toxicological experiments. Biometrics, 34(1): 69–76, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529589.

Lawless:1978:ESN

[1204] J. F. Lawless and K. Singhal. Efficient screening of nonnormal regression models. Biometrics, 34(2):318–327, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530022.

Layard:1978:APD

[1205] M. W. J. Layard and J. N. Arvesen. Analysis of Poisson data in crossover experimental designs. Biometrics, 34(3):421–428, September 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530604.

Matthews:1978:BRC

[1206] J. A. Matthews. Book review: Calculus for the Life Sciences, by R. Sapio. Biometrics, 34(4):722, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530401.

Mead:1978:E

[1207] Roger Mead and Samuel Zahl. [Editorial]. Biometrics, 34(4):714–716, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530395.

Mielke:1978:CAI

[1208] P. W. Mielke. Clarification and appropriate inferences for Mantel and Valand’s nonparametric multivariate analysis technique. Biometrics, 34(2): 277–282, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530017.

Mosimann:1978:SSA

[1209] J. E. Mosimann, J. D. Malley, A. W. Cheever, and C. B. Clark. Size and shape analysis of schistosome egg-counts in Egyptian autopsy data. Biometrics, 34(3):341–356, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530597.

Murray:1978:BRS

[1210] J. D. Murray. Book review: Scale Effects in Animal Locomotion, by T. J. Pedley. Biometrics, 34(2):331–332, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530028.

Neal:1978:BRM

[1211] D. G. Neal. Book review: Mathematics, Statistics and Systems for Health, by N. T. J. Bailey. Biometrics, 34(3):527, September 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530623.

Norwood:1978:MAB

[1212] Paula K. Norwood and Klaus Hinkelmann. Measures of association between disease and genotype. Biometrics, 34(4):593–602, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530378.

OBrien:1978:NTA

[1213] Peter C. O’Brien. A nonparametric test for association with censored data. Biometrics, 34(2):243–250, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530014.

OFallon:1978:STS

[1214] J. R. O’Fallon, S. D. Dubey, D. S. Salsburg, J. H. Edmonson, A. Soffer, and T. Colton. Should there be statistical guidelines for medical research papers? Biometrics, 34(4):687–695, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530390.

Oldham:1978:BRC

[1215] P. D. Oldham. Book review: Clinical Biostatistics, by A. R. Feinstein. Biometrics, 34(3):525–526, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530622.

ONeill:1978:SPU

[1216] R. T. O’Neill and C. W. Chen. A statistical procedure useful in evaluating time of onset and termination of response in clinical trials. Biometrics, 34(3):411–420, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530603.

Patil:1978:WDS

[1217] G. P. Patil and C. R. Rao. Weighted distributions and size-biased sampling with applications to wildlife populations and human families. Biometrics, 34(2):179–189, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530008.

Pollock:1978:FDE

[1218] K. H. Pollock. A family of density estimators for line-transect sampling. Biometrics, 34(3):475–478, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530611.

Prentice:1978:AFT

[1219] R. L. Prentice, J. D. Kalbfleisch, A. V. Peterson, Jr., N. Flournoy, V. T. Farewell, and N. E. Breslow. The analysis of failure times in the presence of competing risks. Biometrics, 34(4):541–554, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530374.

Prentice:1978:RAG

[1220] R. L. Prentice and L. A. Gloeckler. Regression analysis of grouped survival data with application to breast cancer data. Biometrics, 34(1): 57–67, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529588.

Raghavarao:1978:EPW

[1221] D. Raghavarao. On an estimation problem in Warner’s randomized response technique. Biometrics, 34(1):87–90, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529591.

Rao:1978:SMM

[1222] Chennupati R. Rao and A. M. Kshirsagar. A semi-markovian model for predator-prey interactions. Biometrics, 34(4):611–619, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530380.

Russell:1978:EPT

[1223] K. G. Russell. Estimation of the parameters of the Thomas distribution. Biometrics, 34(1):95–99, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529593.

Sampford:1978:BRS

[1224] M. R. Sampford. Book review: Sampling Techniques, by W. G. Cochran. Biometrics, 34(2):332–333, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530031.

Sandvik:1978:EIM

[1225] Leiv Sandvik. Erratum: “Inverse, multiple and sequential sample censuses” (Biometrics 8 (1952), 286–306) by D. G. Chapman. Biometrics, 34(3):523, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Sandvik:1978:NDM

[1226] Leiv Sandvik. A note on direct multiple sample censuses. Biometrics, 34(3):523, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530619.

Schaeffer:1978:SEV

[1227] L. R. Schaeffer, J. W. Wilton, and Robin Thompson. Simultaneous estimation of variance and covariance components from multitrait mixed model equations. Biometrics, 34(2):199–208, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530010.

Seheult:1978:NSS

[1228] Allan Seheult. A note on stand structure of trees. Biometrics, 34(3):522– 523, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530618.

Sen:1978:CLT

[1229] A. R. Sen, J. Tourigny, and G. E. J. Smith. Corrections to: “On the line transect sampling method” (Biometrics 30 (1974), 329–340). Biometrics, 34(2):328–329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [702].

Sinclair:1978:BRP

[1230] C. D. Sinclair. Book review: Prediction and Improved Estimation in Linear Models, by J. Bibby, H. Toutenburg. Biometrics, 34(3):527, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530624.

Smith:1978:BRI

[1231] R. H. Smith. Book review: Integrodifferential Equations and Delay Models in Population Dynamics, by J. M. Cushing. Biometrics, 34(4):721, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530399.

Smith:1978:BRP

[1232] C. Smith. Book review: Proceedings of the International Conference on Quantitative Genetics, by E. Pollak, O. Kempthorne, T. B. Bailey Jr. Biometrics, 34(2):332, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530030.

Soller:1978:EED

[1233] M. Soller and A. Genizi. The efficiency of experimental designs for the detection of linkage between a marker locus and a locus affecting a quantitative trait in segregating populations. Biometrics, 34(1):47–55, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529587.

Spurr:1978:BRA

[1234] B. D. Spurr. Book review: Applied Nonparametric Statistics, by W. W. Daniel. Biometrics, 34(4):721, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530398.

Stanish:1978:AMR

[1235] W. M. Stanish, D. B. Gillings, and G. G. Koch. An application of multivariate ratio methods for the analysis of a longitudinal clinical trial with missing data. Biometrics, 34(2):305–317, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530021.

Stene:1978:AMH

[1236] J. Stene. Ascertainment models for human families selected through several affected children. Biometrics, 34(3):455–457, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530607.

Tallis:1978:EAM

[1237] G. M. Tallis, P. Leppard, and G. Sarfaty. The effect of accessing medical records by date of death on estimates of survival. Biometrics, 34(4):662– 664, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530386.

Temkin:1978:ATS

[1238] Nancy R. Temkin. An analysis for transient states with application to tumor shrinkage. Biometrics, 34(4):571–580, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530376.

Theobald:1978:CCL

[1239] C. M. Theobald and J. R. Mallinson. Comparative calibration, linear structural relationships and congeneric measurements. Biometrics, 34(1): 39–45, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529586.

Thompson:1978:CDE

[1240] R. Thompson. Corrections to: “Design of experiments to estimate heritability when observations are available on parents and offspring” (Biometrics 32 (1976), no. 2, 283–304). Biometrics, 34(2):329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [966].

Thompson:1978:CEV

[1241] R. Thompson. Corrections to: “The estimation of variance and covariance components with an application when records are subject to culling” (Biometrics 29 (1973), 527–550). Biometrics, 34(2):328–329, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [591].

Tolley:1978:CMA

[1242] H. D. Tolley, D. Burdick, K. G. Manton, and E. Stallard. A compartment model approach to the estimation of tumor incidence and growth: Investigation of a model of cancer latency. Biometrics, 34(3):377–389, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530600.

Tolley:1978:LMA

[1243] H. Dennis Tolley, Kenneth G. Manton, and Sharon Sandomirsky Poss. A linear models application of competing risks to multiple causes of death. Biometrics, 34(4):581–591, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530377.

Tomassone:1978:BRT

[1244] R. Tomassone. Book review: Technique de la Description Statistique, by L. Lebart, A. Morineau, N. Tabard. Biometrics, 34(1):147–148, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529605.

Tsutakawa:1978:CTR

[1245] R. K. Tsutakawa and J. E. Hewett. Comparison of two regression lines over a finite interval. Biometrics, 34(3):391–398, September 1978. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530601. See correction [1276].

Turk:1978:QTE

[1246] Ted R. Turk, S. Zahl, G. Van Belle, and T. Zeisig. Query: Testing the effect of an intervention in sequential ecological data. Biometrics, 34(1): 128–137, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529598.

Turnbull:1978:EAD

[1247] Bruce W. Turnbull and Toby J. Mitchell. Exploratory analysis of disease prevalence data from Survival/Sacrifice experiments. Biometrics, 34(4):555–570, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530375.

Turnbull:1978:LRS

[1248] B. W. Turnbull and L. Weiss. A likelihood ratio statistic for testing goodness of fit with randomly censored data. Biometrics, 34(3):367–375, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530599.

Urquhart:1978:LMM

[1249] N. Scott Urquhart and David L. Weeks. Linear models in messy data: Some problems and alternatives. Biometrics, 34(4):696–705, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530391.

VanVleck:1978:GMI

[1250] L. D. Van Vleck. A genetic model involving fetal effects on traits of the dam. Biometrics, 34(1):123–127, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529597.

Volund:1978:AFP

[1251] Aage Vølund. Application of the four-parameter logistic model to bioassay: Comparison with slope ratio and parallel line models. Biometrics, 34(3):357–365, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530598.

Walters:1978:AVH

[1252] D. E. Walters and J. R. Morton. On the analysis of variance of a half diallel table. Biometrics, 34(1):91–94, March 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529592.

Whittaker:1978:FSF

[1253] J. Whittaker and M. Aitkin. A flexible strategy for fitting complex loglinear models. Biometrics, 34(3):487–495, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530614.

Williams:1978:EAW

[1254] J. S. Williams. Efficient analysis of Weibull survival data from experiments on heterogeneous patient populations. Biometrics, 34(2):209–222, June 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530011.

Williams:1978:ECR

[1255] D. A. Williams. An exact confidence region for a relative potency estimated from combined bioassays. Biometrics, 34(4):659–661, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530385.

Williford:1978:EMC

[1256] William O. Williford and Anthony L. Price. Estimation in modified and compound modified discrete distributions. Biometrics, 34(4):631–643, December 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530382.

Wood:1978:CLT

[1257] Constance L. Wood. Comparison of linear trends in binomial proportions. Biometrics, 34(3):496–504, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530615.

Aleong:1979:IGC

[1258] John Aleong and Diantha E. Bartlett. Improved graphs for calculating sample sizes when comparing two independent binomial distributions. Biometrics, 35(4):875–881, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530123.

Alexander:1979:BRF

[1259] J. R. Alexander. Book review: Fetus into Man, by J. M. Tanner. Biometrics, 35(3):707, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530274.

Altham:1979:BRA

[1260] P. M. E. Altham. Book review: The Analysis of Cross-Tabulated Data, by G. J. G. Upton. Biometrics, 35(3):707–708, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530276.

Anderson:1979:BRD

[1261] J. A. Anderson. Book review: Discrete Discriminant Analysis, by M. Goldstein, W. R. Dillon. Biometrics, 35(2):529–530, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530363.

Anderson:1979:BRI

[1262] D. R. Anderson. Book review: Investigating Animal Abundance: Capture-Recapture for Biologists, by M. Begon. Biometrics, 35(4):901, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530126.

Anderson:1979:GMC

[1263] R. L. Anderson, R. J. Monroe, and L. A. Nelson. Gertrude M. Cox — a modern pioneer in statistics. Biometrics, 35(1):3–7, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529931.

Andrich:1979:MCT

[1264] David Andrich. A model for contingency tables having an ordered response classification. Biometrics, 35(2):403–415, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530343.

Anonymous:1979:Aa

[1265] Anonymous. Abstracts. Biometrics, 35(2):525, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530357.

Anonymous:1979:Ab

[1266] Anonymous. Abstracts. Biometrics, 35(3):695–700, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530263.

Anonymous:1979:Ac

[1267] Anonymous. Abstracts. Biometrics, 35(4):883–896, December 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530124.

Anonymous:1979:BRC

[1268] Anonymous. Book review: Contributions to Survey Sampling and Applied Statistics: Papers in Honor of H. O. Hartley by H. A. David. Biometrics, 35(4):904, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530132.

Anonymous:1979:BRI

[1269] Anonymous. Book review: Introduction to Statistics: a Non-parametric Approach for the Social Sciences by C. Leach. Biometrics, 35(4):904, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530134.

Anonymous:1979:BRM

[1270] Anonymous. Book review: Mathematics: an Introduction to Its Spirit and Use — Readings from Scientific American, by M. Kline. Biometrics, 35(3):708–709, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530278.

Anonymous:1979:BRTa

[1271] Anonymous. Book review: Theoretical Chemistry, Periodicities in Chemistry and Biology by H. Eyring, D. Henderson. Biometrics, 35(3):708, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530277.

Anonymous:1979:BSa

[1272] Anonymous. The Biometric Society. Biometrics, 35(1):360, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529966.

Anonymous:1979:BSb

[1273] Anonymous. The Biometric Society. Biometrics, 35(2):533–535, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530368.

Anonymous:1979:BSc

[1274] Anonymous. The Biometric Society. Biometrics, 35(3):701, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530264.

Anonymous:1979:BSF

[1275] Anonymous. The Biometric Society financial statements and auditors’ report January 31, 1979. Biometrics, 35(4):897–899, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530125.

Anonymous:1979:CCT

[1276] Anonymous. Correction: Comparison of two regression lines over a finite interval. Biometrics, 35(4):905, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530137. See [1245].

Anonymous:1979:CMV

[1277] Anonymous. Correction: Missing values in partial diallel cross experiments. Biometrics, 35(4):905, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530135. See [5].

Anonymous:1979:CTD

[1278] Anonymous. Correction: Testing to determine the underlying distribution using randomly censored data. Biometrics, 35(4):905, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530136. See [1340].

Anonymous:1979:BRD

[1279] Anonymous. Book review: Developments in Statistics, by P. R. Krishnaiah. Biometrics, 35(3):709, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530279.

Anonymous:1979:ENE

[1280] Anonymous. Editor’s note [erratum]: R. A. Fisher — the life of a scientist. Biometrics, 35(3):703, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530265. See [1326].

Anonymous:1979:FMa

[1281] Anonymous. Front matter. Biometrics, 35(1):i–iv, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529928.

Anonymous:1979:FMb

[1282] Anonymous. Front matter. Biometrics, 35(2):i–iv, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530339.

Anonymous:1979:FMc

[1283] Anonymous. Front matter. Biometrics, 35(3):i–iv, September 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530243.

Anonymous:1979:FMd

[1284] Anonymous. Front matter. Biometrics, 35(4):i–iv, December 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530101.

Anonymous:1979:MGM

[1285] Anonymous. In memoriam: Gertrude M. Cox, 1900–1978. Biometrics, 35 (1):v–vi, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529929.

Anonymous:1979:NAa

[1286] Anonymous. News and announcements. Biometrics, 35(1):361, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529967.

Anonymous:1979:NAb

[1287] Anonymous. News and announcements. Biometrics, 35(2):537, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530369.

Anonymous:1979:NAc

[1288] Anonymous. News and announcements. Biometrics, 35(3):711–712, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530280.

Anonymous:1979:NAd

[1289] Anonymous. News and announcements. Biometrics, 35(4):913–914, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530142.

Anonymous:1979:PPBa

[1290] Anonymous. Papers to be published in Biometrics. Biometrics, 35(1): 363, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529968.

Anonymous:1979:PPBb

[1291] Anonymous. Papers to be published in Biometrics. Biometrics, 35(2): 538, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530370.

Anonymous:1979:PPBc

[1292] Anonymous. Papers to be published in Biometrics. Biometrics, 35(3): 713, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530281.

Anonymous:1979:BRTb

[1293] Anonymous. Book review: Theoretical Approaches to Complex Systems: Proceedings, Tübingen 1977 by R. Heim, G. Palm. Biometrics, 35(4):904, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530133.

Anonymous:1979:VI

[1294] Anonymous. Volume information. Biometrics, 35(4):v–xxxii, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530100.

Bartlett:1979:BRS

[1295] M. S. Bartlett. Book reviews: Studies in Mathematical Biology: Part I, Cellular Behaviour and the Development of Pattern, by S. A. Levin; Studies in Mathematical Biology: Part II, Populations and Communities, by S. A. Levin. Biometrics, 35(2):530–531, June 1979. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530365.

Beck:1979:SSM

[1296] Gerald J. Beck. Stochastic survival models with competing risks and covariates. Biometrics, 35(2):427–438, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530345.

Becker:1979:UEM

[1297] Niels Becker. The uses of epidemic models. Biometrics, 35(1):295–305, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529951.

Bhapkar:1979:TMS

[1298] Vasant P. Bhapkar. On tests of marginal symmetry and quasi-symmetry in two and three-dimensional contingency tables. Biometrics, 35(2):417– 426, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530344.

Bibby:1979:BRM

[1299] J. Bibby. Book review: Mathematical Tools for Applied Multivariate Analysis, by P. E. Green, J. Douglas Carroll. Biometrics, 35(3):704, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530268.

Billewicz:1979:BRL

[1300] W. Z. Billewicz. Book review: Life Table and Mortality Analysis, by C. L. Chiang. Biometrics, 35(1):356, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529958.

Blackith:1979:BRQ

[1301] R. E. Blackith. Book review: Quantitative Ethology, by P. Colgan. Biometrics, 35(1):355, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529956.

Boyle:1979:MGM

[1302] C. R. Boyle and Robert C. Elston. Multifactorial genetic models for quantitative traits in humans. Biometrics, 35(1):55–68, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529936.

Browne:1979:VAS

[1303] Richard H. Browne. On visual assessment of the significance of a mean difference. Biometrics, 35(3):657–665, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530259.

Burnham:1979:PGH

[1304] K. P. Burnham. A parametric generalization of the Hayne estimator for line transect sampling. Biometrics, 35(3):587–595, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530249.

Cady:1979:EFY

[1305] Foster B. Cady. Editor’s Five-Year Report. Biometrics, 35(4):915–917, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530143.

Cady:1979:PB

[1306] Foster B. Cady. Perspectives in biometry. Biometrics, 35(1):8, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529932.

Cochran:1979:CIB

[1307] William G. Cochran and David J. Finney. Chester Ittner Bliss (1899– 1979). Biometrics, 35(4):715–717, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530102.

Cochran:1979:SR

[1308] William G. Cochran. Some reflections. Biometrics, 35(1):1–2, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529930.

Collett:1979:BRG

[1309] D. Collett. Book review: The GLIM System Release 3, by R. J. Baker, J. A. Nelder. Biometrics, 35(2):527–528, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530359.

Colquhoun:1979:BRC

[1310] D. Colquhoun. Book review: Cooperative Equilibrium in Physical Biochemistry, by D. Poland. Biometrics, 35(3):706, September 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530272.

Colton:1979:BRL

[1311] T. Colton. Book review: Lecture Notes on Medical Statistics, by A. Petrie. Biometrics, 35(4):903–904, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530131.

Cook:1979:DEV

[1312] R. Dennis Cook and Jerald O. Jacobson. A design for estimating visibility bias in aerial surveys. Biometrics, 35(4):735–742, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530104.

Cooke:1979:BRB

[1313] D. Cooke. Book review: Biostatistics, by W. W. Daniel. Biometrics, 35 (1):357, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529960.

Cornelius:1979:CFR

[1314] P. L. Cornelius, W. C. Templeton, Jr., and T. H. Taylor. Curve fitting by regression on smoothed singular vectors. Biometrics, 35(4):849–859, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530119.

Crump:1979:DRP

[1315] Kenny S. Crump. Dose response problems in carcinogenesis. Biometrics, 35(1):157–167, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529942.

Curnow:1979:BRE

[1316] R. N. Curnow. Book review: Evolution and the Genetics of Populations, Volume 4: Variability within and among Natural Populations, by S. Wright. Biometrics, 35(1):359, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529965.

Day:1979:THC

[1317] N. E. Day and D. P. Byar. Testing hypotheses in case-control studies — equivalence of Mantel–Haenszel statistics and logit score tests. Biometrics, 35(3):623–630, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530253.

deVries:1979:GHT

[1318] P. G. de Vries. A generalization of the hayne-type estimator as an application of line intersect sampling. Biometrics, 35(4):743–748, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530105.

Diggle:1979:PEG

[1319] Peter J. Diggle. On parameter estimation and goodness-of-fit testing for spatial point patterns. Biometrics, 35(1):87–101, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529938.

Duncan:1979:ART

[1320] David B. Duncan and James H. Godbold. Approximate k-ratio t tests for differences between unequally replicated treatments. Biometrics, 35(4):749–756, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530106.

Ellerton:1979:MBE

[1321] Roger R. W. Ellerton and Wei-Yann Tsai. Minimum bias estimation and the selection of polynomial terms for response surfaces. Biometrics, 35(3):631–635, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530254.

Engen:1979:SAM

[1322] Steinar Engen. Stochastic abundance models in ecology. Biometrics, 35(1):331–338, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529954.

Evans:1979:ASO

[1323] J. C. Evans and E. A. Roberts. Analysis of sequential observations with applications to experiments on grazing animals and perennial plants. Biometrics, 35(3):687–693, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530262.

Everitt:1979:UPC

[1324] B. S. Everitt. Unresolved problems in cluster analysis. Biometrics, 35(1): 169–181, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529943.

Fienberg:1979:LLR

[1325] Stephen E. Fienberg. Log linear representation for paired comparison models with ties and within-pair order effects. Biometrics, 35(2):479– 481, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530350.

Finney:1979:BRL

[1326] D. J. Finney. Book review: The Life of a Scientist, by R. A. Fisher. Biometrics, 35(1):357–358, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2529962. See erratum [1280].

Gail:1979:COL

[1327] M. H. Gail and J. H. Ware. Comparing observed life table data with a known survival curve in the presence of random censorship. Biometrics, 35(2):385–391, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530341.

Geidel:1979:BRM

[1328] H. Geidel. Book review: Modelle zur Erfassung von Wanderungsbewegungen, by W. Schweitzer. Biometrics, 35(1):358, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529963.

Glaz:1979:PMA

[1329] J. Glaz. Probabilities and moments for absorption in finite homogeneous birth-death processes. Biometrics, 35(4):813–816, December 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530113.

Good:1979:DTE

[1330] Phillip I. Good. Detection of a treatment effect when not all experimental subjects will respond to treatment. Biometrics, 35(2):483–489, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530351. See correction

Goodman:1979:QIT

[1331] Leo A. Goodman. On quasi-independence in triangular contingency tables. Biometrics, 35(3):651–655, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530258.

Griffin:1979:BRC

[1332] Nicholas Griffin. Bertrand Russell and causality. Biometrics, 35(4):909– 911, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530140.

Habbema:1979:SMC

[1333] J. D. F. Habbema. Statistical methods for classification of human chromosomes. Biometrics, 35(1):103–118, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529939.

Haseman:1979:ADR

[1334] J. K. Haseman and L. L. Kupper. Analysis of dichotomous response data from certain toxicological experiments. Biometrics, 35(1):281–293, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529950.

Hauck:1979:LSV

[1335] Walter W. Hauck. The large sample variance of the Mantel–Haenszel estimator of a common odds ratio. Biometrics, 35(4):817–819, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530114.

Herson:1979:PPE

[1336] Jay Herson. Predictive probability early termination plans for phase II clinical trials. Biometrics, 35(4):775–783, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530109.

Hess:1979:SMR

[1337] James L. Hess. Sensitivity of MINQUE with respect to a priori weights. Biometrics, 35(3):645–649, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530257.

Holgate:1979:BRL

[1338] P. Holgate. Book review: Lectures on Nonlinear Differential-Equation Models in Biology, by J. D. Murray. Biometrics, 35(3):705–706, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2530271.

Hollander:1979:CTD

[1339] M. Hollander and F. Proschan. Correction to: “Testing to determine the underlying distribution using randomly censored data” [Biometrics 35 (1979), no. 2, 393–401; MR 81a:62034]. Biometrics, 35(4):905, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1340].

Hollander:1979:TDU

[1340] Myles Hollander and Frank Proschan. Testing to determine the underlying distribution using randomly censored data. Biometrics, 35(2):393– 401, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530342. See cor-

rection [1278, 1339].

Jolly:1979:BRP

[1341] G. M. Jolly. Book review: POPAN-2. A Data Maintenance and Analysis System for Mark-Recapture Data, by A. N. Arnason, L. Baniuk. Biometrics, 35(2):527, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530358.

Jupp:1979:BRM

[1342] P. E. Jupp. Book review: The Maximum Entropy Formalism, by R. D. Levine, M. Tribus. Biometrics, 35(4):902, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530128.

Kemp:1979:BRS

[1343] A. W. Kemp. Book review: Stochastic Abundance Models, by S. Engen. Biometrics, 35(2):529, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530362.

Kempton:1979:SSA

[1344] R. A. Kempton. The structure of species abundance and measurement of diversity. Biometrics, 35(1):307–321, March 1979. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529952.

Kilpatrick:1979:BRM

[1345] S. J. Kilpatrick, Jr. Book review: Mainland’s Elementary Medical Statistics, by Donald Mainland, C. M. Miller. Biometrics, 35(4):902, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530129.

Kirkwood:1979:GMM

[1346] Thomas B. L. Kirkwood. Geometric means and measures of dispersion. Biometrics, 35(4):908–909, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530139.

Knoke:1979:NAS

[1347] James D. Knoke. Normal approximations for serial correlation statistics. Biometrics, 35(2):491–495, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530352.

Korn:1979:MAP

[1348] Edward L. Korn and Alice S. Whittemore. Methods for analyzing panel studies of acute health effects of air pollution. Biometrics, 35(4):795–802, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530111.

Lachenbruch:1979:DA

[1349] P. A. Lachenbruch and M. Goldstein. Discriminant analysis. Biometrics, 35(1):69–85, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529937.

Lagakos:1979:GRC

[1350] S. W. Lagakos. General right censoring and its impact on the analysis of survival data. Biometrics, 35(1):139–156, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529941.

Laska:1979:SPU

[1351] Eugene M. Laska, Morris Meisner, Chao W. Chen, and Robert T. O’Neill. A statistical procedure useful in evaluating time of onset and termination of responses in clinical trials. Biometrics, 35(4):907–908, December 1979.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530138.

Lefkovitch:1979:BRG

[1352] L. P. Lefkovitch. Book review: Graphical Techniques for Multivariate Data, by B. S. Everitt. Biometrics, 35(1):357, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529961.

Lindley:1979:ALT

[1353] D. V. Lindley. Analysis of life tables with grouping and withdrawals. Biometrics, 35(3):605–612, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530251.

Marcus:1979:FPT

[1354] Allan H. Marcus and Stanley Czajkowski, Jr. First passage times as environmental safety indicators: Carboxyhemoglobin from cigarette smoke. Biometrics, 35(3):539–548, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530244.

Matis:1979:SMC

[1355] J. H. Matis and T. E. Wehrly. Stochastic models of compartmental systems. Biometrics, 35(1):199–220, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2529945.

Mead:1979:BRM

[1356] R. Mead. Book review: Mathematics and the Life Sciences, by D. E. Matthews. Biometrics, 35(3):705, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530270.

Mead:1979:CE

[1357] R. Mead. Competition experiments. Biometrics, 35(1):41–54, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529935.

Mitchell:1979:LLM

[1358] Toby J. Mitchell and Bruce W. Turnbull. Log-linear models in the analysis of disease prevalence data from survival/ sacrifice experiments. Biometrics, 35(1):221–234, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529946.

Mountford:1979:BRG

[1359] M. D. Mountford. Book review: Guide to the Study of Animal Populations, by James T. Tanner. Biometrics, 35(3):707, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530275.

Mueller:1979:CTM

[1360] Laurence D. Mueller. A comparison of two methods for making statistical inferences on Nei’s measure of genetic distance. Biometrics, 35(4):757– 763, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530107.

Mussett:1979:BRM

[1361] M. V. Mussett. Book review: Micro-Organisms in Foods 1: Their Significance and Methods of Enumeration, by ICMSF. Biometrics, 35(3):704– 705, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530269.

Nelder:1979:BRR

[1362] J. A. Nelder. Book review: Regression Analysis by Example, by S. Chatterjee, B. Price. Biometrics, 35(1):355–356, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529957.

Nelder:1979:BRS

[1363] J. A. Nelder. Book review: Statistical Analysis, a Computer Oriented Approach, by A. A. Afifi, S. P. Azen. Biometrics, 35(3):703, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530266.

Nicholson:1979:QEM

[1364] M. D. Nicholson. Query: On expressing the mean of rounded data. Biometrics, 35(4):873–874, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530122.

North:1979:MHS

[1365] Philip M. North and Byron J. T. Morgan. Modelling heron survival using weather data. Biometrics, 35(3):667–681, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530260.

OBrien:1979:MTP

[1366] Peter C. O’Brien and Thomas R. Fleming. A multiple testing procedure for clinical trials. Biometrics, 35(3):549–556, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530245.

Ord:1979:BRA

[1367] J. K. Ord. Book review: Applied Time Series Analysis, by D. F. Findley. Biometrics, 35(3):703–704, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530267.

Patil:1979:NEP

[1368] S. A. Patil, K. P. Burnham, and J. L. Kovner. Nonparametric estimation of plant density by the distance method. Biometrics, 35(3):597–604, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530250.

Paul:1979:CBB

[1369] S. R. Paul. A clumped beta-binomial model for the analysis of clustered attribute data. Biometrics, 35(4):821–824, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530115.

Pearce:1979:BRS

[1370] S. C. Pearce. Book review: Statistics for Experimenters: an Introduction to Design, Data Analysis and Model Building, by G. E. P. Box, W. G. Hunter, J. S. Hunter. Biometrics, 35(2):528–529, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530361.

Perry:1979:PID

[1371] J. N. Perry and R. Mead. On the power of the index of dispersion test to detect spatial pattern. Biometrics, 35(3):613–622, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530252.

Pielou:1979:BRF

[1372] E. C. Pielou. Book review: Food Webs and Niche Space, by J. E. Cohen. Biometrics, 35(1):356–357, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529959.

Pierce:1979:DFR

[1373] Donald A. Pierce, William H. Stewart, and Kenneth J. Kopecky. Distribution-free regression analysis of grouped survival data. Biometrics, 35(4):785–793, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530110.

Pocock:1979:APT

[1374] S. J. Pocock. Allocation of patients to treatment in clinical trials. Biometrics, 35(1):183–197, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2529944.

Pollak:1979:SMG

[1375] Edward Pollak. Some models of genetic selection. Biometrics, 35(1):119– 137, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2529940.

Poole:1979:BRG

[1376] R. W. Poole. Book review: The Golden Age of Theoretical Ecology: 1923–1940. A Collection of Works by V. Volterra, V. A. Kostitzin, A. J. Lotka, and A. N. Kolmogoroff, by F. M. Scudo, J. R. Ziegler, V. Volterra, V. A. Kostitzin, A. J. Lotka, A. N. Kolmogoroff. Biometrics, 35(3):706–707, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530273.

Prentice:1979:HRM

[1377] Ross L. Prentice and J. D. Kalbfleisch. Hazard rate models with covariates. Biometrics, 35(1):25–39, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2529934.

Prentice:1979:QDB

[1378] R. L. Prentice and P. Marek. A qualitative discrepancy between censored data rank tests. Biometrics, 35(4):861–867, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530120.

Prescott:1979:MGT

[1379] Robin Prescott and Jun mo Nam. On McNemar’s and Gart’s tests for cross-over clinical trials. Biometrics, 35(2):523–524, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530356.

Purcell:1979:BIP

[1380] Noel J. Purcell and Leslie Kish. A biometrics invited paper. estimation for small domains. Biometrics, 35(2):365–384, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530340.

Purcell:1979:ESD

[1381] Noel J. Purcell and Leslie Kish. Estimation for small domains. Biometrics, 35(2):365–384, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reed:1979:EDC

[1382] George F. Reed and Richard B. McHugh. The estimation of dental caries incidence in the presence of diagnostic error. Biometrics, 35(2):473– 478, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530349.

Regal:1979:CIA

[1383] R. R. Regal and E. J. Cushing. Confidence intervals for absolute pollen counts. Biometrics, 35(3):557–565, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530246.

Rice:1979:AMN

[1384] J. Rice, T. Reich, C. R. Cloninger, and R. Wette. An approximation to the multivariate normal integral: its application to multifactorial qualitative traits. Biometrics, 35(2):451–459, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530347.

Rosner:1979:STI

[1385] B. Rosner, A. Donner, and C. H. Hennekens. Significance testing of interclass correlations from familial data. Biometrics, 35(2):461–471, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530348.

Sandland:1979:BRF

[1386] R. L. Sandland. Book review: Fisheries Mathematics, by J. H. Steele. Biometrics, 35(1):358–359, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529964.

Sandland:1979:SGC

[1387] R. L. Sandland and C. A. McGilchrist. Stochastic growth curve analysis. Biometrics, 35(1):255–271, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529948.

Schenzle:1979:FTN

[1388] D. Schenzle. Fitting the truncated negative binomial distribution without the second sample moment. Biometrics, 35(3):637–639, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530255.

Schoenfeld:1979:DAC

[1389] D. A. Schoenfeld and R. D. Gelber. Designing and analyzing clinical trials which allow institutions to randomize patients to a subset of the treatments under study. Biometrics, 35(4):825–829, December 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530116.

Shah:1979:MIL

[1390] B. K. Shah. On the method of internal least squares. Biometrics, 35(2): 497–502, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530353.

Shoemaker:1979:OTM

[1391] Christine A. Shoemaker. Optimal timing of multiple applications of pesticides with residual toxicity. Biometrics, 35(4):803–812, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530112.

Simon:1979:RRD

[1392] Richard Simon. Restricted randomization designs in clinical trials. Biometrics, 35(2):503–512, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2530354.

Smith:1979:BRM

[1393] C. A. B. Smith. Book review: Mathematische Grundlagen der Genetik, by E. Weber. Biometrics, 35(2):532, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530367.

Smith:1979:SAL

[1394] G. E. J. Smith. Some aspects of line transect sampling when the target populations moves. Biometrics, 35(1):323–329, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529953.

Snee:1979:UMA

[1395] R. D. Snee, S. K. Acuff, and James R. Gibson. A useful method for the analysis of growth studies. Biometrics, 35(4):835–848, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530118.

Sprent:1979:BRM

[1396] P. Sprent. Book review: The Measurement of Biological Shape and Shape Change, by F. L. Bookstein. Biometrics, 35(2):528, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2530360.

Steel:1979:BRB

[1397] G. G. Steel. Book review: Biomathematics and Cell Kinetics, by A. J. Valleron, P. D. M. Macdonald. Biometrics, 35(2):531–532, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530366.

Stoddard:1979:SMP

[1398] Anne M. Stoddard. Standardization of measures prior to cluster analysis. Biometrics, 35(4):765–773, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530108.

Tarter:1979:BMA

[1399] Michael E. Tarter. Biocomputational methodology: an adjunct to theory and applications. Biometrics, 35(1):9–24, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529933.

Taulbee:1979:GMH

[1400] John D. Taulbee. A general model for the hazard rate with covariables. Biometrics, 35(2):439–450, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2530346.

Tavare:1979:NFH

[1401] S. Tavaré. A note on finite homogeneous continuous-time Markov chains. Biometrics, 35(4):831–834, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530117.

Thompson:1979:SE

[1402] R. Thompson. Sire evaluation. Biometrics, 35(1):339–353, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2529955.

Tiede:1979:ARC

[1403] James J. Tiede and Marcello Pagano. The application of robust calibration to radioimmunoassay. Biometrics, 35(3):567–574, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530247.

Upton:1979:BRA

[1404] G. J. G. Upton. Book review: Analyzing Qualitative Categorical Data, by L. A. Goodman. Biometrics, 35(2):530, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2530364.

Usher:1979:BRC

[1405] M. B. Usher. Book review: Caste and Ecology in the Social Insects, by G. F. Oster, E. O. Wilson. Biometrics, 35(4):902–903, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530130.

Usher:1979:BRD

[1406] M. B. Usher. Book review: The Dynamics of Arthropod Predator-Prey Systems, by M. P. Hassell. Biometrics, 35(4):901–902, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530127.

Warren:1979:QAI

[1407] W. G. Warren. Query: Analysis and interpretation of an experiment with a heterogeneous mixture of treatment types. Biometrics, 35(4):869–872, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2530121.

Weir:1979:IAL

[1408] B. S. Weir. Inferences about linkage disequilibrium. Biometrics, 35(1): 235–254, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2529947.

Weiss:1979:WRS

[1409] Gary B. Weiss and Harvey Bunce Iii. Are we ready for statistical guidelines for medical research papers? Biometrics, 35(4):911, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530141.

Wermuth:1979:DSE

[1410] Nanny Wermuth and W. G. Cochran. Detecting systematic errors in multi-clinic observational data. Biometrics, 35(3):683–686, September 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530261.

Westlake:1979:SAC

[1411] W. J. Westlake. Statistical aspects of comparative bioavailability trials. Biometrics, 35(1):273–280, March 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2529949.

Williams:1979:BIP

[1412] J. S. Williams. A biometrics invited paper. A synthetic basis for a comprehensive factor-analysis theory. Biometrics, 35(4):719–733, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530103.

Williams:1979:SBC

[1413] J. S. Williams. A synthetic basis for a comprehensive factor-analysis theory. Biometrics, 35(4):719–733, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolde-Tsadik:1979:CVS

[1414] Girma Wolde-Tsadik and Mimi C. Yu. Concordance in variable-subset discriminant analysis. Biometrics, 35(3):641–644, September 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530256.

Woodbury:1979:LAD

[1415] Max A. Woodbury, Kenneth G. Manton, and Eric Stallard. Longitudinal analysis of the dynamics and risk of coronary heart disease in the Framingham Study. Biometrics, 35(3):575–585, September 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2530248.

Wu:1979:URM

[1416] Margaret Wu and James H. Ware. On the use of repeated measurements in regression analysis with dichotomous responses. Biometrics, 35(2): 513–521, June 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2530355.

Anonymous:1980:CCD

[1417] Anonymous. Corrections: Correction to ‘Detection of a treatment effect when not all experimental subjects will respond to treatment’. Biometrics, 36(4):751, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2556140. See [1330].

Anonymous:1981:CAT

[1418] Anonymous. Correction: Analysis of the Two-Period Repeated Measurements Crossover Design with Application to Clinical Trials. Biometrics, 37(4):875, December 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2530201. See [1105].

Anonymous:1983:CPC

[1419] Anonymous. Corrections: Power computations for designing comparative Poisson trials. Biometrics, 39(4):1137, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531373. See [658].

Anonymous:1987:CNA

[1420] Anonymous. Correction: 333. Note: The analysis of non-independent 2×2 tables using rank sums. Biometrics, 43(1):259, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2531986. See [353].

Anonymous:1988:CEH

[1421] Anonymous. Corrections: Estimation of heritability from experiments with inbred and related individuals. Biometrics, 44(1):321, March 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2531938. See [222].